### v1

# Causal Inference with High-dimensional Discrete Covariates

Abstract
When estimating causal effects from observational studies, researchers often need to adjust for many covariates to deconfound the non-causal relationship between exposure and outcome, among which many covariates are discrete. The behavior of commonly used estimators in the presence of many discrete covariates is not well understood since their properties are often analyzed under structural assumptions including sparsity and smoothness, which do not apply in discrete settings. In this work, we study the estimation of causal effects in a model where the covariates required for confounding adjustment are discrete but high-dimensional, meaning the number of categories $d$ is comparable with or even larger than sample size $n$. Specifically, we show the mean squared error of commonly used regression, weighting and doubly robust estimators is bounded by $\frac{d^{2}}{n^{2}}+\frac{1}{n}$. We then prove the minimax lower bound for the average treatment effect is of order $\frac{d^{2}}{n^{2}\log^{2}n}+\frac{1}{n}$, which characterizes the fundamental difficulty of causal effect estimation in the high-dimensional discrete setting, and shows the estimators mentioned above are rate-optimal up to log-factors. We further consider additional structures that can be exploited, namely effect homogeneity and prior knowledge of the covariate distribution, and propose new estimators that enjoy faster convergence rates of order $\frac{d}{n^{2}}+\frac{1}{n}$, which achieve consistency in a broader regime. The results are illustrated empirically via simulation studies.
Keywords: average treatment effects, categorical data, high-dimensionality, effect homogeneity, minimax lower bounds.
1 Introduction
Researchers performing causal inference based on observational studies often have access to high-dimensional data with a substantial number of samples and many covariates. To draw causal conclusions from such data, typically a large set of covariates (which are potentially confounders that affect both treatment assignment and outcome of interest) needs to be measured and adjusted for. Various approaches exist for estimating causal effects from observational studies. Classic methods include outcome regression (Rubin,, 1979; Hernán and Robins,, 2010), inverse propensity score weighting (Rosenbaum and Rubin,, 1983; Hahn,, 1998; Hirano et al.,, 2003) and doubly robust estimation (also known as augmented inverse probability weighting) (Bang and Robins,, 2005; Kennedy,, 2022). Notably the doubly robust estimator is a consistent estimator of average treatment effect (ATE) even when one of the outcome model or propensity score is misspecified, providing a robust tool for estimating ATE under possible model misspecification. It also achieves semiparametric efficiency bound when both the outcome model and propensity score are correctly specified under regularity conditions. Recent state-of-the-art methods to estimate ATE include double/debiased machine learning (Chernozhukov et al.,, 2018), covariate balancing (Athey et al.,, 2018; Imai and Ratkovic,, 2014; Li et al.,, 2018) and calibration-based estimation (Tan, 2020a, ; Tan, 2020b, ). All these methods require estimating at least one of the outcome model or propensity score. Techniques such as sample splitting or cross-fitting (Chernozhukov et al.,, 2018) are often used to prevent overfitting the nuisance functions and avoid empirical process conditions (Kennedy,, 2022). In the classic setting where the dimension of data (i.e., number of covariates) is fixed, the aforementioned methods and techniques enjoy appealing theoretical properties (e.g., $\sqrt{n}$-consistency and asymptotic normality) and performance in real applications. However, when the number of covariates we need to adjust for is large, additional problems and concerns arise.
Recently there has been much work in the causal inference literature focused on understanding the behavior and properties of various estimation procedures in the high-dimensional setting. Yadlowsky, (2022) derived the excess variance of commonly used estimators due to first-stage nuisance estimation with high-dimensional covariates by assuming a linear outcome model and known propensity score. They also performed a simulation study to show the inflation of variance empirically, hence even doubly robust estimator may not achieve the efficiency bound when the covariates included are high-dimensional. Jiang et al., (2022) established a novel central limit theorem for the doubly robust estimator where the outcome regression model and propensity score are fitted with high-dimensional generalized linear models without assumptions on sparsity, under a stylized setting where covariates are normally distributed and covariate dimension $d$ is of the same order as sample size $n$. They also showed estimates obtained by permuting the folds in cross-fitting are asymptotically correlated in the high-dimensional regime. Celentano and Wainwright, (2023) further proposed a novel debiased method for missing data models in the case $n<d$, where ordinary least square estimation is not feasible. To help overcome the curse of dimensionality and achieve non-trivial rates of convergence, most of the work in the literature combines the debiased machine learning/doubly robust techniques with additional structural assumptions on the nuisance functions, including smoothness (Robins et al.,, 2009; Kennedy,, 2020; Semenova and Chernozhukov,, 2021) or sparsity (Belloni et al.,, 2017; Chernozhukov et al.,, 2018; Athey,, 2018; Bradic et al.,, 2019). See also Maathuis et al., (2009); Lin, (2013); Zhao, (2016); Chakrabortty et al., (2018); Ma et al., (2019); Lei and Ding, (2021); Antonelli et al., (2022); Tang et al., (2023); Du et al., (2024) for relevant discussion on causal inference with high-dimensional data.
Our work enriches the high-dimensional causal literature under another distributional assumption on the covariates, i.e., the covariates are discrete. This stylized setting may uncover properties of different estimators in the high-dimensional regime, where the number of categories $d$ may be comparable with or larger than sample size $n$, and motivate further studies into this regime. The discrete covariate setting can also be viewed as an interesting base case that also provides implications for continuous data, including testing the data generating distribution (Balakrishnan and Wasserman,, 2019) and continuum-armed bandit problems (Kleinberg,, 2004; Liu et al.,, 2021). In applications of medical and health policy research, researchers often adjust for International Classification of Diseases (ICD) codes (Organization,, 2004), which include a large number of indicators. The cartesian product of these binary indicators together with other commonly adjusted covariates (e.g., sex, race, education level) often induces a large number of categories, which corresponds to the high-dimensional regime. Our work establishes the theoretical foundation of adjusting for high-dimensional discrete covariates and helps practitioners evaluate the reliability of their estimates given the sample size and dimension of data.
Estimation of treatment effect under a discrete setting is relevant to estimating functionals of discrete distributions (e.g., entropy, support size) (Paninski,, 2003; Valiant and Valiant,, 2010; Jiao et al.,, 2015; Jiang and Qi,, 2015; Wu and Yang,, 2016, 2019). Recently, important breakthroughs have been made in terms of both improved estimation methodologies and fundamental limits. Using tools in polynomial approximation theory (Timan,, 2014), a best-polynomial-approximation estimator usually enjoys the so-called “effective sample size enlargement” property, meaning its behavior with $n$ samples resembles that of MLE (the naive plug-in estimator) with $n\log n$ samples. The corresponding lower bound is established by a moment matching technique. Estimation based on polynomial approximation theory and deriving minimax lower bounds using moment matching usually yield matching upper and lower bounds since, in the optimization sense, moment matching is the dual problem of best polynomial approximation where strong duality holds. The readers are referred to Wu and Yang, (2016); Luenberger, (1997) for more detailed discussion on this duality phenomenon. While the estimation of causal functionals under the discrete covariate framework has remained largely unexplored, our work helps bridge this gap by applying and extending the techniques in this framework to the estimation of treatment effects.
Specifically, in this paper we study the treatment effects estimation problem when adjusting for discrete covariates with possibly more categories than samples. Our main contributions are summarized as follows:
  • First, we provide finite-sample bounds on the mean squared error of commonly used regression, weighting and doubly robust estimators of ATE estimation. Our results imply $d/n\rightarrow 0$ is a sufficient and necessary condition for these estimators to be consistent under positivity assumption, where $d$ is the number of categories and $n$ is the sample size.
  • We then characterize the fundamental difficulty of ATE estimation under positivity assumption. The minimax lower bound is (in terms of mean squared error) of order $\frac{d^{2}}{n^{2}\log^{2}n}+\frac{1}{n}$, which shows the commonly used estimators aforementioned are minimax optimal up to log-factors. Moreover, $d=o(n\log n)$ is a necessary condition for consistent ATE estimation.
  • Next, we explore the role of effect homogeneity in ATE estimation and prove a faster convergence rate when the treatment effects are homogeneous across categories, which enables consistent estimation if $d/n^{2}\rightarrow 0$.
  • We further study the estimation of treatment effects when prior knowledge of covariate distribution is available. We first provide negative results on ATE estimation, i.e., knowledge of covariate distribution may not help us improve the rate. We then consider the variance-weight average treatment effect and derive its faster estimation rate achieved by a second-order estimator. Similar to the homogeneous effects setting, the second-order estimator is consistent if $d/n^{2}\rightarrow 0$.
The structure of our paper is as follows: After introducing the data-generating process, causal assumptions and notation in Section 2, in Section 3 we study the properties of commonly used regression, weighting and doubly robust estimators in estimating the average treatment effects (ATE). We first show the numerical equivalence between these three estimators and the plug-in estimator. We then provide finite-sample bounds on their mean squared error and characterize the sufficient and necessary conditions for them to be consistent in Section 3.1. In Section 3.2 we study the minimax lower bound and sample complexity of ATE estimation in the high-dimensional setting, based on the moment matching method borrowed from the theoretical computer science literature. Next, we consider two additional structures that one can exploit in the high-dimensional problem: effect homogeneity (Section 4) and prior knowledge of covariate distribution (Section 5). We propose novel estimators that can take advantage of these structures and enjoy faster convergence rates, making consistent estimation possible in the regime where conventional estimators examined in Section 3 fail to be consistent. Finally in Section 6 we perform numerical experiments to verify our results in previous sections. All the proofs and additional complementary results are presented in the supplementary materials. To the best of our knowledge, our work is the first in the literature that directly analyzes the theoretical properties of different estimators of causal effects, provides a minimax lower bound for ATE estimation and explores additional structures that we can exploit to achieve faster convergence rate in the high-dimensional discrete setting.
2 Setup & Notation
In this section, we first introduce the data-generating process in the covariate setting and characterize the distributions of several counting statistics that play an important role in estimating causal effects. Identification assumptions of causal estimands and additional notation are further introduced with discussion.
2.1 Data Generating Process
Suppose we observe $n$ i.i.d. copies of $\mathbf{Z}=(X,A,Y)$ where $X$ is the discrete covariate with $d$ categories, $A\in\{0,1\}$ is the binary treatment and $Y\in\{0,1\}$ is the binary outcome. On the population level, the covariate $X$ has a categorical distribution on $[d]=\{1,\dots,d\}$ with
[…]
Let ${\boldsymbol{p}}=(p_{1},\dots,p_{d})$ be the probability vector. In real applications one may observe multiple discrete covariates. For example, we may observe $K$ binary covariates and it is easy to see that one could encode these binary variables as one categorical variable with $d=2^{K}$ categories. Hence we will assume a single discrete covariate $X$ in our problem. Given the covariate $X=k$, the treatment $A$ follows a Bernoulli distribution with parameter $\pi_{k}$
$A\mid X=k\,\sim\text{Bernoulli}(\pi_{k})\text{ with }\pi_{k}=\mathbb{P}(A=1\mid X=k).$
Conditioned on the covariate $X=k$ and the treatment $A=a$, $Y$ has a Bernoulli distribution with parameter $\mu_{ak}$
$Y\mid A=a,X=k\,\sim\text{Bernoulli}(\mu_{ak})\text{ with }\mu_{ak}=\mathbb{P}(Y=1\mid X=k,A=a).$
$\pi_{k}$ and $\mu_{ak}$ are the propensity score and regression functions, respectively. We further denote $q_{ak}=\mathbb{P}(X=k,A=a,Y=1)=p_{k}\left[a\pi_{k}+(1-a)\left(1-\pi_{k}\right)\right]\mu_{ak}$ and $w_{k}=\mathbb{P}(X=k,A=1)=p_{k}\pi_{k}$. Based on a sample of size $n$, define the following empirical average estimators of the parameters:
$n\widehat{q}_{ak}=$ $\,\sum_{i=1}^{n}I\left(X_{i}=k,A_{i}=a,Y_{i}=1\right)\sim\operatorname{Bin}\left(n,q_{ak}\right),$ (1)
$n\widehat{w}_{k}=$ $\,\sum_{i=1}^{n}I\left(X_{i}=k,A_{i}=1\right)\sim\operatorname{Bin}\left(n,w_{k}\right),$
$n\widehat{p}_{k}=$ $\,\sum_{i=1}^{n}I\left(X_{i}=k\right)\sim\operatorname{Bin}\left(n,p_{k}\right),$
for each $1\leq k\leq d$. The corresponding empirical average estimates of the propensity score $\pi_{k}$ and regression function $\mu_{ak}$ are
$\widehat{\pi}_{k}=$ $\,\frac{\widehat{w}_{k}}{\widehat{p}_{k}}=\frac{\#\{i:X_{i}=k,A_{i}=1\}}{\#\{i:X_{i}=k\}},$ (2)
$\widehat{\mu}_{1k}=$ $\,\frac{\widehat{q}_{1k}}{\widehat{w}_{k}}=\frac{\#\{i:X_{i}=k,A_{i}=1,Y_{i}=1\}}{\#\{i:X_{i}=k,A_{i}=1\}},$
$\widehat{\mu}_{0k}=$ $\,\frac{\widehat{q}_{0k}}{\widehat{p}_{k}-\widehat{w}_{k}}=\frac{\#\{i:X_{i}=k,A_{i}=0,Y_{i}=1\}}{\#\{i:X_{i}=k,A_{i}=0\}},$
where we define $/0=0$ 0 whenever both the numerator and denominator are zero. This may happen when no individual in $k$-th category is assigned to treatment ($\widehat{w}_{k}=0$) and hence the response $Y$ under treatment is unavailable ($\widehat{q}_{1k}=0$). Let $\mathbf{X}^{n}=(X_{1},\dots,X_{n})$ and $\mathbf{A}^{n}=(A_{1},\dots,A_{n})$ be the collection of covariates and treatments in the sample, respectively. According to the sampling schemes we have
$(n\widehat{p}_{1},\dots,n\widehat{p}_{d})\,\sim\text{Multinomial}(n,p_{1},\dots,p_{d}),$
$n\widehat{w}_{k}=n\widehat{p}_{k}\widehat{\pi}_{k}\mid\mathbf{X}^{n}\sim\operatorname{Bin}\left(n\widehat{p}_{k},\pi_{k}\right),$
$n\widehat{q}_{ak}=n\widehat{p}_{k}\left[a\widehat{\pi}_{k}+(1-a)\left(1-\widehat{\pi}_{k}\right)\right]\widehat{\mu}_{ak}\mid\mathbf{X}^{n},\mathbf{A}^{n}\sim\operatorname{Bin}\left(n\widehat{p}_{k}\left[a\widehat{\pi}_{k}+(1-a)\left(1-\widehat{\pi}_{k}\right)\right],\mu_{ak}\right).$
[…]
2.2 Causal Assumptions and Other Notation
To properly define and identify causal estimands of interests, we rely on the potential outcome framework (Rubin,, 1974; Splawa-Neyman et al.,, 1990) and additional identification assumptions to connect counterfactual outcomes with the observed data. We use the random variable $Y^{a}$ to denote the potential/counterfactual outcome we would have observed had a subject received treatment $A=a$, which may be contrary to the observation $Y$. The average treatment effect (ATE) $\psi$ is defined as
$\psi=\mathbb{E}\left[Y^{1}-Y^{0}\right].$
The following identification assumptions are often imposed to identify $\psi$ as a functional of the observed data distribution $\mathbb{P}$.
Assumption 1.
Consistency: $Y=Y^{a}\text{\, if \,}A=a.$
Assumption 2.
Positivity: For any $k\in[d]$, $\pi_{k}\in[\epsilon,1-\epsilon]$ for some constant $\epsilon\in(0,1/2)$.
Assumption 3.
[…]
Under Assumptions 13, the ATE can be identified as
$\psi=\mathbb{E}\left[\mathbb{E}(Y\mid X,A=1)-\mathbb{E}(Y\mid X,A=0)\right]=\sum_{k=1}^{d}p_{k}(\mu_{1k}-\mu_{0k})=\sum_{k=1}^{d}p_{k}\left(\frac{q_{1k}}{w_{k}}-\frac{q_{0k}}{p_{k}-w_{k}}\right).$ (3)
We refer the readers to Hernán and Robins, (2010) for detailed discussion on these identification assumptions. From this point forward, $\psi$ will denote the functional of observed data distribution in (3), which is equal to ATE when Assumptions 13 hold. If Assumption 13 are violated, the functional $\psi$ should only be viewed as the expected difference in the regression functions between the treatment and control group, which may not represent a causal effect. Nonetheless, all our results still hold for $\psi$ in (3) under only the positivity assumption.
In this paper we will consider the following model:
$\mathcal{D}(\epsilon)=\left\{\sum_{k=1}^{d}p_{k}=1,0\leq p_{k}\leq 1,\epsilon\leq\pi_{k}\leq 1-\epsilon,0\leq\mu_{0k},\mu_{1k}\leq 1,\,\forall 1\leq k\leq d\right\}.$ (4)
We will always assume positivity in addition to the basic bounds on model parameters. The binary nature of the outcome variable $Y$ is not essential to our analysis; we anticipate that our results will remain valid as long as $Y$ is bounded.
Remark 1.
It is worth noting that in the high-dimensional regime where $d$ can be large (the regime we are mainly interested in), the positivity assumption 2 can impose additional restriction on the observed distribution (D’Amour et al.,, 2021). However, when positivity is violated, the ATE is not identified and may not be an appropriate estimand to focus on. This together with other possible issues due to violation of positivity are tangential to our main points and thus we may proceed assuming positivity holds. Some of our results characterize the rate in terms of $\epsilon$ explicitly (e.g., Theorem1) and still provide meaningful implications under weak positivity assumption where $\epsilon=\epsilon_{n}$ shrinks to zero.
Remark 2.
Note that in the model $\mathcal{D}(\epsilon)$ some proportions $p_{k}$ may be zero, so the support size $s=|\{k:p_{k}>0\}|$ can be less than $d$. However, when deriving upper bounds in Section 35, we assume $p_{k}>0$ for all $k\in[d]$ so that $d$ is also the support size of $X$. Given $\mathbb{P}\in\mathcal{D}(\epsilon)$ with $p_{k}=0$ for some $k\in[d]$, one could remove categories with $p_{k}=0$ and the results still hold with $d$ replaced by the support size $s$. Since $s$ is clearly upper bounded by $d$, all the rate requirements on $(n,s)$ (e.g., $s/n\rightarrow 0$) can be stated in $(n,d)$ (e.g., $d/n\rightarrow 0$) as well.
[…]
3 Average Treatment Effects
In this section, we focus on the theoretical properties of commonly used estimators of ATE
$\psi=\mathbb{E}\left[\mathbb{E}(Y\mid X,A=1)-\mathbb{E}(Y\mid X,A=0)\right]=\sum_{k=1}^{d}p_{k}(\mu_{1k}-\mu_{0k})=\sum_{k=1}^{d}p_{k}\left(\frac{q_{1k}}{w_{k}}-\frac{q_{0k}}{p_{k}-w_{k}}\right).$
A simple plug-in style estimator based on empirical average estimates of model parameters $(p_{k},\pi_{k},\mu_{1k},\mu_{0k})$ in (1) and (2) is
$\widehat{\psi}=\sum_{k=1}^{d}\widehat{p}_{k}\left(\widehat{\mu}_{1k}-\widehat{\mu}_{0k}\right)=\sum_{k=1}^{d}\widehat{p}_{k}\left(\frac{\widehat{q}_{1k}}{\widehat{w}_{k}}-\frac{\widehat{q}_{0k}}{\widehat{p}_{k}-\widehat{w}_{k}}\right),$ (5)
We also consider three popular estimators in the literature to estimate ATE, namely the outcome regression estimator $\widehat{\psi}_{\text{reg}}$ (Rubin,, 1979), inverse probability weighting estimator $\widehat{\psi}_{\text{ipw}}$ (Rosenbaum and Rubin,, 1983; Hahn,, 1998) and doubly robust estimator $\widehat{\psi}_{\text{dr}}$ (Robins et al.,, 1994; Scharfstein et al.,, 1999) defined as follows:
$\widehat{\psi}_{\text{reg}}=$ $\,\mathbb{P}_{n}\left[\widehat{\mu}_{1X}-\widehat{\mu}_{0X}\right],$ (6)
$\widehat{\psi}_{\text{ipw}}=$ $\,\mathbb{P}_{n}\left[\frac{AY}{\widehat{\pi}_{X}}-\frac{(1-A)Y}{1-\widehat{\pi}_{X}}\right],$
$\widehat{\psi}_{\text{dr}}=$ $\,\mathbb{P}_{n}\left[\frac{A(Y-\widehat{\mu}_{1X})}{\widehat{\pi}_{X}}+\widehat{\mu}_{1X}-\frac{(1-A)(Y-\widehat{\mu}_{0X})}{1-\widehat{\pi}_{X}}-\widehat{\mu}_{0X}\right],$
where again in the IPW and DR estimator we define $/0=0$ 0 whenever it occurs. In this paper, we will focus on the observational study setting where propensity score is unknown and needs to be estimated. In randomized experiments where the treatment process is known, the IPW estimator with known propensity scores is unbiased and $\sqrt{n}$-consistent.
Surprisingly, all these three commonly used estimators of ATE are numerically equivalent to the simple plug-in estimator (5) in the discrete covariate setting, if we use the same sample to estimate the nuisance parameters in (2) and take average over in (6). This property unique to the discrete covariate setting is presented in regression coefficients estimation with missing covariates in Wang et al., (2007). Recently similar equivalence in ATE estimation is shown in Słoczyński et al., (2023) when the correct parametric model for propensity score is specified. The discrete covariate setting can be viewed as a special case where the design matrix only contains indicators specifying the category membership of each sample. We restate their result in the discrete covariate setting and emphasize that such equivalence still holds in the high-dimensional setting where some categories may not have treated/untreated samples observed, as long as we define $/0=0$ 0 whenever it appears.
Proposition 1.
Supposed $\mathcal{D}=\{(X_{i},A_{i},Y_{i}),1\leq i\leq n\}$ is a sample of size $n$ and $X\in[d]$ is discrete. If the nuisance estimators are the empirical averages defined in equation (2) using $\mathcal{D}$ and we take average over $\mathcal{D}$ in (6). Then we have
$\widehat{\psi}_{\text{reg}}=\widehat{\psi}_{\text{ipw}}=\widehat{\psi}_{\text{dr}}=\widehat{\psi}.$
So the regression, weighting and doubly robust estimators are numerically equivalent in the discrete covariate setting.
The numerical equivalence in Proposition 1 does not necessarily hold when the covariates have continuous components and smoothing is applied to construct estimates of propensity scores and regression functions. One explanation is that in the discrete case, the arguments used to show $\mathbb{E}[{\mu}_{1X}]=\mathbb{E}[AY/\pi_{X}]=\psi_{1}$ (See e.g., Chapter 2 in Hernán and Robins, (2010)) can be applied to the empirical distribution $\mathbb{P}_{n}$ (i.e. replace every expectation and conditional expectation in the argument with sample averages) to show $\mathbb{P}_{n}[\widehat{\mu}_{1X}]=\mathbb{P}_{n}[AY/\widehat{\pi}_{X}]=\widehat{\psi}$.
Proposition 1 shows that we only need to consider the properties of plug-in style estimator $\widehat{\psi}$ and all the results hold for estimators in (6) as well. In the following discussion, we will first derive the upper bound on the mean squared error of $\widehat{\psi}$ in Section 3.1. Then we characterize the minimax lower bound in Section 3.2.
3.1 Upper Bound
In this section, we study the behavior of $\widehat{\psi}$ in a potentially high-dimensional regime and characterize its mean squared error in terms of $(n,d)$. In the low-dimensional regime where $d$ is fixed, the empirical average estimators of nuisance functions (2) have parametric convergence rates. One can expect this ideal property to propagate to plug-in style estimator $\widehat{\psi}$, which crucially depends on the empirical average estimates. We provide a central limit theorem of $\widehat{\psi}$ in Appendix A when $d$ is fixed to avoid distracting the readers from the high-dimensional regime of main interest. In the following discussions, our main results hold non-asymptotically for all $n,d$ sufficiently large. We will focus on the point estimation theory in this work and characterize the condition under which $\widehat{\psi}$ is consistent. Let $\psi_{a}=\mathbb{E}[Y^{a}]=\mathbb{E}[\mathbb{E}(Y\mid X,A=a)]=\sum_{k=1}^{d}p_{k}\mu_{ak}$ be the mean of potential outcome $Y^{a}$ and $\widehat{\psi}_{a}=\sum_{k=1}^{d}\widehat{p}_{k}\widehat{\mu}_{ak}$ be the corresponding plug-in style estimator. We first summarize the exact bias of $\widehat{\psi}_{a}$ in the following proposition.
Proposition 2.
The exact bias of $\widehat{\psi}_{a}$ is
$\mathbb{E}\left[\widehat{\psi}_{1}-\psi_{1}\right]=$ $\,-\sum_{k=1}^{d}\mu_{1k}p_{k}(1-\pi_{k})(1-p_{k}\pi_{k})^{n-1},$
$\mathbb{E}\left[\widehat{\psi}_{0}-\psi_{0}\right]=$ $\,-\sum_{k=1}^{d}\mu_{0k}p_{k}\pi_{k}(1-p_{k}+p_{k}\pi_{k})^{n-1}.$
Hence the exact bias of $\widehat{\psi}=\widehat{\psi}_{1}-\widehat{\psi}_{0}$ is
$\mathbb{E}\left[\widehat{\psi}-\psi\right]=-\sum_{k=1}^{d}\left[\mu_{1k}p_{k}(1-\pi_{k})(1-p_{k}\pi_{k})^{n-1}-\mu_{0k}p_{k}\pi_{k}(1-p_{k}+p_{k}\pi_{k})^{n-1}\right]$
From the proof of Proposition 2, the bias of $\widehat{\psi}_{a}$ comes from categories with no subjects receiving treatment $A=a$, and hence it is not possible to obtain an unbiased estimator of regression functions $\mu_{ak}$ within these categories. It is unclear from the exact bias how fast $d$ can grow with $n$ while still having vanishing bias. To make the dependency of bias on $(n,d)$ explicit, we maximize the bias over model $\mathcal{D}(\epsilon)$ and characterize the worst-case bias in the next proposition. In the following discussion of this section, we will mainly consider the functional $\psi_{1}$, with the understanding that analogous arguments can be applied to $\psi_{0}$ to obtain the same rate/property.
Proposition 3.
The worst-case bias of $\widehat{\psi}_{1}$ is
$\sup_{\mathbb{P}\in\mathcal{D(\epsilon)}}\big{|}\mathbb{E}_{\mathbb{P}}[\widehat{\psi}_{1}-\psi_{1}]\big{|}=\sup_{p_{k}}\sum_{k=1}^{d}p_{k}(1-\epsilon)(1-\epsilon p_{k})^{n-1}.$
As a consequence, we have the following bounds on the worst-case bias:
$(1-\epsilon)\exp\left(-\frac{\epsilon(n-1)}{d-\epsilon}\right)\leq\sup_{\mathbb{P}\in\mathcal{D(\epsilon)}}\big{|}\mathbb{E}_{\mathbb{P}}[\widehat{\psi}_{1}-\psi_{1}]\big{|}\leq\frac{1-\epsilon}{\epsilon}\frac{d}{n}.$
Furthermore if $n\geq 1+1/\epsilon$, then we have
$\sup_{\mathbb{P}\in\mathcal{D(\epsilon)}}\big{|}\mathbb{E}_{\mathbb{P}}[\widehat{\psi}_{1}-\psi_{1}]\big{|}\geq\frac{1-\epsilon}{8}\left(\frac{d-1}{\epsilon(n-1)}\wedge 1\right)$
Proposition 3 shows that a sufficient and necessary condition for the bias to vanish as $n\rightarrow\infty$ in the worst case over model class $\mathcal{D}(\epsilon)$ is $d/n\rightarrow 0$. The following theorem further characterizes a bound on the variance of $\widehat{\psi}_{1}$ and arrives at the MSE of the plug-in-style estimator.
Theorem 1.
The variance of $\widehat{\psi}_{1}$ is upper bounded as
$\sup_{\mathbb{P}\in\mathcal{D}(\epsilon)}\operatorname{Var}_{\mathbb{P}}(\widehat{\psi}_{1})\leq\frac{C}{\epsilon n},$
where $C>0$ is an absolute constant. Hence the MSE is upper bounded as
$\sup_{\mathbb{P}\in\mathcal{D}(\epsilon)}\mathbb{E}_{\mathbb{P}}\left[\left(\widehat{\psi}_{1}-\psi_{1}\right)^{2}\right]\leq\frac{(1-\epsilon)^{2}}{\epsilon^{2}}\frac{d^{2}}{n^{2}}+\frac{C}{\epsilon n}.$
Under positivity condition 2, the marginal probability for a subject to receive treatment is $\mathbb{P}(A=1)\geq\epsilon$. Thus on average the total number of treated samples is lower bounded by $\epsilon n$. Since the denominator of the bound on variance is exactly $\epsilon n$, intuitively $\epsilon n$ acts as the effective sample size for estimating $\psi_{1}$. The results in Proposition 3 and 1 imply:
  • The plug-in estimator $\widehat{\psi}_{1}$ is consistent in non-classical high-dimensional regimes where $d\rightarrow\infty$ as long as $d/n\rightarrow 0$ holds.
  • $d/n\rightarrow 0$ is also a necessary condition for the worst-case bias to vanish as $n\rightarrow\infty$. Thus consistency of $\widehat{\psi}_{1}$ in the high-dimensional regime $n\lesssim d$ is not achievable without further assumptions.
The intuition is that with $n$ samples and $d$ categories, there are $n/d$ samples within each category on average. Without further assumptions, one needs to consistently estimate the regression functions $\mu_{1k}$ in each category to achieve consistent estimation of $\psi_{1}=\sum_{k}p_{k}\mu_{1k}$. And consistent estimation of $\mu_{1k}$ requires infinite samples assigned to each category, i.e. $n/d\rightarrow\infty$. However, our results in this section do not preclude the existence of other consistent estimates. In order to conclusively determine whether consistent estimates exist in the high-dimensional regime, one needs to characterize the minimax lower bound for $\psi_{1}$.
3.2 Minimax Lower Bound
In Section 3.1 we showed that $d/n\rightarrow 0$ is a sufficient and necessary condition for the plug-in estimator $\widehat{\psi}$ to be consistent. In this section, we study the existence of consistent estimators in the high-dimensional regime by considering the minimax lower bound for the mean of potential outcome $\psi_{1}=\mathbb{E}[Y^{1}]$, with the understanding that similar arguments show $\psi_{0}=\mathbb{E}[Y^{0}]$ and ATE share the same minimax rate as $\psi_{1}$. Recall the model we consider is:
$\mathcal{D}(\epsilon)=\left\{0\leq p_{k}\leq 1,\epsilon\leq\pi_{k}\leq 1-\epsilon,0\leq\mu_{0k},\mu_{1k}\leq 1,\,\forall 1\leq k\leq d\right\},$
which corresponds to the setting where statisticians have knowledge on a set of possible values of $X$ but some categories may have zero proportion.
By Theorem 1, the MSE of plug-in-style estimator satisfies
$\mathbb{E}_{\mathbb{P}}\left[\left(\widehat{\psi}_{1}-\psi_{1}\right)^{2}\right]\lesssim\frac{d^{2}}{n^{2}}+\frac{1}{n},\forall\,\mathbb{P}\in\mathcal{D}(\epsilon),$ (7)
which shows a sufficient condition for the plug-in estimator to be consistent is $d/n\rightarrow 0$, i.e. we require sample size $n\gg d$ to achieve consistency. The following theorem characterizes the minimax lower bound of the estimation error of $\psi_{1}$ over model class $\mathcal{D}(\epsilon)$.
Theorem 2.
In the regime $d\lesssim n\log n$, we have
$R^{*}(d,n):=\inf_{\widehat{\psi}_{1}}\sup_{\mathbb{P}\in\mathcal{D}(\epsilon)}\mathbb{E}_{\mathbb{P}}\left[\left(\widehat{\psi}_{1}-\psi_{1}\right)^{2}\right]\gtrsim\left(\frac{d}{n\log n}\right)^{2}+\frac{1}{n},$
where the constant behind $\gtrsim$” depends on $\epsilon$.
The proof adopts the idea of moment matching commonly used in deriving minimax lower bounds for discrete functionals (Jiao et al.,, 2015; Wu and Yang,, 2016, 2019) and is included in the Appendix. Theorem 2 shows in the regime $d\gtrsim n\log n$, consistent estimator of $\psi_{1}$ does not exist over model class $\mathcal{D}(\epsilon)$. Compared with the upper bound in (7), we conclude that up to log-factors the plug-in estimator $\widehat{\psi}_{1}$ is minimax optimal and one needs sample size at least of order $d$ to consistently estimate $\psi_{1}$ over model class $\mathcal{D}(\epsilon)$. In applications, if the observed number of categories of the discrete covariate $X$ is comparable to sample size $n$, then we should interpret the estimated ATE carefully since it is possible that our estimator is not consistent in that high-dimensional regime.
It is worth noting that the lower bound in Theorem 2 does not exactly match the upper bound provided by the plug-in-style estimator with the difference being a log-factor. It is possible that some estimator of $\psi_{1}$ based on polynomial approximation, which further reduces the bias, enjoys the “effective sample size enlargement” property (Wu and Yang,, 2016; Jiao et al.,, 2015; Wu and Yang,, 2019) and achieves the minimax lower bound. We leave the exploration of such estimators to future investigation.
In the following two sections, we will consider two structures, namely effect homogeneity and prior knowledge of covariate distribution, that one can exploit to obtain faster rates and achieve consistent estimation of causal effects in the regime $n\lesssim d$ under certain scaling conditions on $n$ and $d$.
4 Role of Effect Homogeneity
In this section, we study the role of effect homogeneity in consistent estimation of treatment effects in the regime $n\lesssim d$. Let $\tau_{k}=\mathbb{E}[Y|X=k,A=1]-\mathbb{E}[Y|X=k,A=0]=\mu_{1k}-\mu_{0k}$ be the conditional average treatment effect (CATE) in the $k$-th category. In the high-dimensional regime where $d=d_{n}\rightarrow\infty$, $\tau_{k}$’s should be viewed as a triangular array $\{\tau_{nk},1\leq k\leq d_{n},n\geq 1\}$ and we will slightly abuse the notation to denote $\tau_{k}=\tau_{nk}$. Homogeneous effects (i.e. when $\tau_{k}=\psi$ for all $k\in[d]$) can be helpful in terms of estimation. Intuitively, since the treatment effects are the same within each level of covariate, there is no need to consistently estimate CATE $\tau_{k}$ for all possible $d$ categories. One only needs to estimate the CATE within a few categories accurately and due to effect homogeneity, these estimated CATEs generalize to other levels of covariate, which potentially reduces the sample size required for consistent estimation. In this work, we adopt a novel form of approximate effect homogeneity via parameter capturing the extent of heterogeneity. Mathematically, denote
$\sigma_{n}:=\max_{1\leq k\leq d_{n}}|\tau_{k}-\psi|$ (8)
as the maximal effect heterogeneity. Note that $\sigma_{n}=0$ corresponds to the constant conditional average treatment effects and $\sigma_{n}=2$ imposes no restriction on the model class. Our parameterization can interpolate between these two extremes as $\sigma_{n}$ varies.
The estimator we propose to take advantage of effect homogeneity is
$\widehat{\tau}=\frac{\sum_{k=1}^{d}\widehat{t}_{k}\widehat{\tau}_{k}}{\sum_{k=1}^{d}\widehat{t}_{k}}$ (9)
where $\widehat{t}_{k}=I(\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1)$ is the indicator of whether both treated and untreated samples are observed in the $k$-th level and $\widehat{\tau}_{k}=\widehat{\mu}_{1k}-\widehat{\mu}_{0k}$. The idea is to only estimate the CATE $\tau_{k}$ within those categories with both treated and untreated units and take the average over such categories. We restrict our attention to these categories so that unbiased estimation of $(\mu_{1k},\mu_{0k})$ (and hence $\tau_{k}$) is possible. We again define $/0=0$ 0 if $\sum_{k=1}^{d}\widehat{t}_{k}=0$, i.e. there is no such category that contains both treated and untreated units. Clearly on the event $\left\{\sum_{k=1}^{d}\widehat{t}_{k}=0\right\}$, we cannot obtain any information on treatment effects from $\widehat{\tau}$. Under appropriate scaling conditions, the probability of this “adverse” event converges to 00 even in the case $n\lesssim d$, as summarized in the following lemma.
Lemma 1.
The chance of having every category consist of either all treated or all untreated units (i.e., no collisions of any treated and untreated units at any category) is upper-bounded as
$\mathbb{P}\left(\sum_{k=1}^{d}\widehat{t}_{k}=0\right)\leq 2\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right)$
where $C(\epsilon)>0$ is a constant depending on $\epsilon$.
This lemma is not only critical to our analysis of $\widehat{\tau}$ but also of independent interests in the literature of applied probability. It is closely related to the birthday problem (Clevenson and Watkins,, 1991) and can be viewed as the occupancy problem (Wendl,, 2003; Nakata,, 2014) under a different sampling scheme. In the classic occupancy problem, the number of treated and untreated units are fixed first and then they are assigned to different categories of $X$ with probability vectors ${\boldsymbol{p}}^{\text{T}}$ and ${\boldsymbol{p}}^{\text{C}}$ separately. While in our setting people’s covariates are first sampled, following which their treatment assignments are determined. Importantly, Lemma 1 shows the probability that the denominator of $\widehat{\tau}$ is 00 vanishes even in the high-dimensional regime $n\lesssim d$, as long as $d/n^{2}\rightarrow 0$. With Lemma 1 we can derive the following theorem bounding the MSE of $\widehat{\tau}$.
Theorem 3.
The bias and variance of $\widehat{\tau}$ are bounded as
$|\mathbb{E}[\widehat{\tau}-\psi]|\leq\sigma_{n}+2\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right),$
$\operatorname{Var}(\widehat{\tau})\leq\mathbb{E}\left[\operatorname{Var}\left(\widehat{\tau}\big{|}\mathbf{X}^{n},\mathbf{A}^{n}\right)\right]+\sigma_{n}^{2}+2\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right)+4\sigma_{n}\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$
If we further assume that there exists a constant $\alpha\in(0,1]$ such that $p_{k}\geq\alpha/d$ for all $k\in[d]$, then in the regime $n\leq\frac{3\epsilon(1-\epsilon)d}{\alpha}$, we have
$\mathbb{E}\left[\operatorname{Var}\left(\widehat{\tau}\big{|}\mathbf{X}^{n},\mathbf{A}^{n}\right)\right]\leq\frac{4d}{\alpha^{2}\epsilon(1-\epsilon)n^{2}}+\frac{3}{2}\exp\left(-\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{2d}\right)+\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$
Hence the estimator is consistent if $\sigma_{n}\rightarrow 0$ and $d/n^{2}\rightarrow 0$ in the regime $n\lesssim d$.
The assumption $p_{k}\geq\alpha/d$ is required in our proof on bounding the expected conditional variance. We allow $p_{k}$ to approach zero but the rate should be slower than $1/d$. An important special case is $\alpha=1$ which corresponds to the uniform distribution on $[d]$. It is common to impose some constraints on the proportion probability of each level in categorical data analysis (Zeng et al., 2023a, ; Gu and Xu,, 2023). In the literature of regression adjustment with high-dimensional covariates, this assumption may also be required. For instance, Lei and Ding, (2021) considered regression adjustment estimation of sample average treatment effects and they impose conditions on the maximum leverage score of the projection matrix $\mathbf{H}=\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top}$, where $\mathbf{X}$ is the design matrix. Their condition translates to lower bounds on the proportions $p_{k}$’s when the covariates are discrete. Furthermore, another plausible condition that requires $p_{k}$’s are comparable
$\delta\leq p_{i}/p_{j}\leq\delta^{-1},\,i\neq j,$
with $\delta\in(0,1)$ being a constant implies $p_{k}\gtrsim 1/d$. Such conditions can be frequently found in the high-dimensional testing literature (Jiang and Qi,, 2015; He et al.,, 2021).
Theorem 3 shows that if asymptotic effect homogeneity holds (i.e. $\sigma_{n}\rightarrow 0$), then the estimator $\widehat{\tau}$ is consistent in the regime $d/n^{2}\rightarrow 0$. Compared with $d/n\rightarrow 0$ required for consistency by the plug-in estimator $\widehat{\psi}$, $\widehat{\tau}$ has a faster convergence rate and enables us to achieve consistency in a wider regime under asymptotic effect homogeneity.
5 Prior Knowledge of Covariate Distribution
In this section, we consider a different structure that could be exploited in causal effects estimation: the covariate distribution is known or can be estimated at a fast rate. The role of covariate distribution has been studied in both ATE estimation (Robins et al.,, 2008, 2017) and CATE estimation (Kennedy,, 2020), where an improved rate is achievable with knowledge of covariate density. However, when the covariate is discrete, faster rate of ATE estimation is not achievable with information on covariate distribution, as shown in Section 5.1. The fundamental difficulty is that the measure underlying ATE estimation is the product of covariate distribution and propensity score, whereas propensity score is unknown in our observational study setting. We thus switch attention to variance-weighted average treatment effects (WATE) in Section 5.2, which itself also receives an increasing amount of attention in the literature. The underlying measure of WATE is the covariate distribution and a faster rate of estimation is achieved with covariate information.
5.1 Average Treatment Effects
In this section, we introduce a minimax lower bound for $\psi_{1}=\mathbb{E}[\mathbb{E}(Y|X,A=1)]$ under the condition where $p_{k}=1/d$ for every $k\in[d]$, implying that when the covariate distribution is known to be uniform, the significant improvement on the rate of ATE estimation, such as $d/n^{2}$ discussed in Section 4, cannot be attained. This highlights the limitations of leveraging covariate distribution knowledge in improving the efficiency of ATE estimation. Consider the following model class where the covariate distribution is uniform over $[d]$:
$\mathcal{D}^{U}(\epsilon)=\left\{p_{k}=1/d,\epsilon\leq\pi_{k}\leq 1-\epsilon,0\leq\mu_{0k},\mu_{1k}\leq 1,\,\forall 1\leq k\leq d\right\}.$
The minimax lower bound of $\psi_{1}$ over model class $\mathcal{D}^{U}(\epsilon)$ is summarized in the following theorem.
Theorem 4.
For any fixed constant $\beta\in(0,1)$, in the regime $n\lesssim d^{1-\beta}$ we have
$\inf_{\widehat{\psi}_{1}}\sup_{\mathbb{P}\in\mathcal{D}^{U}(\epsilon)}\mathbb{E}_{\mathbb{P}}\left[\left(\widehat{\psi}_{1}-\psi_{1}\right)^{2}\right]\gtrsim 1,$
where the constant behind $\gtrsim$” depends on $\beta,\epsilon$.
Theorem 4 establishes that the minimax rate for estimating $\psi_{1}$ is lower bounded away from zero within the regime $n\lesssim d^{1-\beta}$, provided the model class includes $\mathcal{D}^{U}(\epsilon)$ as a subclass. This result indicates that enhancing the estimation rate of $\psi_{1}$ by incorporating knowledge of the covariate distribution may not yield significant improvements when $n\lesssim d$. Specifically, achieving consistency in the scenario where $d/n^{2}\rightarrow 0$—analogous to the situation of effect homogeneity discussed in Section 4—is unattainable solely with insights into the covariate distribution.
The negative result in Theorem 4 can be explained by the fact that the underlying measure of ATE estimation is the product of covariate distribution and propensity score, as shown in analyzing its second-order estimator (Robins et al.,, 2009; Zeng et al., 2023b, ). In observational studies where propensity score is unknown, it’s hard to obtain nearly unbiased estimates of $p_{k}\mu_{1k}$ (the summand in $\psi_{1}$) with the observed data. One may write $p_{k}\mu_{1k}=p_{k}\pi_{k}\mu_{1k}/\pi_{k}$ with $p_{k}\pi_{k}\mu_{1k}$ unbiasedly estimable. The extra term $1/\pi_{k}$ could be approximated by a polynomial with possibly diverging degrees
$\frac{1}{\pi_{k}}\approx\sum_{j=0}^{L}(1-\pi_{k})^{j},$
where $L=L_{n}$ determines the approximation error. Since the product $(p_{k}\pi_{k})^{j}$ can be estimated unbiasedly and $p_{k}$ is known, this strategy could reduce the bias. However, this approach suffers from a large variance, particularly in situations where $n\lesssim d$. In fact, our construction in the proof of Theorem 4 indeed replies on the difficulty of approximating $1/\pi_{k}$ with a polynomial of $\pi_{k}$, thereby highlighting the fundamental barrier in ATE estimation even with information on the covariate distribution.
5.2 Variance-weighted Average Treatment Effects
In this section we switch our attention to a popular variance-weighted average treatment effect estimand (WATE), defined as
$\theta=\frac{\mathbb{E}[\operatorname{Cov}(Y,A\mid X)]}{\mathbb{E}[\operatorname{Var}(A\mid X)]}=\frac{\mathbb{E}[\operatorname{Var}(A\mid X)\tau_{X}]}{\mathbb{E}[\operatorname{Var}(A\mid X)]},$
where the weight $\operatorname{Var}(A\mid X)$ minimizes the asymptotic variance among all weighted average treatment effects with known weight under homoskedasticity (Crump et al.,, 2006). Another interpretation of $\theta$ in terms of robustness is that if we assume a partial linear (homogeneous effect) model for the outcome and use E-estimator (Robins et al.,, 1992) to estimate the parametric component, under model misspecification (i.e. the partial linear model is wrong and effect of treatment is heterogeneous), one can show the E-estimator converges to WATE (Vansteelandt and Daniel,, 2014). The WATE has also been derived under different frameworks recently. Zhou and Opacic, (2022) showed the marginal interventional effect under incremental propensity score intervention (Kennedy,, 2019) coincides with WATE. The expected conditional covariance also appears in conditional independence test (Shah and Peters,, 2020) and can be interpreted as a causal effect under a stochastic intervention (Díaz,, 2023). In contrast to ATE, the underlying measure of $\theta$ is the covariate distribution alone, making improvements from knowledge of covariate distribution possible.
The strategy of estimating $\theta$ is to deal with the numerator and the denominator separately. Let $\eta=\mathbb{E}[\operatorname{Cov}(Y,A\mid X)]$ and $\rho=\mathbb{E}[\operatorname{Var}(A\mid X)]$. Further denote the regression function in $k$-th category as $\mu_{k}:=\mathbb{E}[Y|X=k]$. The first-order and second-order influence functions of $\eta$ under a nonparametric model, if they exist, have form
$\varphi_{1}(\mathbf{Z})=(A-\pi_{X})(Y-\mu_{X}),$
$\varphi_{2}(\mathbf{Z}_{1},\mathbf{Z}_{2})=-\frac{(A_{1}-\pi_{X_{1}})I(X_{1}=X_{2})(Y_{2}-\mu_{X_{2}})}{p_{X_{1}}}.$
Similar forms of influence functions of $\rho$ can be obtained by replacing $(Y,\mu)$ with $(A,\pi)$. However, in general settings where covariates have continuous components, the second-order influence function does not exist for $\eta$ (Robins et al.,, 2009). Intuitively, for two independent samples $X_{1},X_{2}$ from a continuous distribution, we have $X_{1}\neq X_{2}$ with probability $1$ and hence $\varphi_{2}(\mathbf{Z}_{1},\mathbf{Z}_{2})=0$. So the “second-order influence function” is always 00 when covariates have continuous components and cannot help us improve the estimator. On the other hand, when $X$ is discrete it is possible that $X_{1},X_{2}$ fall into the same category and $\varphi_{2}(\mathbf{Z}_{1},\mathbf{Z}_{2})\neq 0$. Then the idea is to use second-order estimator based on $\varphi_{1},\varphi_{2}$ to simultaneously correct for the first-order and second-order bias of plug-in-style estimator $\widehat{\eta}=\mathbb{P}_{n}[YA-\widehat{\mu}_{X}\widehat{\pi}_{X}]$. Consider the following second-order estimator of $\eta$ and $\rho$,
$\widehat{\eta}=\mathbb{P}_{n}[(A-\widehat{\pi}_{X})(Y-\widehat{\mu}_{X})]-\mathbb{U}_{n}\left[\frac{(A_{1}-\widehat{\pi}_{X_{1}})I(X_{1}=X_{2})(Y_{2}-\widehat{\mu}_{X_{2}})}{\widehat{p}_{X_{1}}}\right],$ (10)
$\widehat{\rho}=\mathbb{P}_{n}\left[(A-\widehat{\pi}_{X})^{2}\right]-\mathbb{U}_{n}\left[\frac{(A_{1}-\widehat{\pi}_{X_{1}})I(X_{1}=X_{2})(A_{2}-\widehat{\pi}_{X_{2}})}{\widehat{p}_{X_{1}}}\right],$ (11)
where $\mathbb{P}_{n},\mathbb{U}_{n}$ are empirical and U-statistic measures. The following theorem establishes the estimation guarantees of second-order estimators when the nuisance functions are estimated from a separate independent sample agnostically, i.e., we do not specify the way to estimate $\mu$ and $\pi$.
Theorem 5.
Suppose the nuisance functions $\pi_{k},\mu_{k}$ and covariate distribution $p_{k}$ are estimated from a separate independent sample $D$ as $\widehat{\pi}_{k},\widehat{\mu}_{k},\widehat{p}_{k}$ with $\widehat{\mu}_{k},\widehat{\pi}_{k}\in[0,1]$. Then for second-order estimators in (10)–(11) we have
$|\mathbb{E}\left[\widehat{\eta}\mid D\right]-\eta|\leq$ $\,\|\widehat{\mu}-\mu\|_{2}\|\widehat{\pi}-\pi\|_{2}\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|,$
$\operatorname{Var}\left(\widehat{\eta}\mid D\right)\lesssim$ $\,\frac{1}{n}\left(1+\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|\right)^{2}+\frac{d}{n^{2}}\left(1+\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|\right)^{2},$
$|\mathbb{E}\left[\widehat{\rho}\mid D\right]-\rho|\leq$ $\,\|\widehat{\pi}-\pi\|_{2}^{2}\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|,$
$\operatorname{Var}\left(\widehat{\rho}\mid D\right)\lesssim$ $\,\frac{1}{n}\left(1+\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|\right)^{2}+\frac{d}{n^{2}}\left(1+\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|\right)^{2},$
[…]
Theorem 5 shows the advantage of second-order estimators: after correcting for the first-order and second-order bias, the conditional bias of $\widehat{\eta}$ depends on the product of estimation error of $\mu,\pi,p$ and is “third-order small”. In the following theorem, we parameterize the estimation rate of covariate distribution to show the consistency of second-order estimators in the high-dimensional regime $n\lesssim d$ with different choices of nuisance estimators $\widehat{\pi},\widehat{\mu}$.
Theorem 6.
Suppose the nuisance functions $\pi_{k},\mu_{k}$ and covariate distribution $p_{k}$ are estimated from a separate independent sample $D$ as $\widehat{\pi}_{k},\widehat{\mu}_{k},\widehat{p}_{k}$. Further assume the estimated covariate distribution $\widehat{p}_{k}$ satisfies
$\max_{1\leq k\leq d}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|\leq\xi_{n}<1.$
If nuisance functions $\pi_{k},\mu_{k}$ are estimated from a sample of size $n$ using empirical averages, then the second-order estimators of $\eta$ and $\rho$ in (10)–(11) satisfy
$\mathbb{E}\left[(\widehat{\eta}-\eta)^{2}\right]\lesssim\xi_{n}^{2}\frac{d\wedge n}{n}+\frac{d}{n^{2}}+\frac{1}{n},$
$\mathbb{E}\left[(\widehat{\rho}-\rho)^{2}\right]\lesssim\xi_{n}^{2}\frac{d\wedge n}{n}+\frac{d}{n^{2}}+\frac{1}{n}.$
If we set the nuisance functions $\widehat{\pi}_{k},\widehat{\mu}_{k}$ as 00, then the second-order estimators of $\eta$ and $\rho$ satisfy
$\mathbb{E}\left[(\widehat{\eta}-\eta)^{2}\right]\lesssim\xi_{n}^{2}+\frac{d}{n^{2}}+\frac{1}{n},$
$\mathbb{E}\left[(\widehat{\rho}-\rho)^{2}\right]\lesssim\xi_{n}^{2}+\frac{d}{n^{2}}+\frac{1}{n}.$
where the constant behind $\lesssim$” is an absolute constant.
As a consequence, under positivity $\epsilon\leq\pi_{k}\leq 1-\epsilon$ and assume $\widehat{\rho}\geq\epsilon(1-\epsilon)$, the estimator $\widehat{\theta}=\widehat{\eta}/\widehat{\rho}$ satisfies
$\mathbb{E}\left[(\widehat{\theta}-\theta)^{2}\right]\lesssim$ $\,\mathbb{E}\left[(\widehat{\eta}-\eta)^{2}\right]+\mathbb{E}\left[(\widehat{\rho}-\rho)^{2}\right]$
$\lesssim$ $\,\xi_{n}^{2}+\frac{d}{n^{2}}+\frac{1}{n},$
where the constant behind $\lesssim$” depends on $\epsilon$.
We note that Theorem 6 implies a faster convergence rate than $d^{2}/n^{2}+1/n$ when $\xi_{n}$ is small, which is hard to achieve for estimators $\widehat{p}_{k}$’s constructed from a sample of size $n$ in the high-dimensional regime $n\lesssim d$. For example, in the uniform case $p_{k}=1/d,1\leq k\leq d$, one can use Chernoff bound to show that with probability $1-\delta$,
$\max_{1\leq k\leq d}|d\widehat{p}_{k}-1|\leq\sqrt{\frac{3d\log(2d/\delta)}{n}}.$
Thus we need a sample size of order $d\log d$ to guarantee small uniform estimation errors of $p_{k}$, which is not achievable in the case $n\lesssim d$. The improvements on the convergence rate of $\widehat{\eta}$ should come from prior knowledge/assumptions on covariate distribution. The results imply that consistency is still possible in the regime $n\lesssim d$ when such information on covariate distribution is available. One example is that the covariate distribution is known to statisticians, then we have $\widehat{p}_{k}=p_{k},\xi_{n}=0$ and the estimators are consistent as long as $d/n^{2}\rightarrow 0$. Another useful setting is semi-supervised causal inference (Chakrabortty et al.,, 2022; Zhang et al.,, 2023; Kallus and Mao,, 2020; Zeng et al.,, 2024): there is a large number ($\gg n$) of individuals in the database who were not selected into the randomized trial/observational study, hence treatment and outcome are not available for them but one can use the database to estimate the covariate distribution very accurately thanks to the large sample size. Then $\xi_{n}$ is small and the estimation error of $\widehat{\eta}$ will be small according to Theorem 6.
When the covariate distribution can be estimated very well and $\xi_{n}$ is small, we can use arbitrary (bounded) nuisance estimators and still maintain consistency. As shown in Theorem 6, simply setting $\widehat{\pi},\widehat{\mu}$ as 00 leads to consistent estimation as long as $\xi_{n}$ vanishes and $d/n^{2}\rightarrow 0$. Using empirical average estimators has advantage over setting $\widehat{\pi},\widehat{\mu}$ as 00 only when $d\ll n$. The interpretation is that when $\xi_{n}$ is small, the conditional bias in Theorem 5 is negligible regardless of the convergence rate of $\widehat{\pi},\widehat{\mu}$. The conditional variance, whose order is independent of the nuisance estimation rate, is the dominant term in the MSE.
We conclude this section with additional discussion on second-order estimation of $\psi_{1}=\mathbb{E}[\mathbb{E}(Y\mid X,A=1)]$ in the discrete covariate setting. The forms of first-order and second-order influence functions are
$\phi_{1}(\mathbf{Z})=\frac{A(Y-\mu_{1X})}{\pi_{X}}+\mu_{1X},$
$\phi_{2}(\mathbf{Z}_{1},\mathbf{Z}_{2})=-\left(\frac{A_{1}}{\pi_{X_{1}}}-1\right)\frac{I(X_{1}=X_{2})}{p_{X_{1}}\pi_{X_{2}}}A_{2}(Y_{2}-\mu_{1X_{2}}).$
The second-order estimator is then
$\mathbb{P}_{n}\left[\widehat{\phi}_{1}(\mathbf{Z})\right]+\mathbb{U}_{n}\left[\widehat{\phi}_{2}(\mathbf{Z}_{1},\mathbf{Z}_{2})\right].$ (12)
Assume we use a separate independent sample $D$ to estimate nuisance functions $\pi,\mu_{1}$ and covariate distribution $p$, the conditional bias of second-order estimator (12) is
$\mathbb{E}\left[(\widehat{\mu}_{1X}-\mu_{1X})\left(1-\frac{\pi_{X}}{\widehat{\pi}_{X}}\right)\left(1-\frac{\pi_{X}p_{X}}{\widehat{\pi}_{X}\widehat{p}_{X}}\right)\mid D\right].$
In order to obtain a faster rate by making use of information on covariate distribution, one also needs an accurate estimator of propensity score $\pi$ to ensure
$\max_{1\leq k\leq d}\bigg{|}1-\frac{\pi_{k}p_{k}}{\widehat{\pi}_{k}\widehat{p}_{k}}\bigg{|}$ (13)
is small. In observational studies where the treatment process is unknown, it is hard to estimate the product of propensity score and covariate distribution well and make (13) small in the regime $n\lesssim d$, due to the same reason discussed after Theorem 6. This further demonstrates the difficulty in ATE estimation with solely information on covariate distribution. In applications, WATE also serves as an important and useful summary of treatment effects to estimate and report.
6 Simulation Study
In this section, we provide numerical results to verify the theoretical properties of the estimators discussed in Section 35. Consider the following data generating process: $X$ has a categorical distribution with uniform probability ${\boldsymbol{p}}=(1/d,\dots,1/d)^{\top}\in\mathbb{R}^{d}$. For each category $X=k$, the propensity score $\pi_{k}=1/2$ and the conditional means of outcome within treated and untreated groups are $\mu_{1k}=1/2,\mu_{0k}=1/4$, respectively. Hence the treatment effects are homogeneous across different levels (i.e. $\tau_{k}=1/4$ for all $k\in[d]$) in our simulation setting and expected conditional covariance between treatment and outcome $\eta=1/16$ (considered in Section 5). For each estimator we generate data containing $n$ observations with $n\in\{100,500,1000,5000,10000\}$, apply the proposed method, and compute the estimated RMSE. Such procedure is repeated $M=500$ times and the average RMSE is reported in each plot.
We first consider the plug-in estimator discussed in Section 3. For each sample size $n$ let the number of categories $d=\lfloor n^{\gamma}\rfloor$ with $\gamma\in\{0.5,0.55,0.6,\dots,1.5\}$. The idea is to evaluate how large the order of $d$ can be to maintain consistency (He et al.,, 2021). The average treatment effect is $\psi=1/4$ and we estimate RMSE as
$\widehat{\text{RMSE}}=\left[\frac{1}{M}\sum_{m=1}^{M}(\widehat{\psi}^{m}-\psi)^{2}\right]^{1/2},$
where $\widehat{\psi}^{m}$ is the estimated ATE from $m$-th repetition. The relationship between RMSE and $\gamma$ is summarized in Figure 1. For a fixed sample size $n$, as $\gamma$ increases (i.e. $d$ increases) the estimated RMSE also increases as expected. We see a clear phase transition in the plot: in the region $\gamma<0.8$, the RMSE is quite stable of order $1/\sqrt{n}$; around $\gamma=1$ the RMSE increases drastically, indicating the plug-in estimator starts to have larger error. This corresponds to our theoretical results in Section 3: the plug-in estimator is consistent if and only if $d/n\rightarrow 0$.
Figure 1: RMSE V.S. $\gamma$ for the plug-in estimator, where $\gamma$ controls the order of $d$.
Then we consider the estimator (9) proposed in Section 4 under effect homogeneity. Again for each fixed sample size $n$ we let $d=\lfloor n^{\gamma}\rfloor$ with $\gamma\in\{0.5,0.55,0.6,\dots,2\}$. Note that the treatment effects are homogeneous in our setting (i.e. $\sigma_{n}=0$) and we expect $\widehat{\tau}$ in equation (9) to be consistent in a wider regime.
The relationship between RMSE and $\gamma$ is summarized in Figure 2. The phase transition occurs in the region $\gamma>1$ (instead of at $\gamma=1$) and the estimator in (9) has a smaller error in the region $1.251<\gamma<$ 1.25 compared with the plug-in estimator when $n\geq 1000$. For instance, in the case $n=1000,\gamma=1$ the plug-in estimator has RMSE around $0.08$ while the estimator under effect homogeneity has RMSE around $0.05$. This coincides with our theoretical results in Section 4 that the estimator in equation (9) has a faster rate and is consistent in a wider regime.
Figure 2: RMSE V.S. $\gamma$ for the estimator under effect homogeneity, where $\gamma$ controls the order of $d$.
To further understand the order of the RMSE, for $n\in\{1000,10000\}$ we include the theoretical upper bound on RMSE, which is of order $C_{1}\sqrt{d/n^{2}}=C_{1}n^{\gamma/2-1}$, in the plot as a benchmark. Here the constant $C$ is chosen as $1.5$ to fit the empirical RMSE curve. The results are summarized in Figure 3. When $\gamma<1.5$, the empirical RMSE fits the theoretical upper bound quite well. When $\gamma>1.5$ the empirical RMSE starts to deviate from the theoretical bound. In our experiments we found when $\gamma>1.5$ and $d$ is large, the denominator in the estimator (9) is usually small (i.e., only a few categories have both treated and untreated samples) and the variation is large, which may explain the deviation of the empirical RMSE from the theoretical bound.
(a) n=1000
(b) n=10000
Figure 3: Comparison of theoretical and empirical order of RMSE
Finally, we evaluate the performance of the second-order estimator in (10) of expected conditional covariance $\eta$ (the estimator (11) of expected conditional variance of treatment is expected to have similar performance). In our setting $\eta=1/16$ and for each fixed sample size $n$, let $d=\lfloor n^{\gamma}\rfloor$ with $\gamma\in\{0.5,0.55,0.6,\dots,2\}$. We set the estimates of probabilities $\widehat{p}_{k}$’s as the true $p_{k}=1/d$ and hence $\xi_{n}=0$ in Theorem 6. The estimates $\widehat{\pi}_{k},\widehat{\mu}_{k}$’s are all set to 00. The results are summarized in Figure 4. Similar to Figure 2, the estimated RMSE is quite stable in the region $\gamma<1.25$ and the phase transition seems to happen in the region $\gamma>1$.
Figure 4: RMSE V.S. $\gamma$ for the second-order estimator (10) using true covariate distribution ${\boldsymbol{p}}$, where $\gamma$ controls the order of $d$.
We also plot the relationship between empirical RMSE and theoretical bound $C_{2}\sqrt{d/n^{2}}=C_{2}n^{\gamma/2-1}$ with $C_{2}=0.5$ in Figure 5. We see the empirical RMSE matches the theoretical bound very well.
(a) n=1000
(b) n=10000
Figure 5: Comparison of theoretical and empirical order of RMSE
7 Discussion
In this paper, we studied the treatment effects estimation problem in the context of high-dimensional discrete covariates. Theoretical properties of commonly used regression, weighting and doubly robust estimators are examined in this non-classic regime. We also evaluated the role of effect homogeneity and covariate distribution in treatment effects estimation and proposed estimators that can properly take advantage of these structures and achieve faster convergence rates. Finally, we explored the fundamental limits of treatment effects estimation and showed consistent estimation of ATE is a difficult task on high-dimensional data. The discrete covariate setting is not only an interesting base case but also informative for the general dataset with continuous components. We hope our work can help researchers appropriately understand and interpret the treatment effects estimated from datasets with many covariates.
There are several possible extensions for future work. In this paper, we borrowed the moment matching tools from the theoretical computer science literature (Jiao et al.,, 2015; Wu and Yang,, 2016, 2019) to study the minimax lower bound on ATE estimation. It would be interesting to explore how to construct estimators of ATE based on polynomial approximation theory and realize the “effective sample size enlargement” phenomenon, which is feasible in entropy and support size estimation. Moreover, the minimax lower bound for ATE we proved is an initial result that does not exactly match the upper bound. More effort is needed to come up with new construction and tighten the lower bound if possible. Examining how other structural assumptions can help us achieve faster estimation rates is also an interesting topic. Other extensions could involve the estimation of different causal estimands in a similar discrete setting, including time-varying treatment effects, generalizability and transportability, optimal treatment regimes, instrumental variable and more. Developing a comprehensive understanding on the properties of popular estimators and fundamental limits of different causal functionals in the discrete setting could be an interesting avenue left for future investigation.
Appendix A Asymptotic Normality of Estimators
In this section, we provide a central limit theorem for the plug-in estimator when the number of categories $d$ is fixed as a constant, summarized in the following theorem.
Theorem 7.
Supposed $X\in[d]$ is discrete with $d$ fixed and the nuisance estimators are the empirical averages defined in (2). Then we have
$\sqrt{n}\left(\widehat{\psi}-\psi\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(0,\operatorname{Var}(\varphi(\mathbf{Z}))\right),$
where
$\varphi(\mathbf{Z})=\mu_{1X}-\mu_{0X}+\left(\frac{A}{\pi_{X}}-\frac{1-A}{1-\pi_{X}}\right)\left(Y-\mu_{AX}\right)$
is the first-order influence function of ATE $\psi$ under a nonparametric model.
Theorem 7 implies that when the covariate is discrete with fixed dimension $d$, the plug-in style estimator $\widehat{\psi}$ is $\sqrt{n}$-consistent and asymptotically normal. Hence in low-dimensional problems, the plug-in estimator $\widehat{\psi}$ enjoys appealing properties and we can construct confidence intervals and perform statistical tests on ATE based on Theorem 7. It is worth noting that asymptotic normality also holds in the regime $d=o(\sqrt{n})$. However, truncating the propensity score estimates at $\epsilon$ and $1-\epsilon$ is required to avoid instability induced by imprecise estimation of $\pi_{k}$ in the high-dimensional regime. Moreover, the empirical influence function $\widehat{\varphi}$ belongs to a Donsker class when $d$ is fixed since it could be expressed as finite-dimensional parametric models (the dimension depends on $d$). As $d=d_{n}$ grows with $n$, they may not belong to a Donsker class and sample splitting is required to control the empirical process term.
Appendix B Proof of Main Results
B.1 Proof of Proposition 1
Proof.
We will prove three estimators for $\psi_{1}=\mathbb{E}[Y^{1}]=\sum_{k=1}^{d}{p}_{k}{\mu}_{1k}$ are equal and similar results hold for $\mathbb{E}[Y^{0}]=\sum_{k=1}^{d}{p}_{k}{\mu}_{0k}$. First consider the regression estimator
$\widehat{\psi}_{1,\text{reg}}=$ $\,\frac{1}{n}\sum_{i=1}^{n}\widehat{\mu}_{1X_{i}}$
$=$ $\,\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}I(X_{i}=k)\widehat{\mu}_{1X_{i}}$
$=$ $\,\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}I(X_{i}=k)\widehat{\mu}_{1k}$
$=$ $\,\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}I(X_{i}=k)\right)\widehat{\mu}_{1k}$
$=$ $\,\sum_{k=1}^{d}\widehat{p}_{k}\widehat{\mu}_{1k}=\widehat{\psi},$
where the second equation follows from the fact $\sum_{k=1}^{d}I(X_{i}=k)=1$. In fact, given $n$ samples, we have $\sum_{k=1}^{d}I(X_{i}=k)=\sum_{k:\widehat{p}_{k}>0}I(X_{i}=k)=1$. With this in mind, for the inverse probability weighting estimator we have
$\widehat{\psi}_{1,\text{ipw}}=$ $\,\frac{1}{n}\sum_{i=1}^{n}\frac{A_{i}Y_{i}}{\widehat{\pi}_{X_{i}}}$
$=$ $\,\frac{1}{n}\sum_{i=1}^{n}\sum_{k:\widehat{p}_{k}>0}I(X_{i}=k)\frac{A_{i}Y_{i}}{\widehat{\pi}_{X_{i}}}$
$=$ $\,\frac{1}{n}\sum_{i=1}^{n}\sum_{k:\widehat{p}_{k}>0}I(X_{i}=k)\frac{A_{i}Y_{i}}{\widehat{\pi}_{k}}$
$=$ $\,\sum_{k:\widehat{p}_{k}>0}\frac{1}{n}\sum_{i=1}^{n}\frac{I(X_{i}=k)A_{i}Y_{i}}{\widehat{\pi}_{k}}$
$=$ $\,\sum_{k:\widehat{p}_{k}>0}\widehat{p}_{k}\frac{1}{n}\sum_{i=1}^{n}\frac{I(X_{i}=k)A_{i}Y_{i}}{\widehat{p}_{k}\widehat{\pi}_{k}}$
$=$ $\,\sum_{k:\widehat{p}_{k}>0}\widehat{p}_{k}\frac{\widehat{q}_{1k}}{\widehat{w}_{k}}$
$=$ $\,\sum_{k:\widehat{p}_{k}>0}\widehat{p}_{k}\widehat{\mu}_{1k}$
$=$ $\,\sum_{k=1}^{d}\widehat{p}_{k}\widehat{\mu}_{1k}=\widehat{\psi},$
where $\widehat{q}_{1k}=\frac{1}{n}\sum_{i=1}^{n}I(X_{i}=k)A_{i}Y_{i}$ and $\widehat{q}_{1k}/\widehat{w}_{k}=\widehat{\mu}_{1k}$ by definition in (2). Note that all the equations still hold when some categories have no treated samples (i.e. $\widehat{w}_{k}=0$) since we define $/0=0$ 0 whenever it appears. The last equation holds since the categories with $\widehat{p}_{k}=0$ do not contribute to the estimation. Finally, we consider the doubly robust estimator. Note that
$\,\mathbb{P}_{n}\left[\frac{A\widehat{\mu}_{1X}}{\widehat{\pi}_{X}}\right]$
$=$ $\,\frac{1}{n}\sum_{i=1}^{n}\sum_{k:\widehat{p}_{k}>0}I(X_{i}=k)\frac{A_{i}\widehat{\mu}_{1X_{i}}}{\widehat{\pi}_{X_{i}}}$
$=$ $\,\sum_{k:\widehat{p}_{k}>0}\left(\frac{\sum_{i=1}^{n}A_{i}I(X_{i}=k)}{n\widehat{p}_{k}\widehat{\pi}_{k}}\right)\widehat{p}_{k}\widehat{\mu}_{1k}.$
Note that $n\widehat{p}_{k}\widehat{\pi}_{k}=\sum_{i=1}^{n}A_{i}I(X_{i}=k)$. If $n\widehat{p}_{k}\widehat{\pi}_{k}>0$ then
$\frac{\sum_{i=1}^{n}A_{i}I(X_{i}=k)}{n\widehat{p}_{k}\widehat{\pi}_{k}}=1.$
If $n\widehat{p}_{k}\widehat{\pi}_{k}>0$ by our definition on $/0=0$ 0
$\frac{\sum_{i=1}^{n}A_{i}I(X_{i}=k)}{n\widehat{p}_{k}\widehat{\pi}_{k}}=0,\,\widehat{\mu}_{1k}=\frac{\widehat{q}_{1k}}{\widehat{w}_{k}}=0$
and hence we always have
$\left(\frac{\sum_{i=1}^{n}A_{i}I(X_{i}=k)}{n\widehat{p}_{k}\widehat{\pi}_{k}}\right)\widehat{p}_{k}\widehat{\mu}_{1k}=\widehat{p}_{k}\widehat{\mu}_{1k}.$
We conclude
$\mathbb{P}_{n}\left[\frac{A\widehat{\mu}_{1X}}{\widehat{\pi}_{X}}\right]=\sum_{k:\widehat{p}_{k}>0}\widehat{p}_{k}\widehat{\mu}_{1k}=\sum_{k=1}^{d}\widehat{p}_{k}\widehat{\mu}_{1k}.$
This together with
$\mathbb{P}_{n}\left[\frac{AY}{\widehat{\pi}_{X}}\right]=\widehat{\psi}$
shows
$\widehat{\psi}_{1,\text{dr}}=\mathbb{P}_{n}\left[\frac{A(Y-\widehat{\mu}_{1X})}{\widehat{\pi}_{X}}+\widehat{\mu}_{1X}\right]=\widehat{\psi}.$
B.2 Proof of Proposition 2
Proof.
We focus on the bias of $\widehat{\psi}_{1}$ and $\widehat{\psi}_{0}$ can be similarly analyzed. We may rewrite the plug-in estimator as
$\widehat{\psi}_{1}=\sum_{k=1}^{d}\frac{\widehat{q}_{1k}\widehat{p}_{k}}{\widehat{w}_{k}}I(\widehat{w}_{k}>0)$
to emphasize the definition $/0=0$ 0 in each term. On the event $\{n\widehat{w}_{k}=0\}$, we automatically have $\widehat{q}_{1k}=0$. On the event $\{n\widehat{w}_{k}>0\}$,
$\mathbb{E}[n\widehat{q}_{1k}\mid\mathbf{X}^{n},\mathbf{A}^{n}]=n\widehat{w}_{k}\mu_{1k}=n\frac{q_{1k}}{w_{k}}\widehat{w}_{k}$
since $n\widehat{q}_{1k}\mid\mathbf{X}^{n},\mathbf{A}^{n}\sim\text{B}(n\widehat{w}_{k},\mu_{1k})$. Hence we have
$\mathbb{E}\left[\frac{\widehat{q}_{1k}\widehat{p}_{k}}{\widehat{w}_{k}}I(\widehat{w}_{k}>0)\right]=\mathbb{E}\left[\frac{\widehat{p}_{k}}{\widehat{w}_{k}}I(\widehat{w}_{k}>0)\mathbb{E}(\widehat{q}_{1k}\mid\mathbf{X}^{n},\mathbf{A}^{n})\right]=\mathbb{E}[\widehat{p}_{k}I(\widehat{w}_{k}>0)]\frac{q_{1k}}{w_{k}}.$
The bias of each individual term is
$\,\mathbb{E}\left[\frac{\widehat{q}_{1k}\widehat{p}_{k}}{\widehat{w}_{k}}I(\widehat{w}_{k}>0)\right]-\frac{q_{1k}p_{k}}{w_{k}}$
$=$ $\,\mathbb{E}[\widehat{p}_{k}I(\widehat{w}_{k}>0)]\frac{q_{1k}}{w_{k}}-\frac{q_{1k}p_{k}}{w_{k}}$
$=$ $\,-\mathbb{E}[\widehat{p}_{k}I(\widehat{w}_{k}=0)]\frac{q_{1k}}{w_{k}}$
Further note that $n\widehat{w}_{k}\mid\mathbf{X}^{n}\sim\text{B}(n\widehat{p}_{k},\pi_{k})$,
$\,\mathbb{E}[\widehat{p}_{k}I(\widehat{w}_{k}=0)]$
$=$ $\,\mathbb{E}[\widehat{p}_{k}I(n\widehat{w}_{k}=0)]$
$=$ $\,\mathbb{E}[\widehat{p}_{k}\mathbb{P}(n\widehat{w}_{k}=0\mid\mathbf{X}^{n})]$
$=$ $\,\mathbb{E}[\widehat{p}_{k}(1-\pi_{k})^{n\widehat{p}_{k}}].$
We then use the fact $\mathbb{E}[Vc^{V}]=npc(1-p+cp)^{n-1}$ for $V\sim\text{Bin}(n,p)$ and obtain
$\,\mathbb{E}[\widehat{p}_{k}(1-\pi_{k})^{n\widehat{p}_{k}}]$
$=$ $\,\frac{1}{n}\mathbb{E}[n\widehat{p}_{k}(1-\pi_{k})^{n\widehat{p}_{k}}]$
$=$ $\,\frac{1}{n}\times np_{k}(1-\pi_{k})[1-p_{k}+(1-\pi_{k})p_{k}]^{n-1}$
$=$ $\,p_{k}(1-\pi_{k})(1-p_{k}\pi_{k})^{n-1}$
Hence the bias for individual terms is
$\mathbb{E}\left[\frac{\widehat{q}_{1k}\widehat{p}_{k}}{\widehat{w}_{k}}I(\widehat{w}_{k}>0)\right]-\frac{q_{1k}p_{k}}{w_{k}}=-\mu_{1k}p_{k}(1-\pi_{k})(1-p_{k}\pi_{k})^{n-1}$
and we conclude
$\mathbb{E}[\widehat{\psi}_{1}-\psi_{1}]=-\sum_{k=1}^{d}\mu_{1k}p_{k}(1-\pi_{k})(1-p_{k}\pi_{k})^{n-1}.$
Similarly (or by symmetry) for $\widehat{\psi}_{0}$ we have
$\mathbb{E}[\widehat{\psi}_{0}-\psi_{0}]=-\sum_{k=1}^{d}\mu_{0k}p_{k}\pi_{k}(1-p_{k}+p_{k}\pi_{k})^{n-1}$
B.3 Proof of Proposition 3
Proof.
$\sup_{\mathbb{P}\in\mathcal{D}_{+}(\epsilon)}|\mathbb{E}_{\mathbb{P}}[\widehat{\psi}_{1}-\psi_{1}]|=$ $\,\sup_{\mu_{1k},\pi_{k},p_{k}}\sum_{k=1}^{d}\mu_{1k}p_{k}(1-\pi_{k})(1-p_{k}\pi_{k})^{n-1}$
$=$ $\,\sup_{\pi_{k},p_{k}}\sum_{k=1}^{d}p_{k}(1-\pi_{k})(1-p_{k}\pi_{k})^{n-1}$
$=$ $\,\sup_{p_{k}}\sum_{k=1}^{d}p_{k}(1-\epsilon)(1-\epsilon p_{k})^{n-1}$
The first equation follows from $\mu_{1k}\leq 1$ and the second follows from $\pi_{k}\geq\epsilon$. For the upper bound consider the function $f_{1}(x)=x(1-\epsilon x)^{n-1}$ and $f_{1}^{\prime}(x)=(1-\epsilon nx)(1-\epsilon x)^{n-2}$. Hence for $x\in[0,1]$ we always have
$f_{1}(x)\leq f_{1}(1/\epsilon n)=\frac{1}{\epsilon n}\left(1-\frac{1}{n}\right)^{n-1}<\frac{1}{\epsilon n}.$
We thus conclude
$\sup_{p_{k}}\sum_{k=1}^{d}p_{k}(1-\epsilon)(1-\epsilon p_{k})^{n-1}\leq\frac{1-\epsilon}{\epsilon}\frac{d}{n}.$
For the lower bound we plug in the uniform distribution on $[d]$ and obtain
$\sup_{p_{k}}\sum_{k=1}^{d}p_{k}(1-\epsilon)(1-\epsilon p_{k})^{n-1}\geq(1-\epsilon)\left(1-\frac{\epsilon}{d}\right)^{n-1}.$
Now by inequality
$\log(1+x)\geq\frac{x}{1+x},\,x\geq-1$
we have
$(n-1)\log\left(1-\frac{\epsilon}{d}\right)\geq-\frac{\epsilon(n-1)}{d-\epsilon}$
and hence
$(1-\epsilon)\left(1-\frac{\epsilon}{d}\right)^{n-1}\geq(1-\epsilon)\exp\left[-\frac{\epsilon(n-1)}{d-\epsilon}\right].$
For the lower bound when $n\geq 1/\epsilon+1$ consider two cases:
Case 1: $\frac{d-1}{\epsilon(n-1)}\leq 1$, one could set $p_{1}=\dots=p_{d-1}=\frac{1}{\epsilon(n-1)},p_{d}=1-\frac{d-1}{\epsilon(n-1)}$. We have
$\sum_{k=1}^{d}p_{k}(1-\epsilon p_{k})^{n-1}\geq\frac{d-1}{\epsilon(n-1)}\left(1-\frac{1}{n-1}\right)^{n-1}\geq\frac{1}{4\epsilon}\frac{d-1}{n-1},$
since $\left(1-\frac{1}{n-1}\right)^{n-1}\geq\frac{1}{4}$ when $n\geq 3$.
Case 2: $\frac{d-1}{\epsilon(n-1)}>1$, we have $d\geq\lfloor\epsilon(n-1)\rfloor+1$ and we set
$p_{1}=\dots=p_{\lfloor\epsilon(n-1)\rfloor}=\frac{1}{\epsilon(n-1)},p_{\lfloor\epsilon(n-1)\rfloor+1}=1-\frac{\lfloor\epsilon(n-1)\rfloor}{\epsilon(n-1)}.$
$\sum_{k=1}^{d}p_{k}(1-\epsilon p_{k})^{n-1}\geq\frac{\lfloor\epsilon(n-1)\rfloor}{\epsilon(n-1)}\left(1-\frac{1}{n-1}\right)^{n-1}\geq\frac{1}{8},$
since $\lfloor x\rfloor/x\geq 1/2$ when $x\geq 1$. Combining two cases we have
$\sum_{k=1}^{d}p_{k}(1-\epsilon p_{k})^{n-1}\geq\frac{1}{8}\left(\frac{d-1}{\epsilon(n-1)}\wedge 1\right)$
when $n\geq 1/\epsilon+1$.
B.4 Proof of Theorem 1
Proof.
$\operatorname{Var}(\widehat{\psi}_{1})=\mathbb{E}[\operatorname{Var}(\widehat{\psi}_{1}\mid\mathbf{X}^{n},\mathbf{A}^{n})]+\operatorname{Var}(\mathbb{E}[\widehat{\psi}_{1}\mid\mathbf{X}^{n},\mathbf{A}^{n}]).$ (14)
We start with the first term. By the conditional independence of $\widehat{q}_{11},\dots,\widehat{q}_{1d}$ given $\mathbf{X}^{n},\mathbf{A}^{n}$, we have
$\,\operatorname{Var}(\widehat{\psi}_{1}\mid\mathbf{X}^{n},\mathbf{A}^{n})$
$=$ $\,\operatorname{Var}\left(\sum_{k=1}^{d}\frac{\widehat{q}_{1k}\widehat{p}_{k}I(\widehat{w}_{k}>0)}{\widehat{w}_{k}}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)$
$=$ $\,\sum_{k=1}^{d}\frac{\widehat{p}_{k}^{2}I(\widehat{w}_{k}>0)}{\widehat{w}_{k}^{2}}\operatorname{Var}(\widehat{q}_{1k}\mid\mathbf{X}^{n},\mathbf{A}^{n})$
$=$ $\,\sum_{k=1}^{d}\frac{\widehat{p}_{k}^{2}I(\widehat{w}_{k}>0)}{\widehat{w}_{k}^{2}}\frac{\widehat{w}_{k}\mu_{1k}(1-\mu_{1k})}{n}$
$=$ $\,\frac{1}{n}\sum_{k=1}^{d}\frac{\widehat{p}_{k}^{2}I(\widehat{w}_{k}>0)}{\widehat{w}_{k}}\mu_{1k}(1-\mu_{1k})$
where we use the fact $n\widehat{q}_{1k}\mid\mathbf{X}^{n},\mathbf{A}^{n}\sim\text{Bin}(n\widehat{w}_{k},\mu_{1k})$.
Note that $n\widehat{w}_{k}\sim\text{Bin}(n\widehat{p}_{k},\pi_{k})\mid\mathbf{X}^{n}$ and apply Lemma 7, we have
$\,\mathbb{E}[\operatorname{Var}(\widehat{\psi}_{1}\mid\mathbf{X}^{n},\mathbf{A}^{n})]$
$=$ $\,\frac{1}{n}\sum_{k=1}^{d}\mathbb{E}\left[\frac{\widehat{p}_{k}^{2}I(\widehat{w}_{k}>0)}{\widehat{w}_{k}}\right]\mu_{1k}(1-\mu_{1k})$
$=$ $\,\sum_{k=1}^{d}\mathbb{E}\left\{\widehat{p}_{k}^{2}\mathbb{E}\left[\frac{I(\widehat{w}_{k}>0)}{n\widehat{w}_{k}}\mid\mathbf{X}^{n}\right]\right\}\mu_{1k}(1-\mu_{1k})$
$\leq$ $\,\sum_{k=1}^{d}\mathbb{E}\left\{\widehat{p}_{k}^{2}\frac{2}{(n\widehat{p}_{k}+1)\pi_{k}}\right\}\mu_{1k}(1-\mu_{1k})$
$\leq$ $\,\frac{2}{n}\sum_{k=1}^{d}\mathbb{E}\left\{\frac{\widehat{p}_{k}}{\pi_{k}}\right\}\mu_{1k}(1-\mu_{1k})$
$\leq$ $\,\frac{1}{2n\epsilon}\sum_{k=1}^{d}\mathbb{E}[\widehat{p}_{k}]$
$=$ $\,\frac{1}{2n\epsilon},$
where we use the fact $\mu_{1k}(1-\mu_{1k})\leq 1/4$. Then we evaluate the second term in (14)
$\mathbb{E}[\widehat{\psi}_{1}\mid\mathbf{X}^{n},\mathbf{A}^{n}]=\sum_{k=1}^{d}\mathbb{E}\left[\frac{\widehat{q}_{1k}\widehat{p}_{k}}{\widehat{w}_{k}}I(\widehat{w}_{k}>0)\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]=\sum_{k=1}^{d}\widehat{p}_{k}I(\widehat{w}_{k}>0)\mu_{1k}.$
By further using the property of conditional variance we have
$\operatorname{Var}(\mathbb{E}[\widehat{\psi}_{1}\mid\mathbf{X}^{n},\mathbf{A}^{n}])=\mathbb{E}\left[\operatorname{Var}\left(\sum_{k=1}^{d}\widehat{p}_{k}I(\widehat{w}_{k}>0)\mu_{1k}\mid\mathbf{X}^{n}\right)\right]+\operatorname{Var}\left(\mathbb{E}\left[\sum_{k=1}^{d}\widehat{p}_{k}I(\widehat{w}_{k}>0)\mu_{1k}\mid\mathbf{X}^{n}\right]\right)$ (15)
We analyze the first term in (15). By the conditional independence of $n\widehat{w}_{1},\dots,n\widehat{w}_{d}$ given $\mathbf{X}^{n}$ we have
$\,\operatorname{Var}\left(\sum_{k=1}^{d}\widehat{p}_{k}I(\widehat{w}_{k}>0)\mu_{1k}\mid\mathbf{X}^{n}\right)$
$=$ $\,\sum_{k=1}^{d}\widehat{p}_{k}^{2}\mu_{1k}^{2}\operatorname{Var}(I(\widehat{w}_{k}>0)\mid\mathbf{X}^{n})$
$\leq$ $\,\sum_{k=1}^{d}\widehat{p}_{k}^{2}\mu_{1k}^{2}(1-\pi_{k})^{n\widehat{p}_{k}}$
$\leq$ $\,\sum_{k=1}^{d}\widehat{p}_{k}^{2}\mu_{1k}^{2}\frac{1}{\pi_{k}(1+n\widehat{p}_{k})}$
$\leq$ $\,\frac{1}{\epsilon n}\sum_{k=1}^{d}\widehat{p}_{k}$
$=$ $\,\frac{1}{\epsilon n},$
where we use the fact
[…]
Hence we have
$\mathbb{E}\left[\operatorname{Var}\left(\sum_{k=1}^{d}\widehat{p}_{k}I(\widehat{w}_{k}>0)\mu_{1k}\mid\mathbf{X}^{n}\right)\right]\leq\frac{1}{\epsilon n}.$
For the second term in (15), we have
$\,\mathbb{E}\left[\sum_{k=1}^{d}\widehat{p}_{k}I(\widehat{w}_{k}>0)\mu_{1k}\mid\mathbf{X}^{n}\right]$
$=$ $\,\sum_{k=1}^{d}\widehat{p}_{k}\mu_{1k}(1-\mathbb{P}(n\widehat{w}_{k}=0\mid\mathbf{X}^{n}))$
$=$ $\,\sum_{k=1}^{d}\mu_{1k}\widehat{p}_{k}[1-(1-\pi_{k})^{n\widehat{p}_{k}}]$
So we need to evaluate
$\,\operatorname{Var}\left(\mathbb{E}\left[\sum_{k=1}^{d}\widehat{p}_{k}I(\widehat{w}_{k}>0)\mu_{1k}\mid\mathbf{X}^{n}\right]\right)$
$=$ $\,\operatorname{Var}\left(\sum_{k=1}^{d}\mu_{1k}\widehat{p}_{k}[1-(1-\pi_{k})^{n\widehat{p}_{k}}]\right)$
$=$ $\,\sum_{k=1}^{d}\mu_{1k}^{2}\operatorname{Var}\left(\widehat{p}_{k}[1-(1-\pi_{k})^{n\widehat{p}_{k}}]\right)+\sum_{i\neq j}\mu_{1i}\mu_{1j}\operatorname{Cov}\left(\widehat{p}_{i}[1-(1-\pi_{i})^{n\widehat{p}_{i}}],\widehat{p}_{j}[1-(1-\pi_{j})^{n\widehat{p}_{j}}]\right).$
We note that
$\mathbb{E}[\widehat{p}_{k}[1-(1-\pi_{k})^{n\widehat{p}_{k}}]]=p_{k}-p_{k}(1-\pi_{k})(1-\pi_{k}p_{k})^{n-1}$
and
$\mathbb{E}\left\{\widehat{p}_{k}^{2}[1-(1-\pi_{k})^{n\widehat{p}_{k}}]^{2}\right\}\leq\mathbb{E}[\widehat{p}_{k}^{2}]=\frac{p_{k}(1-p_{k})}{n}+p_{k}^{2}.$
This yields the following bound
$\,\operatorname{Var}(\widehat{p}_{k}[1-(1-\pi_{k})^{n\widehat{p}_{k}}])$
$\leq$ $\,\frac{p_{k}(1-p_{k})}{n}+2p_{k}^{2}(1-\pi_{k})(1-\pi_{k}p_{k})^{n-1}$
$\leq$ $\,\frac{p_{k}(1-p_{k})}{n}+2p_{k}^{2}(1-\pi_{k})\frac{1}{n\pi_{k}p_{k}}$
$\leq$ $\,\frac{p_{k}(1-p_{k})}{n}+\frac{2p_{k}}{n\epsilon}$
$\leq$ $\,\left(1+\frac{2}{\epsilon}\right)\frac{p_{k}}{n}.$
So we have
$\sum_{k=1}^{d}\mu_{1k}^{2}\operatorname{Var}\left(\widehat{p}_{k}[1-(1-\pi_{k})^{n\widehat{p}_{k}}]\right)\leq\left(1+\frac{2}{\epsilon}\right)\frac{1}{n}.$
For the covariance part, the computations are more involved. We need to compute
$\,\mathbb{E}\left\{\widehat{p}_{i}[1-(1-\pi_{i})^{n\widehat{p}_{i}}]\widehat{p}_{j}[1-(1-\pi_{j})^{n\widehat{p}_{j}}]\right\}$
$=$ $\,\mathbb{E}[\widehat{p}_{i}\widehat{p}_{j}]-\mathbb{E}[\widehat{p}_{i}\widehat{p}_{j}(1-\pi_{i})^{n\widehat{p}_{i}}]-\mathbb{E}[\widehat{p}_{i}\widehat{p}_{j}(1-\pi_{j})^{n\widehat{p}_{j}}]+\mathbb{E}[\widehat{p}_{i}\widehat{p}_{j}(1-\pi_{i})^{n\widehat{p}_{i}}(1-\pi_{j})^{n\widehat{p}_{j}}]$
For three-dimensional multinomial distribution $(X_{1},X_{2},X_{3})\sim$ Multinomial$(n,p_{1},p_{2},p_{3})$ the probability generating function is
$\mathbb{E}[z_{1}^{X_{1}}z_{2}^{X_{2}}z_{3}^{X_{3}}]=(p_{1}z_{1}+p_{2}z_{2}+p_{3}z_{3})^{n}.$
From this formula and differentiation we have
$\mathbb{E}[X_{1}X_{2}z_{1}^{X_{1}}z_{2}^{X_{2}}]=n(n-1)p_{1}z_{1}p_{2}z_{2}(p_{1}z_{1}+p_{2}z_{2}+p_{3})^{n-2}.$
We can obtain the expectations appearing in the covariance as:
$\mathbb{E}[\widehat{p}_{i}\widehat{p}_{j}]=\frac{n-1}{n}p_{i}p_{j}$
$\mathbb{E}[\widehat{p}_{i}\widehat{p}_{j}(1-\pi_{i})^{n\widehat{p}_{i}}]=\frac{n-1}{n}p_{i}p_{j}(1-\pi_{i})(1-\pi_{i}p_{i})^{n-2}$
$\mathbb{E}[\widehat{p}_{i}\widehat{p}_{j}(1-\pi_{i})^{n\widehat{p}_{i}}(1-\pi_{j})^{n\widehat{p}_{j}}]=\frac{n-1}{n}p_{i}p_{j}(1-\pi_{i})(1-\pi_{j})(1-\pi_{i}p_{i}-\pi_{j}p_{j})^{n-2}$
Now the covariance of pair $(i,j)$ is
$\,\operatorname{Cov}\left(\widehat{p}_{i}[1-(1-\pi_{i})^{n\widehat{p}_{i}}],\widehat{p}_{j}[1-(1-\pi_{j})^{n\widehat{p}_{j}}]\right)$
$=$ $\,\mathbb{E}\left[\widehat{p}_{i}[1-(1-\pi_{i})^{n\widehat{p}_{i}}]\widehat{p}_{j}[1-(1-\pi_{j})^{n\widehat{p}_{j}}]\right]-\mathbb{E}\left[\widehat{p}_{i}[1-(1-\pi_{i})^{n\widehat{p}_{i}}]\right]\mathbb{E}\left[\widehat{p}_{j}[1-(1-\pi_{j})^{n\widehat{p}_{j}}]\right]$
$=$ $\,\frac{n-1}{n}p_{i}p_{j}-\frac{n-1}{n}p_{i}p_{j}(1-\pi_{i})(1-\pi_{i}p_{i})^{n-2}-\frac{n-1}{n}p_{i}p_{j}(1-\pi_{j})(1-\pi_{j}p_{j})^{n-2}$
$\,+\frac{n-1}{n}p_{i}p_{j}(1-\pi_{i})(1-\pi_{j})(1-\pi_{i}p_{i}-\pi_{j}p_{j})^{n-2}$
$\,-[p_{i}-p_{i}(1-\pi_{i})(1-\pi_{i}p_{i})^{n-1}][p_{j}-p_{j}(1-\pi_{j})(1-\pi_{j}p_{j})^{n-1}]$
$=$ $\,-\frac{1}{n}p_{i}p_{j}+p_{i}p_{j}(1-\pi_{i})(1-\pi_{i}p_{i})^{n-2}\left(\frac{1}{n}-\pi_{i}p_{i}\right)+p_{i}p_{j}(1-\pi_{j})(1-\pi_{j}p_{j})^{n-2}\left(\frac{1}{n}-\pi_{j}p_{j}\right)$
$\,+\frac{1}{n}p_{i}p_{j}(1-\pi_{i})(1-\pi_{j})\left[(n-1)(1-\pi_{i}p_{i}-\pi_{j}p_{j})^{n-2}-n(1-\pi_{i}p_{i})^{n-1}(1-\pi_{j}p_{j})^{n-1}\right]$
where in the last equation we combine the four terms in $\mathbb{E}\left[\widehat{p}_{i}[1-(1-\pi_{i})^{n\widehat{p}_{i}}]\widehat{p}_{j}[1-(1-\pi_{j})^{n\widehat{p}_{j}}]\right]$ and $\mathbb{E}\left[\widehat{p}_{i}[1-(1-\pi_{i})^{n\widehat{p}_{i}}]\right]\mathbb{E}\left[\widehat{p}_{j}[1-(1-\pi_{j})^{n\widehat{p}_{j}}]\right]$ correspondingly. We proceed as taking summations
$\,\bigg{|}\sum_{i\neq j}\mu_{1i}\mu_{1j}\operatorname{Cov}\left(\widehat{p}_{i}[1-(1-\pi_{i})^{n\widehat{p}_{i}}],\widehat{p}_{j}[1-(1-\pi_{j})^{n\widehat{p}_{j}}]\right)\bigg{|}$
$\leq$ $\,\sum_{i\neq j}\frac{1}{n}p_{i}p_{j}+2\sum_{i\neq j}p_{i}p_{j}(1-\pi_{i})(1-\pi_{i}p_{i})^{n-2}\bigg{|}\frac{1}{n}-\pi_{i}p_{i}\bigg{|}$
$\,+\sum_{i\neq j}\frac{1}{n}p_{i}p_{j}(1-\pi_{i})(1-\pi_{j})\big{|}(n-1)(1-\pi_{i}p_{i}-\pi_{j}p_{j})^{n-2}-n(1-\pi_{i}p_{i})^{n-1}(1-\pi_{j}p_{j})^{n-1}\big{|}$
For each individual term in the expression above, we have
$\sum_{i\neq j}\frac{1}{n}p_{i}p_{j}=\sum_{i}\frac{1}{n}p_{i}(1-p_{i})\leq\frac{1}{n}.$
$\,\sum_{i\neq j}p_{i}p_{j}(1-\pi_{i})(1-\pi_{i}p_{i})^{n-2}\bigg{|}\frac{1}{n}-\pi_{i}p_{i}\bigg{|}$
$\leq$ $\,\sum_{i\neq j}\frac{1}{n}p_{i}p_{j}+\sum_{i\neq j}p_{i}p_{j}(1-\pi_{i})(1-\pi_{i}p_{i})^{n-2}\pi_{i}p_{i}$
$\leq$ $\,\frac{1}{n}+\sum_{i\neq j}p_{i}p_{j}(1-\pi_{i})\frac{1}{\pi_{i}p_{i}(n-1)}\pi_{i}p_{i}$
$\leq$ $\,\frac{1}{n}+\frac{1}{n-1}.$
For the last term we need an auxiliary lemma as follows:
Lemma 2.
For $n\geq 3$, consider the function
$f_{2}(x,y)=(n-1)(1-x-y)^{n-2}-n(1-x)^{n-1}(1-y)^{n-1}$
defined on the triangle $D=\{x\geq 0,y\geq 0,x+y\leq 1\}$. Then we have
$|f_{2}(x,y)|\leq 1.$
For the last term we invoke Lemma 2 and have
$\,\sum_{i\neq j}\frac{1}{n}p_{i}p_{j}(1-\pi_{i})(1-\pi_{j})\big{|}(n-1)(1-\pi_{i}p_{i}-\pi_{j}p_{j})^{n-2}-n(1-\pi_{i}p_{i})^{n-1}(1-\pi_{j}p_{j})^{n-1}\big{|}$
$\leq$ $\,\sum_{i\neq j}\frac{1}{n}p_{i}p_{j}\leq\frac{1}{n}.$
We thus showed
$\big{|}\sum_{i\neq j}\mu_{1i}\mu_{1j}\operatorname{Cov}\left(\widehat{p}_{i}[1-(1-\pi_{i})^{n\widehat{p}_{i}}],\widehat{p}_{j}[1-(1-\pi_{j})^{n\widehat{p}_{j}}]\right)\big{|}\leq\frac{4}{n}+\frac{2}{n-1}.$
$\operatorname{Var}\left(\mathbb{E}\left[\sum_{k=1}^{d}\widehat{p}_{k}I(\widehat{w}_{k}>0)\mu_{1k}\bigg{|}\mathbf{X}^{n}\right]\right)\leq\left(1+\frac{2}{\epsilon}\right)\frac{1}{n}+\frac{4}{n}+\frac{2}{n-1}.$
$\operatorname{Var}(\mathbb{E}[\widehat{\psi}_{1}\mid\mathbf{X}^{n},\mathbf{A}^{n}])\leq\left(1+\frac{3}{\epsilon}\right)\frac{1}{n}+\frac{4}{n}+\frac{2}{n-1}.$
$\operatorname{Var}(\widehat{\psi}_{1})\leq\left(1+\frac{7}{2\epsilon}\right)\frac{1}{n}+\frac{4}{n}+\frac{2}{n-1}.$
B.5 Proof of Theorem 2
Proof.
The proof strategy follows from the moment matching method commonly used in theoretical computer science literature (Wu and Yang,, 2016, 2019; Jiao et al.,, 2015). We first define a “relaxed” model class with proportion vector ${\boldsymbol{p}}=(p_{1},p_{2},\dots)$ being approximately a probability vector. Mathematically, for $\delta>0$ define
$\mathcal{D}(\epsilon,\delta)=\left\{\left|\sum_{k=1}^{d}p_{k}-1\right|\leq\delta,p_{k}\in\left[0,1\right],\pi_{k}\in[\epsilon,1-\epsilon],\forall k\in[d]\right\},$ (16)
where we relax the assumption $\sum_{k}p_{k}=1$ to $\left|\sum_{k}p_{k}-1\right|\leq\delta$ so that we can set $p_{k}$’s to be random variables in our construction and use the method of fuzzy hypotheses (Tsybakov,, 2009; Le Cam,, 2012). Under $\mathbb{P}\in\mathcal{D}(\epsilon,\delta)$, we assume $n\widehat{p}_{k}\sim\text{Poi}(np_{k})$ and $(n\widehat{p}_{1},n\widehat{p}_{2},\dots)$ are independent, i.e. we again rely on a Poisson sampling model to prove our results. The treatment assignment $A$ and outcome $Y$ have the same distribution as in Section 2.1 conditioned on the category each sample falls into. Define the minimax lower bound over $\mathcal{D}(\epsilon,\delta)$ as
$\widetilde{R}^{*}(d,n,\delta):=\inf_{\widehat{\psi}_{1}}\sup_{\mathbb{P}\in\mathcal{D}(\epsilon,\delta)}\mathbb{E}_{\mathbb{P}}\left[\left(\widehat{\psi}_{1}-\psi_{1}\right)^{2}\right]$
The following lemma allows us to relate ${R}^{*}(d,n)$ with $\widetilde{R}^{*}(d,n,\delta)$.
Lemma 3.
For any $\delta\in[0,1/3)$,
${R}^{*}(d,n/2)\geq\frac{1}{2}\widetilde{R}^{*}(d,n,\delta)-\exp(-n/50)-\delta^{2}.$
We then present an auxiliary result characterizing the prior distribution on the parameters in the method of fuzzy hypotheses. In the following proof of this section, we will set $\mu_{0k}=0$ and abbreviate $\mu_{1k}$ as $\mu_{k}$, which is different from the mean of outcome in $k$-th category $\mathbb{E}[Y|X=k]$ as in Section 5.
Lemma 4.
There exists constants $c,c^{\prime}>0$ such that for any constants $c_{1},c_{2},c_{3}>0$ satisfying $\frac{cc_{1}c_{3}}{c_{2}^{2}}\leq 1$ and $d\leq c_{3}n\log n$, there exist two distributions $\mu_{0}$ and $\mu_{1}$ on $(p,\pi,\mu)$ satisfying the following properties:
  1. 1.
    $\mu_{0}$ a.s. and $\mu_{1}$ a.s.
[…]
  2. 2.
    For $i,j,k\geq 0$ and $i+j+k\leq 3K$ with $K=c_{2}\log n$,
    $\mathbb{E}_{\mu_{0}}\left[p^{i}(p\pi)^{j}(p\pi\mu)^{k}\right]=\mathbb{E}_{\mu_{1}}\left[p^{i}(p\pi)^{j}(p\pi\mu)^{k}\right].$
  3. 3.
    $a_{0}:=\mathbb{E}_{\mu_{0}}[p]=\mathbb{E}_{\mu_{1}}[p]\leq\frac{1}{d}.$
  4. 4.
    $|\mathbb{E}_{\mu_{0}}[p\mu]-\mathbb{E}_{\mu_{1}}[p\mu]|\geq\frac{c_{4}}{n\log n},$
    where $c_{4}=\frac{cc^{\prime}c_{1}}{c_{2}^{2}}$.
The proof of Lemma 3 and 4 are provided in Section C. Note that in the construction of $\mu_{i}$, $\pi$ and $\mu$ are actually functions of $p$. Under null hypothesis $P$, let $(p_{k},\pi_{k},\mu_{k})$ i.i.d. $\sim\mu_{0},1\leq k\leq d$. We add one more category with $p_{d+1}=1-da_{0},\pi_{d+1}=\epsilon,\mu_{d+1}=0$. Under the alternative hypothesis $P^{\prime}$, let $(p_{k}^{\prime},\pi_{k}^{\prime},\mu_{k}^{\prime})$ i.i.d. $\sim\mu_{1},1\leq k\leq d$ and $p_{d+1}^{\prime}=1-da_{0},\pi_{d+1}^{\prime}=\epsilon,\mu_{d+1}^{\prime}=0$. Obviously, adding one more category will not affect the final rate. The sufficient statistics for $\psi_{1}$ are
#$\#\{i:X_{i}=k,A_{i}=1,Y_{i}=1\},\#\{i:X_{i}=k,A_{i}=1,Y_{i}=0\},\#\{i:X_{i}=k,A_{i}=0\},1\leq k\leq d+1.$
By the property of Poisson distribution these counting statistics are independent under the Poisson sampling model and under the null hypothesis $P$ (given $\{p_{1},\dots,p_{d}\}$, which is equivalent to conditioning on $\{(p_{k},\pi_{k},\mu_{k}),1\leq k\leq d\}$ since $\pi,\mu$ are functions of $p$)
$N_{k11}:=\#\{i:X_{i}=k,A_{i}=1,Y_{i}=1\}\sim\text{Poi}(np_{k}\pi_{k}\mu_{k}),$
$N_{k10}:=\#\{i:X_{i}=k,A_{i}=1,Y_{i}=0\}\sim\text{Poi}(np_{k}\pi_{k}(1-\mu_{k})),$
$N_{k0}:=\#\{i:X_{i}=k,A_{i}=0\}\sim\text{Poi}(np_{k}(1-\pi_{k})).$
Similarly under the alternative hypothesis $P^{\prime}$,
$N_{k11}^{\prime}:=\#\{i:X_{i}=k,A_{i}=1,Y_{i}=1\}\sim\text{Poi}(np_{k}^{\prime}\pi_{k}^{\prime}\mu_{k}^{\prime}),$
$N_{k10}^{\prime}:=\#\{i:X_{i}=k,A_{i}=1,Y_{i}=0\}\sim\text{Poi}(np_{k}^{\prime}\pi_{k}^{\prime}(1-\mu_{k}^{\prime})),$
$N_{k0}^{\prime}:=\#\{i:X_{i}=k,A_{i}=0\}\sim\text{Poi}(np_{k}^{\prime}(1-\pi_{k}^{\prime})).$
Denote $\mathbf{N}_{k}=(N_{k11},N_{k10},N_{k0})$ as the sufficient statistics in $k$-th category and $\mathbf{N}=(\mathbf{N}_{1},\dots,\mathbf{N}_{d},\mathbf{N}_{d+1})$ as the collection of these sufficient statistics. Define $\mathbf{N}^{\prime}$ similarly for the alternative hypothesis. The total variation distance between the marginal distribution of $\mathbf{N}$ and $\mathbf{N}^{\prime}$ (marginalize over the distribution of $\{(p_{k},\pi_{k},\mu_{k}),1\leq k\leq d\}$ and $\{(p_{k}^{\prime},\pi_{k}^{\prime},\mu_{k}^{\prime}),1\leq k\leq d\}$) can be bounded by triangle inequality as (since $\mathbf{N}_{k}$ depends only on $(p_{k},\pi_{k},\mu_{k})$ for each $k$ and $\{(p_{k},\pi_{k},\mu_{k}),1\leq k\leq d\}$ are independent, $(\mathbf{N}_{1},\dots,\mathbf{N}_{d},\mathbf{N}_{d+1})$ are also independent)
$\text{TV}(\mathbf{N},\mathbf{N}^{\prime})\leq\,\sum_{k=1}^{d}\text{TV}(\mathbf{N}_{k},\mathbf{N}_{k}^{\prime}).$
Note that the marginal distributions of components of $\mathbf{N}_{k}$ are not independent since $N_{k11},N_{k10},N_{k0}$ all depend on $(p_{k},\pi_{k},\mu_{k})$. Conditioned on $(p_{k},\pi_{k},\mu_{k})$, $(N_{k11},N_{k10},N_{k0})$ are conditionally independent with each being a Poisson distribution. By definition of total variation distance,
$\,\text{TV}(\mathbf{N}_{k},\mathbf{N}_{k}^{\prime})$
$=$ $\,\frac{1}{2}\sum_{i,j,\ell=0}^{\infty}\big{|}\mathbb{P}(N_{k11}=i,N_{k10}=j,N_{k0}=\ell)-\mathbb{P}(N_{k11}^{\prime}=i,N_{k10}^{\prime}=j,N_{k0}^{\prime}=\ell)\big{|}.$
Conditioning on $(p_{k},\pi_{k},\mu_{k})$ we have
$\,\mathbb{P}(N_{k11}=i,N_{k10}=j,N_{k0}=\ell)$
$=$ $\,\mathbb{E}\left[\exp(-np_{k}\pi_{k}\mu_{k})\frac{(np_{k}\pi_{k}\mu_{k})^{i}}{i!}\exp(-np_{k}\pi_{k}(1-\mu_{k}))\frac{[np_{k}\pi_{k}(1-\mu_{k})]^{j}}{j!}\right.$
$\,\left.\exp(-np_{k}(1-\pi_{k}))\frac{[np_{k}(1-\pi_{k})]^{\ell}}{\ell!}\right]$
$=$ $\,\mathbb{E}\left[\exp(-np_{k})\frac{(np_{k}\pi_{k}\mu_{k})^{i}[np_{k}\pi_{k}(1-\mu_{k})]^{j}[np_{k}(1-\pi_{k})]^{\ell}}{i!j!\ell!}\right]$
$=$ $\,\mathbb{E}\left[\sum_{t=0}^{\infty}\frac{(-np_{k})^{t}(np_{k}\pi_{k}\mu_{k})^{i}[np_{k}\pi_{k}(1-\mu_{k})]^{j}[np_{k}(1-\pi_{k})]^{\ell}}{i!j!\ell!t!}\right].$
Hence the total variation distance can be written as
$\,\text{TV}(\mathbf{N}_{k},\mathbf{N}_{k}^{\prime})$
$=$ $\,\frac{1}{2}\sum_{i,j,\ell}\left|\mathbb{E}\left[\sum_{t=0}^{\infty}\frac{(-np_{k})^{t}(np_{k}\pi_{k}\mu_{k})^{i}[np_{k}\pi_{k}(1-\mu_{k})]^{j}[np_{k}(1-\pi_{k})]^{\ell}}{i!j!\ell!t!}\right]\right.$
$\,\left.-\mathbb{E}\left[\sum_{t=0}^{\infty}\frac{(-np_{k}^{\prime})^{t}(np_{k}^{\prime}\pi_{k}^{\prime}\mu_{k}^{\prime})^{i}[np_{k}^{\prime}\pi_{k}^{\prime}(1-\mu_{k}^{\prime})]^{j}[np_{k}^{\prime}(1-\pi_{k}^{\prime})]^{\ell}}{i!j!\ell!t!}\right]\right|.$
Note that
$(np_{k})^{t}(np_{k}\pi_{k}\mu_{k})^{i}[np_{k}\pi_{k}(1-\mu_{k})]^{j}[np_{k}(1-\pi_{k})]^{\ell}$
is a polynomial of $(p_{k},p_{k}\pi_{k},p_{k}\pi_{k}\mu_{k})$ with degree $t+i+j+\ell$. By moment matching property $1$ in Theorem 4, we have
$\,\text{TV}(\mathbf{N}_{k},\mathbf{N}_{k}^{\prime})$
$\leq$ $\,\frac{1}{2}\left\{\sum_{i+j+\ell+t>3K}\mathbb{E}\left[\frac{(np_{k})^{t}(np_{k}\pi_{k}\mu_{k})^{i}[np_{k}\pi_{k}(1-\mu_{k})]^{j}[np_{k}(1-\pi_{k})]^{\ell}}{i!j!\ell!t!}\right]\right.$
$\,\left.+\mathbb{E}\left[\frac{(np_{k}^{\prime})^{t}(np_{k}^{\prime}\pi_{k}^{\prime}\mu_{k}^{\prime})^{i}[np_{k}^{\prime}\pi_{k}^{\prime}(1-\mu_{k}^{\prime})]^{j}[np_{k}^{\prime}(1-\pi_{k}^{\prime})]^{\ell}}{i!j!\ell!t!}\right]\right\}.$
Note that
$\,\frac{(np_{k})^{t}(np_{k}\pi_{k}\mu_{k})^{i}[np_{k}\pi_{k}(1-\mu_{k})]^{j}[np_{k}(1-\pi_{k})]^{\ell}}{i!j!\ell!t!}$
$=$ $\,\exp(2np_{k})\exp(-np_{k})\frac{(np_{k})^{t}}{t!}\exp(-np_{k}\pi_{k}\mu_{k})\frac{(np_{k}\pi_{k}\mu_{k})^{i}}{i!}$
$\,\exp(-np_{k}\pi_{k}(1-\mu_{k}))\frac{[np_{k}\pi_{k}(1-\mu_{k})]^{j}}{j!}\exp(-np_{k}(1-\pi_{k}))\frac{[np_{k}(1-\pi_{k})]^{\ell}}{\ell!}$
$=$ $\,\exp(2np_{k})\mathbb{P}(V_{1}=t,V_{2}=i,V_{3}=j,V_{4}=\ell\mid p_{k},\pi_{k},\mu_{k}),$
where conditioning on $(p_{k},\pi_{k},\mu_{k})$,
$V_{1}\sim$ $\,\text{Poi}(np_{k}),$
$V_{2}\sim$ $\,\text{Poi}(np_{k}\pi_{k}\mu_{k}),$
$V_{3}\sim$ $\,\text{Poi}(np_{k}\pi_{k}(1-\mu_{k})),$
$V_{4}\sim$ $\,\text{Poi}(np_{k}(1-\pi_{k})),$
and $(V_{1},V_{2},V_{3},V_{4})$ are independent. Further let $V=V_{1}+V_{2}+V_{3}+V_{4}\sim\text{Poi}(2np_{k})$ given $(p_{k},\pi_{k},\mu_{k})$. Thus we have
$\,\sum_{i+j+\ell+t>3K}\mathbb{E}\left[\frac{(np_{k})^{t}(np_{k}\pi_{k}\mu_{k})^{i}[np_{k}\pi_{k}(1-\mu_{k})]^{j}[np_{k}(1-\pi_{k})]^{\ell}}{i!j!\ell!t!}\right]$
$=$ $\,\mathbb{E}[\exp(2np_{k})\mathbb{P}(V_{1}+V_{2}+V_{3}+V_{4}>3K\mid p_{k},\pi_{k},\mu_{k})]$
$=$ $\,\mathbb{E}[\exp(2np_{k})\mathbb{P}(V>3K\mid p_{k},\pi_{k},\mu_{k})]$
$=$ $\,\mathbb{E}\left[\exp(2np_{k})\sum_{\ell>3K}\exp(-2np_{k})\frac{(2np_{k})^{\ell}}{\ell!}\right]$
$=$ $\,\mathbb{E}\left[\sum_{\ell>3K}\frac{(2np_{k})^{\ell}}{\ell!}\right]$
$\leq$ $\,\sum_{\ell>3K}\frac{(2c_{1}\log n)^{\ell}}{\ell!}$
$=$ $\,\exp(2c_{1}\log n)\sum_{\ell>3K}\exp(-2c_{1}\log n)\frac{(2c_{1}\log n)^{\ell}}{\ell!}$
$=$ $\,n^{2c_{1}}\mathbb{P}(W>3K),$
where $W\sim\text{Poi}(2c_{1}\log n)$. Apply the following Chernoff bound: For $L>eM$
$\mathbb{P}(\text{Poi}(M)>L)\leq\exp(-M)\left(\frac{eM}{L}\right)^{L}.$ (17)
When $3K>2ec_{1}\log n$, we have
$\mathbb{P}(W>3K)\leq\exp(-2c_{1}\log n)\left(\frac{2ec_{1}}{3c_{2}}\right)^{3c_{2}\log n},$
$n^{2c_{1}}\mathbb{P}(W>3K)\leq\left(\frac{2ec_{1}}{3c_{2}}\right)^{3c_{2}\log n}=n^{3c_{2}\log\left(\frac{2ec_{1}}{3c_{2}}\right)}\leq n^{-2}\leq\frac{c_{3}\log n}{nd}$
as long as we choose constant $c_{1},c_{2}$ satisfying $3c_{2}\log\left(\frac{3c_{2}}{2ec_{1}}\right)\geq 2$ and $c_{3}$ is the constant in Lemma 4. Thus the total variation distance is bounded as (similar inequality holds under the alternative hypothesis)
$\text{TV}(\mathbf{N}_{k},\mathbf{N}_{k}^{\prime})\leq\frac{c_{3}\log n}{nd}$
$\text{TV}(\mathbf{N},\mathbf{N}^{\prime})\leq\frac{c_{3}\log n}{n}.$
The functional separation is
$\psi_{1}(P)-\psi_{1}(P^{\prime})=\sum_{k=1}^{d}(p_{k}\mu_{k}-p_{k}^{\prime}\mu_{k}^{\prime}),$
$\left|\mathbb{E}\left[\psi_{1}(P)-\psi_{1}(P^{\prime})\right]\right|=d|\mathbb{E}[p\mu-p^{\prime}\mu^{\prime}]|\geq\frac{c_{4}d}{n\log n}:=q.$
Consider the following events:
$E=\left\{\bigg{|}\sum_{k=1}^{d}p_{k}-da_{0}\bigg{|}\leq\delta,\bigg{|}\sum_{k=1}^{d}p_{k}\mu_{k}-d\mathbb{E}_{\mu_{0}}[p\mu]\bigg{|}\leq q/4\right\},$
$E^{\prime}=\left\{\bigg{|}\sum_{k=1}^{d}p_{k}^{\prime}-da_{0}\bigg{|}\leq\delta,\bigg{|}\sum_{k=1}^{d}p_{k}^{\prime}\mu_{k}^{\prime}-d\mathbb{E}_{\mu_{1}}[p^{\prime}\mu^{\prime}]\bigg{|}\leq q/4\right\},$
By Chebyshev’s inequality, we have
$\mathbb{P}(E^{c})\leq$ $\,\mathbb{P}\left(\bigg{|}\sum_{k=1}^{d}p_{k}-da_{0}\bigg{|}>\delta\right)$
$\,+\mathbb{P}\left(\bigg{|}\sum_{k=1}^{d}p_{k}\mu_{k}-d\mathbb{E}_{\mu_{0}}[p\mu]\bigg{|}>q/4\right)$
$\leq$ $\,\frac{d\operatorname{Var}_{\mu_{0}}(p)}{\delta^{2}}+\frac{16d\operatorname{Var}_{\mu_{0}}(p\mu)}{q^{2}}$
$\leq$ $\,\frac{c_{1}^{2}d(\log n)^{2}}{\delta^{2}n^{2}}+\frac{16c_{1}^{2}d(\log n)^{2}}{n^{2}q^{2}}$
$=$ $\,\frac{1}{16}+\frac{16c_{1}^{2}(\log n)^{4}}{c_{4}^{2}d},$
where the third inequality follows from the bound $|p\mu|\leq|p|\leq c_{1}\log n/n$ and the last equation follows by setting $\delta=\frac{4c_{1}\sqrt{d}\log n}{n}$. Similarly, we have
$\mathbb{P}(E^{\prime c})\leq\frac{1}{16}+\frac{16c_{1}^{2}(\log n)^{4}}{c_{4}^{2}d}.$
We put the following prior distributions induced by $\{(p_{k},\pi_{k},\mu_{k}),1\leq k\leq d\}$ and $\{(p_{k}^{\prime},\pi_{k}^{\prime},\mu_{k}^{\prime}),1\leq k\leq d\}$ on $P$ and $P^{\prime}$, respectively:
$\pi\stackrel{{\scriptstyle d}}{{=}}P\mid E,\pi^{\prime}\stackrel{{\scriptstyle d}}{{=}}P^{\prime}\mid E^{\prime}.$
Note that under $\pi,\pi^{\prime}$,
$|\psi_{1}(P)-\psi_{1}(P^{\prime})|\geq q/2.$
By triangle inequality, the total variation distance of the sufficient counting statistics $\mathbf{N}$ and $\mathbf{N}^{\prime}$ under two priors is bounded by
$\text{TV}\left({\mathbf{N}\mid E},{\mathbf{N}^{\prime}\mid E^{\prime}}\right)\leq$ $\,\text{TV}\left({\mathbf{N}\mid E},{\mathbf{N}}\right)+\text{TV}({\mathbf{N}},{\mathbf{N}^{\prime}})+\text{TV}\left({\mathbf{N}^{\prime}\mid E^{\prime}},{\mathbf{N}^{\prime}}\right)$
$\leq$ $\,\mathbb{P}(E^{c})+\mathbb{P}(E^{\prime c})+\text{TV}({\mathbf{N}},{\mathbf{N}^{\prime}})$
$\leq$ $\,\frac{1}{8}+\frac{32c_{1}^{2}(\log n)^{4}}{c_{4}^{2}d}+\frac{c_{3}\log n}{n}.$
By method of fuzzy hypotheses (Section 2.7.4 of Tsybakov, (2009)) we conclude
$\,\widetilde{R}^{*}(d,n,\delta)$
$\geq$ $\,\frac{q^{2}}{32}\left(1-\text{TV}\left({\mathbf{N}\mid E},{\mathbf{N}^{\prime}\mid E^{\prime}}\right)\right).$
$\geq$ $\,\frac{q^{2}}{32}\left(\frac{7}{8}-\frac{32c_{1}^{2}(\log n)^{4}}{c_{4}^{2}d}-\frac{c_{3}\log n}{n}\right).$
Hence in the regime $d\gtrsim(\log n)^{4}$, we have
$\widetilde{R}^{*}(d,n,\delta)\gtrsim\frac{d^{2}}{(n\log n)^{2}}.$
By Lemma 3 we conclude
${R}^{*}(d,n)\gtrsim$ $\,\frac{d^{2}}{(n\log n)^{2}}-\exp(-n/50)-\frac{d(\log n)^{2}}{n^{2}}$
$\gtrsim$ $\,\frac{d^{2}}{(n\log n)^{2}}.$
The overall requirements on the constants are
[…]
Clearly for $c,c_{3}>0$, one can choose $c_{1}$ sufficiently small and $c_{2}$ sufficiently large to satisfy these conditions.
The lower bound $1/n$ can be proved using a two-point method. Without loss of generality, assume that $n\geq 8$. Under the null $P$ and alternative hypothesis $P^{\prime}$, set
$p_{k}=\frac{1}{d},\pi_{k}=\frac{1}{2},\mu_{0k}=0,\mu_{1k}(P)=\frac{1}{2},\mu_{1k}(P^{\prime})=\frac{1}{2}+\delta,$
with $\delta=1/\sqrt{n}$, i.e., the covariate distribution and propensity score are the same. By Le Cam’s two-point method we have
$\inf_{\widehat{\psi}_{1}}\sup_{\mathbb{P}\in\mathcal{D}(\epsilon)}\mathbb{E}_{\mathbb{P}}\left[\left(\widehat{\psi}_{1}-\psi_{1}\right)^{2}\right]\geq\frac{1}{4}(\psi_{1}(P)-\psi_{1}(P^{\prime}))^{2}\exp(-nD(P\|P^{\prime})).$ (18)
Note that the functional separation is
$|\psi_{1}(P)-\psi_{1}(P^{\prime})|=\delta=\frac{1}{\sqrt{n}}.$
Under the null hypothesis $P$, we have
$P(X=k,A=a,Y=y)=\left\{\begin{array}[]{rl}\frac{1}{4d}&\text{if }a=1,y=1,\\ \frac{1}{4d}&\text{if }a=1,y=0,\\ \frac{1}{2d}&\text{if }a=0,y=0.\end{array}\right.$
Under the alternative hypothesis $P^{\prime}$, we have
$P^{\prime}(X=k,A=a,Y=y)=\left\{\begin{array}[]{rl}\frac{1}{2d}\left(\frac{1}{2}+\delta\right)&\text{if }a=1,y=1,\\ \frac{1}{2d}\left(\frac{1}{2}-\delta\right)&\text{if }a=1,y=0,\\ \frac{1}{2d}&\text{if }a=0,y=0.\end{array}\right.$
The K-L divergence between $P$ and $P^{\prime}$ is
$D(P\|P^{\prime})=\sum_{k=1}^{d}\left[-\frac{1}{4d}\log(1+2\delta)-\frac{1}{4d}\log(1-2\delta)\right]=-\frac{1}{4}\log(1-4\delta^{2}).$
Using the inequality
$\log(1-x)\geq-2x,x\in[0,1/2],$
we have
$D(P\|P^{\prime})=-\frac{1}{4}\log\left(1-\frac{4}{n}\right)\leq\frac{2}{n}.$
Plug in the functional separation and bound on K-L divergence into (18), we conclude
$\inf_{\widehat{\psi}_{1}}\sup_{\mathbb{P}\in\mathcal{D}(\epsilon)}\mathbb{E}_{\mathbb{P}}\left[\left(\widehat{\psi}_{1}-\psi_{1}\right)^{2}\right]\gtrsim\frac{1}{n}.$
B.6 Proof of Lemma 1
Proof.
By the definition of $\widehat{t}_{k}=I(\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1)$ we have
$\mathbb{P}\left(\sum_{j=1}^{d}\widehat{t}_{j}=0\right)$ $=\mathbb{P}\left(\widehat{w}_{k}\in\left\{0,\widehat{p}_{k}\right\}\forall k\text{ with }\widehat{p}_{k}>0\right)$
$=\mathbb{E}\left\{\mathbb{P}\left(\widehat{w}_{k}\in\left\{0,\widehat{p}_{k}\right\}\forall k\text{ with }\widehat{p}_{k}>0\mid\mathbf{X}^{n}\right)\right\}$
$=\mathbb{E}\left\{\prod_{k:\widehat{p}_{k}>0}\mathbb{P}\left(\widehat{w}_{k}\in\left\{0,\widehat{p}_{k}\right\}\mid\mathbf{X}^{n}\right)\right\}$
$=\mathbb{E}\left(\prod_{k=1}^{d}\left[\left\{\left(1-\pi_{k}\right)^{n\widehat{p}_{k}}+\pi_{k}^{n\widehat{p}_{k}}\right\}I\left(\widehat{p}_{k}>0\right)+I\left(\widehat{p}_{k}=0\right)\right]\right)$
$=\mathbb{E}\left[\prod_{k=1}^{d}\left\{\left(1-\pi_{k}\right)^{n\widehat{p}_{k}}+\pi_{k}^{n\widehat{p}_{k}}-I\left(\widehat{p}_{k}=0\right)\right\}\right]$
The proof relies on the poissonization technique to bound the above expectation of the product. Poissonization allows us to replace $\left(n\widehat{p}_{1},\ldots,n\widehat{p}_{d}\right)\sim\operatorname{Multinomial}\left(n,p_{1},\ldots,p_{d}\right)$ with $n\widehat{p}_{k}\sim\text{Poisson}(np_{k})$ and $n\widehat{p}_{1},\dots,n\widehat{p}_{d}$ are independent. The following lemma connects the expectation in the multinomial case with that in the independent Poisson case.
Lemma 5.
(Theorem 5.10 in Mitzenmacher and Upfal, (2017)) Let $\mathbf{X}^{n}\in\mathbb{R}^{d}\sim\text{Multinomial}\left(n,p_{1},\ldots,p_{d}\right)$, $\mathbf{Y}^{n}\in\mathbb{R}^{d}$ and $Y_{i}^{n}\sim\text{Poisson}(np_{i})$ and $Y_{1}^{n},\dots,Y_{d}^{n}$ are independent. Consider a non-negative function $f(x_{1},\dots,x_{d})$, if $\mathbb{E}[f(X_{1}^{n},\dots,X_{d}^{n})]$ is monotonely non-increasing with $n$, then
$\mathbb{E}[f(X_{1}^{n},\dots,X_{d}^{n})]\leq 2\mathbb{E}[f(Y_{1}^{n},\dots,Y_{d}^{n})].$
The proof is left as an exercise in Mitzenmacher and Upfal, (2017) and we include it in Appendix C.4. Let
$a_{n}=$ $\,\mathbb{E}\left[\prod_{k=1}^{d}\left\{\left(1-\pi_{k}\right)^{n\widehat{p}_{k}}+\pi_{k}^{n\widehat{p}_{k}}-I\left(\widehat{p}_{k}=0\right)\right\}\right]$
$=$ $\,\mathbb{E}\left[\prod_{k=1}^{d}\left\{\left(1-\pi_{k}\right)^{\sum_{i=1}^{n}I(X_{i}=k)}+\pi_{k}^{\sum_{i=1}^{n}I(X_{i}=k)}-I\left(\sum_{i=1}^{n}I(X_{i}=k)=0\right)\right\}\right].$
First we verify the monotonicity of $a_{n}$. We claim
$\,\left(1-\pi_{k}\right)^{\sum_{i=1}^{n+1}I(X_{i}=k)}+\pi_{k}^{\sum_{i=1}^{n+1}I(X_{i}=k)}-I\left(\sum_{i=1}^{n+1}I(X_{i}=k)=0\right)$
$\leq$ $\,\left(1-\pi_{k}\right)^{\sum_{i=1}^{n}I(X_{i}=k)}+\pi_{k}^{\sum_{i=1}^{n}I(X_{i}=k)}-I\left(\sum_{i=1}^{n}I(X_{i}=k)=0\right).$
On the event $\left\{\sum_{i=1}^{n+1}I(X_{i}=k)=0\right\}$ then $\sum_{i=1}^{n}I(X_{i}=k)=0$ and
$\left(1-\pi_{k}\right)^{\sum_{i=1}^{n+1}I(X_{i}=k)}+\pi_{k}^{\sum_{i=1}^{n+1}I(X_{i}=k)}-I\left(\sum_{i=1}^{n+1}I(X_{i}=k)=0\right)=1$
$\left(1-\pi_{k}\right)^{\sum_{i=1}^{n}I(X_{i}=k)}+\pi_{k}^{\sum_{i=1}^{n}I(X_{i}=k)}-I\left(\sum_{i=1}^{n}I(X_{i}=k)=0\right)=1.$
On the event $\left\{\sum_{i=1}^{n}I(X_{i}=k)=0,X_{n+1}=k\right\}$,
$\left(1-\pi_{k}\right)^{\sum_{i=1}^{n+1}I(X_{i}=k)}+\pi_{k}^{\sum_{i=1}^{n+1}I(X_{i}=k)}-I\left(\sum_{i=1}^{n+1}I(X_{i}=k)=0\right)=1-\pi_{k}+\pi_{k}=1$
$\left(1-\pi_{k}\right)^{\sum_{i=1}^{n}I(X_{i}=k)}+\pi_{k}^{\sum_{i=1}^{n}I(X_{i}=k)}-I\left(\sum_{i=1}^{n}I(X_{i}=k)=0\right)=1.$
On the event $\left\{\sum_{i=1}^{n}I(X_{i}=k)>0\right\}$,
$\,\left(1-\pi_{k}\right)^{\sum_{i=1}^{n+1}I(X_{i}=k)}+\pi_{k}^{\sum_{i=1}^{n+1}I(X_{i}=k)}-I\left(\sum_{i=1}^{n+1}I(X_{i}=k)=0\right)$
$=$ $\,\left(1-\pi_{k}\right)^{\sum_{i=1}^{n+1}I(X_{i}=k)}+\pi_{k}^{\sum_{i=1}^{n+1}I(X_{i}=k)}$
$\leq$ $\,\left(1-\pi_{k}\right)^{\sum_{i=1}^{n}I(X_{i}=k)}+\pi_{k}^{\sum_{i=1}^{n}I(X_{i}=k)}$
$=$ $\,\left(1-\pi_{k}\right)^{\sum_{i=1}^{n}I(X_{i}=k)}+\pi_{k}^{\sum_{i=1}^{n}I(X_{i}=k)}-I\left(\sum_{i=1}^{n}I(X_{i}=k)=0\right).$
Hence the claim is verified and $a_{n}$ is non-increasing. Apply Lemma 5 we can now assume $n\widehat{p}_{k}\sim\text{Poisson}(np_{k})$ and $n\widehat{p}_{1},\dots,n\widehat{p}_{d}$ are independent (with an additional factor 2)
$\,\mathbb{P}\left(\sum_{j=1}^{d}\widehat{t}_{j}=0\right)$
$\leq$  $\,2\mathbb{E}\left[\prod_{k=1}^{d}\left\{\left(1-\pi_{k}\right)^{n\widehat{p}_{k}}+\pi_{k}^{n\widehat{p}_{k}}-I\left(\widehat{p}_{k}=0\right)\right\}\right]$
$=$  $\,2\prod_{k=1}^{d}\mathbb{E}\left[\left(1-\pi_{k}\right)^{n\widehat{p}_{k}}+\pi_{k}^{n\widehat{p}_{k}}-I\left(\widehat{p}_{k}=0\right)\right]$
$=$  $\,2\prod_{k=1}^{d}\left[\exp(-n\pi_{k}p_{k})+\exp(-n(1-\pi_{k})p_{k})-\exp(-np_{k})\right]$
$\leq$  $\,2\prod_{k=1}^{d}\left[\exp(-\epsilon np_{k})+\exp(-(1-\epsilon)np_{k})-\exp(-np_{k})\right].$
The first equation follows from independence and second one follows from probability generating function of Poisson distribution and the last inequality follows since $f_{3}(x)=\exp(-cx)+\exp(-c(1-x))$ (for constant $c>0$) is decreasing on $[0,1/2]$ and increasing on $[1/2,1]$. For simplicity define
$Z_{k}=\exp(-\epsilon np_{k})+\exp(-(1-\epsilon)np_{k})-\exp(-np_{k}).$
1. In the first case $\epsilon np_{k}\geq 2\log 2$ we have (recall $20<\epsilon<1/$ 2)
$Z_{k}\leq 2\exp(-\epsilon np_{k})\leq\exp\left(-\frac{\epsilon np_{k}}{2}\right).$
2. In the second case $np_{k}\leq\epsilon$, we use the following inequalities for $t\geq 0$:
$\exp(-t)\leq 1-t+t^{2}/2,$
$\exp(-t)\geq 1-t+t^{2}/2-t^{3}/6$
and obtain
$Z_{k}\leq$  $\,1-\epsilon np_{k}+\frac{\epsilon^{2}n^{2}p_{k}^{2}}{2}+1-(1-\epsilon)np_{k}+\frac{(1-\epsilon)^{2}n^{2}p_{k}^{2}}{2}$
$\,-\left(1-np_{k}+\frac{n^{2}p_{k}^{2}}{2}-\frac{n^{3}p_{k}^{3}}{6}\right)$
$=$  $\,1-\epsilon(1-\epsilon)n^{2}p_{k}^{2}+\frac{n^{3}p_{k}^{3}}{6}$
$\leq$  $\,1-\left(\frac{5}{6}\epsilon-\epsilon^{2}\right)n^{2}p_{k}^{2}$
$\leq$  $\,1-\frac{\epsilon n^{2}p_{k}^{2}}{3}$
$\leq$ $\,\exp\left(-\frac{\epsilon n^{2}p_{k}^{2}}{3}\right).$
3. In the last case $\epsilon<np_{k}<\frac{2\log 2}{\epsilon}$, since $f_{4}(x)=\exp(-\epsilon x)+\exp(-(1-\epsilon)x)-\exp(-x)$ is monotonely non-increasing on $[0,+\infty]$ (one can check this by taking the first-order derivative easily), we have
$Z_{k}\leq\exp(-\epsilon^{2})+\exp(-\epsilon(1-\epsilon))-\exp(-\epsilon)=\exp(-C_{1}(\epsilon))$
where $C_{1}(\epsilon)=-\log(\exp(-\epsilon^{2})+\exp(-\epsilon(1-\epsilon))-\exp(-\epsilon))\in(0,\epsilon^{2})$. Note that
$1>\frac{\epsilon np_{k}}{2\log 2}.$
We have
$\exp(-C_{1}(\epsilon))\leq\exp(-C_{2}(\epsilon)np_{k})$
where
$C_{2}(\epsilon)=\frac{C_{1}(\epsilon)\epsilon}{2\log 2}<\frac{\epsilon^{3}}{2\log 2}<\frac{\epsilon}{2}$
Hence we can combine the first case with the third case as
$Z_{k}\leq\exp(-C_{2}(\epsilon)np_{k}).$
Let $I_{1}$ include the index $k\in[d]$ in the first and third case, $I_{2}$ include the index $k$ in the second case. Denote $S_{1}=\sum_{i\in I_{1}}p_{i},T_{2}=\sum_{i\in I_{2}}p_{i}^{2}$. We thus have
$\prod_{k=1}^{d}Z_{k}\leq\exp(-C_{2}(\epsilon)nS_{1})\exp\left(-\frac{\epsilon n^{2}T_{2}}{3}\right).$
In the case $n\geq d$ we have
$1-S_{1}=\sum_{i\in I_{2}}p_{i}\leq\frac{\epsilon|I_{2}|}{n}\leq\frac{\epsilon d}{n}\leq\epsilon,$
i.e. $S_{1}>1-\epsilon$ and we have
$\prod_{k=1}^{d}Z_{k}\leq\exp(-C_{2}(\epsilon)(1-\epsilon)n)\leq\exp\left(-\frac{C_{2}(\epsilon)n}{2}\right).$
In the case $n<d$, if $S_{1}\geq 1/2$ then we also have
$\prod_{k=1}^{d}Z_{k}\leq\exp\left(-\frac{C_{2}(\epsilon)n}{2}\right).$
In the case $S_{1}<1/2$, then by Cauchy-Schwarz inequality we have
$\frac{1}{4}\leq(1-S_{1})^{2}=\left(\sum_{i\in I_{2}}p_{i}\right)^{2}\leq|I_{2}|T_{2}\leq dT_{2},$
i.e. $T_{2}\geq\frac{1}{4d}$, this further implies
$\prod_{k=1}^{d}Z_{k}\leq\exp\left(-\frac{\epsilon n^{2}}{12d}\right).$
So we conclude that
$\prod_{k=1}^{d}Z_{k}\leq\max\left\{\exp\left(-\frac{C_{2}(\epsilon)n}{2}\right),\exp\left(-\frac{\epsilon n^{2}}{12d}\right)\right\}\leq\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right)$
where
$C(\epsilon)=\min\left(\frac{C_{2}(\epsilon)}{2},\frac{\epsilon}{12}\right).$
B.7 Proof of Theorem 3
Proof.
We first bound the bias. By conditioning on $\mathbf{X}^{n},\mathbf{A}^{n}$ and noting $\mathbb{E}[\widehat{\mu}_{1k}\mid\mathbf{X}^{n},\mathbf{A}^{n}]=\mu_{1k}I(\widehat{p}_{k}\widehat{\pi}_{k}>0)$, $\mathbb{E}[\widehat{\mu}_{0k}\mid\mathbf{X}^{n},\mathbf{A}^{n}]=\mu_{0k}I(\widehat{p}_{k}(1-\widehat{\pi}_{k})>0)$ we have
$\mathbb{E}\left[\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]=\frac{\sum_{k=1}^{d}\widehat{t}_{k}\tau_{k}}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right).$
Note that
$\psi=\psi I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)+\psi I\left(\sum_{j=1}^{d}\widehat{t}_{j}=0\right)=\frac{\sum_{k=1}^{d}\widehat{t}_{k}\psi}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)+\psi I\left(\sum_{j=1}^{d}\widehat{t}_{j}=0\right).$
Hence the bias can be expressed as
$\,\mathbb{E}\left[\widehat{\psi}-\psi\right]$
$=$ $\,\mathbb{E}\left[\frac{\sum_{k=1}^{d}\widehat{t}_{k}(\tau_{k}-\psi)}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)\right]-\psi\mathbb{P}\left(\sum_{j=1}^{d}\widehat{t}_{j}=0\right)$
Using $\sigma_{n}$ in (8) to parameterize the effect homogeneity and Lemma 1, we have
$\,\left|\mathbb{E}\left[\widehat{\psi}-\psi\right]\right|$
$\leq$ $\,\mathbb{E}\left[\frac{\sum_{k=1}^{d}\widehat{t}_{k}|\tau_{k}-\psi|}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)\right]+|\psi|\mathbb{P}\left(\sum_{j=1}^{d}\widehat{t}_{j}=0\right)$
$\leq$ $\,\sigma_{n}+2\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$
We then consider the variance. By the property of conditional variance, we have
$\operatorname{Var}(\widehat{\psi})=\operatorname{Var}\left(\mathbb{E}\left[\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]\right)+\mathbb{E}\left[\operatorname{Var}\left(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right].$ (19)
Rewrite
$\mathbb{E}\left[\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]=\frac{\sum_{k=1}^{d}\widehat{t}_{k}\tau_{k}}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)=\frac{\sum_{k=1}^{d}\widehat{t}_{k}(\tau_{k}-\psi)}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)+\psi I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right).$
For a bounded random variable $X$ satisfying $m\leq X\leq M$ we have $\operatorname{Var}(X)\leq(M-m)^{2}/4$. Note that
$\left|\frac{\sum_{k=1}^{d}\widehat{t}_{k}(\tau_{k}-\psi)}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)\right|\leq\sigma_{n}$
and we have
$\operatorname{Var}\left(\frac{\sum_{k=1}^{d}\widehat{t}_{k}(\tau_{k}-\psi)}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)\right)\leq\sigma_{n}^{2}.$
Further notice
$\operatorname{Var}\left(\psi I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)\right)\leq\psi^{2}\mathbb{P}\left(\sum_{j=1}^{d}\widehat{t}_{j}=0\right)\leq 2\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$
The covariance can be expressed as
$\,\operatorname{Cov}\left(\frac{\sum_{k=1}^{d}\widehat{t}_{k}(\tau_{k}-\psi)}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right),\psi I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)\right)$
$=$ $\,\psi\mathbb{E}\left[\frac{\sum_{k=1}^{d}\widehat{t}_{k}(\tau_{k}-\psi)}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)\right]-\psi\mathbb{E}\left[\frac{\sum_{k=1}^{d}\widehat{t}_{k}(\tau_{k}-\psi)}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)\right]\mathbb{P}\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)$
$=$ $\,\psi\mathbb{E}\left[\frac{\sum_{k=1}^{d}\widehat{t}_{k}(\tau_{k}-\psi)}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)\right]\mathbb{P}\left(\sum_{j=1}^{d}\widehat{t}_{j}=0\right)$
and hence we can bound the covariance as follows:
$\,\left|\operatorname{Cov}\left(\frac{\sum_{k=1}^{d}\widehat{t}_{k}(\tau_{k}-\psi)}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right),\psi I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)\right)\right|$
$\leq$  $\,2\sigma_{n}\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$
Combining these three terms above we have
$\operatorname{Var}\left(\mathbb{E}\left[\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]\right)\leq\sigma_{n}^{2}+2\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right)+4\sigma_{n}\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$
To complete the proof we need to bound
$\mathbb{E}\left[\operatorname{Var}\left(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right].$
Note that conditioning on $\mathbf{X}^{n},\mathbf{A}^{n}$ the estimated $\{\widehat{\tau}_{k},1\leq k\leq d\}$ are independent with
$\operatorname{Var}(\widehat{\tau}_{k}\mid\mathbf{X}^{n},\mathbf{A}^{n})=\frac{\mu_{1k}(1-\mu_{1k})}{n\widehat{p}_{k}\widehat{\pi}_{k}}I(\widehat{p}_{k}\widehat{\pi}_{k}>0)+\frac{\mu_{0k}(1-\mu_{0k})}{n\widehat{p}_{k}(1-\widehat{\pi}_{k})}I\left(\widehat{p}_{k}(1-\widehat{\pi}_{k})>0\right).$
We have
$\operatorname{Var}(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n})=\frac{I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)}{\left(\sum_{j=1}^{d}\widehat{t}_{j}\right)^{2}}\sum_{k=1}^{d}\widehat{t}_{k}\left(\frac{\mu_{1k}(1-\mu_{1k})}{n\widehat{p}_{k}\widehat{\pi}_{k}}+\frac{\mu_{0k}(1-\mu_{0k})}{n\widehat{p}_{k}(1-\widehat{\pi}_{k})}\right).$
Let $W=\sum_{k=1}^{d}\widehat{t}_{k}\in\sigma(\mathbf{X}^{n},\mathbf{A}^{n})$ be the number of categories that contain both treated and untreated units. On the event $\{W=m\}$ for some $m\in[d]$ we have
$\,\operatorname{Var}(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n})I(W=m)$
$\leq$ $\,\frac{I(W=m)}{4m^{2}}\sum_{k=1}^{d}\widehat{t}_{k}\left(\frac{1}{n\widehat{p}_{k}\widehat{\pi}_{k}}+\frac{1}{n\widehat{p}_{k}(1-\widehat{\pi}_{k})}\right).$
Notice that $\widehat{t}_{k}=I(\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1)$ and on the event $\{\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1\}$ we have $n\widehat{p}_{k}\widehat{\pi}_{k}\geq 1,n\widehat{p}_{k}(1-\widehat{\pi}_{k})\geq 1$. We thus have
$\operatorname{Var}(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n})I(W=m)\leq\frac{I(W=m)}{2m^{2}}\sum_{k=1}^{d}\widehat{t}_{k}=\frac{I(W=m)}{2m}.$
Simple observation shows $\operatorname{Var}(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n})I(W=0)=0$. Thus we have
$\,\mathbb{E}\left[\operatorname{Var}\left(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right]$
$=$ $\,\sum_{m=1}^{d}\mathbb{E}\left[I(W=m)\operatorname{Var}\left(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right]$
$\leq$ $\,\sum_{m=1}^{d}\frac{\mathbb{P}(W=m)}{2m}$
$=$ $\,\frac{1}{2}\mathbb{E}\left[\frac{I(W>0)}{W}\right],$
where the estimated proportion $n\widehat{{\boldsymbol{p}}}\sim\text{Multinomial}(n,{\boldsymbol{p}})$. In the following analysis, we first use Poissonization to reduce the problem to Poisson sampling scheme. Let
$f_{5}(x)=\frac{I(x>0)}{x}+I(x=0)$
and note that $f_{5}$ is non-increasing on $\{0,1,\dots,d\}$. Also as the sample size $n$ increases, the number of levels with both treated and untreated units will not decrease, i.e. $W$ is non-decreasing as we collect more samples in one experiment. This shows
$\mathbb{E}[f_{5}(W)]$
is non-increasing in $n$ and hence by Poissonization Lemma 5 we have
$\mathbb{E}_{n\widehat{{\boldsymbol{p}}}\sim\text{Multi}(n,{\boldsymbol{p}})}[f_{5}(W)]\leq 2\mathbb{E}_{n\widehat{{\boldsymbol{p}}}\sim\text{Poi}(n{\boldsymbol{p}})}[f_{5}(W)]$
where we use $n\widehat{{\boldsymbol{p}}}\sim\text{Poi}(n{\boldsymbol{p}})$ to denote $n\widehat{p}_{i}\sim\text{Poi}(np_{i})$ and $\{n\widehat{p}_{1},\dots,n\widehat{p}_{d}\}$ are independent. This implies
$\,\mathbb{E}\left[\operatorname{Var}\left(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right]$
$\leq$ $\,\frac{1}{2}\mathbb{E}_{n\widehat{{\boldsymbol{p}}}\sim\text{Multi}(n,{\boldsymbol{p}})}\left[\frac{I(W>0)}{W}\right]$
$\leq$ $\,\mathbb{E}_{n\widehat{{\boldsymbol{p}}}\sim\text{Poi}(n{\boldsymbol{p}})}\left[\frac{I(W>0)}{W}+I(W=0)\right]$
$\leq$ $\,\mathbb{E}_{n\widehat{{\boldsymbol{p}}}\sim\text{Poi}(n{\boldsymbol{p}})}\left[\frac{I(W>0)}{W}\right]+\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right),$
where the last inequality follows since in the proof of Lemma 1 we show
$\mathbb{P}_{n\widehat{{\boldsymbol{p}}}\sim\text{Poi}(n{\boldsymbol{p}})}(W=0)\leq\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$
In the following derivation we will assume $n\widehat{{\boldsymbol{p}}}\sim\text{Poi}(n{\boldsymbol{p}})$ and bound
$\mathbb{E}\left[\frac{I(W>0)}{W}\right].$
Now $W=\sum_{k=1}^{d}I(\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1)$ is a summation of independent Bernoulli random variable with potentially different mean. Let
$q_{k}=\mathbb{P}(\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1)=1+\exp(-np_{k})-\exp(-np_{k}(1-\pi_{k}))-\exp(-np_{k}\pi_{k}),$
$\bar{q}=\frac{1}{d}\sum_{k=1}^{d}q_{k}$ and $\overline{W}\sim\text{B}(d,\bar{q})$. We will use the following lemma to relate $W$ with $\overline{W}$.
Lemma 6.
(Theorem 3 in Hoeffding, (1956)) Let $X=\sum_{i=1}^{n}\xi_{i}$ where $\xi_{1},\dots,\xi_{n}$ are independent Bernoulli random variables with parameters $p_{1},\dots,p_{n}$. Let $\bar{p}=\frac{1}{n}\sum_{i=1}^{n}p_{i}$ and $\overline{X}\sim\text{B}(n,\bar{p})$. For any convex function $g$ in the sense that
[…]
we have
$\mathbb{E}[g(X)]\leq\mathbb{E}[g(\overline{X})]$
To use Lemma 6, for any $\delta>0$ let
$f_{\delta}(x)=\frac{I(x>0)}{x}+I(x=0)\left(\frac{3}{2}+\delta\right).$
One can verify $f_{\delta}$ is convex in the sense of Lemma 6 and hence $\mathbb{E}[f_{\delta}(X)]\leq\mathbb{E}[f_{\delta}(\overline{X})]$, i.e.
$\mathbb{E}\left[\frac{I(W>0)}{W}\right]+\left(\frac{3}{2}+\delta\right)\mathbb{P}(W=0)\leq\mathbb{E}\left[\frac{I(\overline{W}>0)}{\overline{W}}\right]+\left(\frac{3}{2}+\delta\right)\mathbb{P}(\overline{W}=0).$
Since $\delta>0$ is arbitrary we conclude
$\mathbb{E}\left[\frac{I(W>0)}{W}\right]+\frac{3}{2}\mathbb{P}(W=0)\leq\mathbb{E}\left[\frac{I(\overline{W}>0)}{\overline{W}}\right]+\frac{3}{2}\mathbb{P}(\overline{W}=0).$
By Lemma 7 we have
$\mathbb{E}\left[\frac{I(\overline{W}>0)}{\overline{W}}\right]\leq\frac{2}{d\bar{q}}$
Since $\pi_{k}\in[\epsilon,1-\epsilon]$, we have
$q_{k}\in[1+\exp(-np_{k})-\exp(-np_{k}(1-\epsilon))-\exp(-np_{k}\epsilon),1+\exp(-np_{k})-2\exp(-np_{k}/2)].$
Note that we assume $p_{k}\geq\alpha/d$ for all $k$ and by using the following two inequalities for $t>0$
$\exp(-t)\leq 1-t+t^{2}/2,$
$\exp(-t)\geq 1-t+t^{2}/2-t^{3}/6$
we have
$q_{k}\geq$  $\,1+\exp\left(-\frac{\alpha n}{d}\right)-\exp\left(-\frac{\epsilon\alpha n}{d}\right)-\exp\left(-\frac{(1-\epsilon)\alpha n}{d}\right)$
$\geq$  $\,1+1-\frac{\alpha n}{d}+\frac{\alpha^{2}n^{2}}{2d^{2}}-\frac{\alpha^{3}n^{3}}{6d^{3}}-1+\frac{\epsilon\alpha n}{d}-\frac{\epsilon^{2}\alpha^{2}n^{2}}{2d^{2}}-1+\frac{(1-\epsilon)\alpha n}{d}-\frac{(1-\epsilon)^{2}\alpha^{2}n^{2}}{2d^{2}}$
$=$ $\,\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{d^{2}}-\frac{\alpha^{3}n^{3}}{6d^{3}}$
$\geq$ $\,\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{2d^{2}}$
in the regime $n\leq\frac{3\epsilon(1-\epsilon)d}{\alpha}$. Thus we have
$\frac{2}{d\bar{q}}\leq\frac{4d}{\alpha^{2}\epsilon(1-\epsilon)n^{2}},$
$\mathbb{P}(\overline{W}=0)=(1-\bar{q})^{d}\leq\left(1-\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{2d^{2}}\right)^{d}\leq\exp\left(-\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{2d}\right).$
Finally we obtain the following bound on $\mathbb{E}\left[\frac{I(W>0)}{W}\right]$ under Poissonization
$\mathbb{E}\left[\frac{I(W>0)}{W}\right]\leq\frac{4d}{\alpha^{2}\epsilon(1-\epsilon)n^{2}}+\frac{3}{2}\exp\left(-\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{2d}\right),$
which yields a bound
$\mathbb{E}\left[\operatorname{Var}\left(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right]\leq\frac{4d}{\alpha^{2}\epsilon(1-\epsilon)n^{2}}+\frac{3}{2}\exp\left(-\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{2d}\right)+\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$
Combining this inequality with equation (19) completes the proof. ∎
B.8 Proof of Theorem 4
Proof.
By the same argument in Lemma 3, we only need to prove the result under the Poisson sampling model, where $(n\widehat{p}_{1},\dots,n\widehat{p}_{d})\text{ i.i.d}\sim\text{Poi}(\lambda)$ with $\lambda=\frac{n}{d}$. Let $\nu_{0},\nu_{1}$ be two prior distributions on propensity score $\pi$ defined on $[\epsilon,1-\epsilon]$ satisfying
  • $\mathbb{E}_{\nu_{0}}[\pi^{j}]=\mathbb{E}_{\nu_{1}}[\pi^{j}],1\leq j\leq 2L,L=\lfloor\frac{1}{\beta}\rfloor+1$.
  • $\mathbb{E}_{\nu_{0}}\left[\frac{1}{\pi}\right]-\mathbb{E}_{\nu_{1}}\left[\frac{1}{\pi}\right]=c_{1}(\beta,\epsilon)>0.$
The existence of $\nu_{0},\nu_{1}$ follows from the duality between moment matching and best polynomial approximation. In fact, one can show
$c_{1}(\beta,\epsilon)=2\mathbb{E}_{2L}\left(\frac{1}{x};[\epsilon,1-\epsilon]\right),$ (20)
where $E_{n}(f;S)$ is the best polynomial (with order no greater than $n$) approximation error of $f$ on the interval $S$. Since $L$ depends on $\beta$, the RHS of (20) is a constant only depending on $\beta,\epsilon$.
Under null hypothesis $P$, let $(\pi_{1},\dots,\pi_{d})$ i.i.d. $\sim\nu_{0}$ and $\mu_{k}=\epsilon/\pi_{k}$. The sufficient statistics for $\psi_{1}$ (conditioning on $(\pi_{1},\dots,\pi_{d})$) are
$N_{k11}:=\#\{i:X_{i}=k,A_{i}=1,Y_{i}=1\}\sim\text{Poi}(\epsilon\lambda),$
$N_{k10}:=\#\{i:X_{i}=k,A_{i}=1,Y_{i}=0\}\sim\text{Poi}\left(\lambda(\pi_{k}-\epsilon)\right),$
$N_{k0}:=\#\{i:X_{i}=k,A_{i}=0\}\sim\text{Poi}\left(\lambda(1-\pi_{k})\right)$
with $N_{k11},N_{k10},N_{k0}$ conditionally independent. Similarly under the alternative hypothesis $P^{\prime}$, let $(\pi_{1}^{\prime},\dots,\pi_{d}^{\prime})$ i.i.d. $\sim\nu_{1}$ and $\mu_{k}^{\prime}=\epsilon/\pi_{k}^{\prime}$. The sufficient statistics for $\psi_{1}$ (conditioning on $(\pi_{1}^{\prime},\dots,\pi_{d}^{\prime})$) are
$N_{k11}^{\prime}:=\#\{i:X_{i}=k,A_{i}=1,Y_{i}=1\}\sim\text{Poi}(\epsilon\lambda),$
$N_{k10}^{\prime}:=\#\{i:X_{i}=k,A_{i}=1,Y_{i}=0\}\sim\text{Poi}\left(\lambda(\pi_{k}^{\prime}-\epsilon)\right),$
$N_{k0}^{\prime}:=\#\{i:X_{i}=k,A_{i}=0\}\sim\text{Poi}\left(\lambda(1-\pi_{k}^{\prime})\right).$
Using the same notation as in proof of Theorem 2, the total variation distance between the sufficient statistics is
$\,\text{TV}(\mathbf{N}_{k},\mathbf{N}_{k}^{\prime})$
$=$ $\,\frac{1}{2}\sum_{i,j,\ell=0}^{\infty}\big{|}\mathbb{P}(N_{k11}=i,N_{k10}=j,N_{k0}=\ell)-\mathbb{P}(N_{k11}^{\prime}=i,N_{k10}^{\prime}=j,N_{k0}^{\prime}=\ell)\big{|}.$
Conditioning on $\pi_{k}$ we have
$\,\mathbb{P}(N_{k11}=i,N_{k10}=j,N_{k0}=\ell)$
$=$ $\,\mathbb{E}\left[\exp(-\epsilon\lambda)\frac{(\epsilon\lambda)^{i}}{i!}\exp(-\lambda(\pi_{k}-\epsilon))\frac{[\lambda(\pi_{k}-\epsilon)]^{j}}{j!}\right.$
$\,\left.\exp(-\lambda(1-\pi_{k}))\frac{[\lambda(1-\pi_{k})]^{\ell}}{\ell!}\right]$
$=$ $\,\mathbb{E}\left[\exp(-\lambda)\frac{(\epsilon\lambda)^{i}[\lambda(\pi_{k}-\epsilon)]^{j}[\lambda(1-\pi_{k})]^{\ell}}{i!j!\ell!}\right].$
Hence the total variation distance can be written as
$\,\text{TV}(\mathbf{N}_{k},\mathbf{N}_{k}^{\prime})$
$=$ $\,\frac{1}{2}\sum_{i,j,\ell}\left|\mathbb{E}\left[\exp(-\lambda)\frac{(\epsilon\lambda)^{i}[\lambda(\pi_{k}-\epsilon)]^{j}[\lambda(1-\pi_{k})]^{\ell}}{i!j!\ell!}\right]\right.$
$\,\left.-\mathbb{E}\left[\exp(-\lambda)\frac{(\epsilon\lambda)^{i}[\lambda(\pi_{k}^{\prime}-\epsilon)]^{j}[\lambda(1-\pi_{k}^{\prime})]^{\ell}}{i!j!\ell!}\right]\right|$
$=$ $\,\frac{1}{2}\sum_{j,\ell}\left|\mathbb{E}\left[\exp(-(1-\epsilon)\lambda)\frac{[\lambda(\pi_{k}-\epsilon)]^{j}[\lambda(1-\pi_{k})]^{\ell}}{j!\ell!}\right]\right.$
$\,\left.-\mathbb{E}\left[\exp(-(1-\epsilon)\lambda)\frac{[\lambda(\pi_{k}^{\prime}-\epsilon)]^{j}[\lambda(1-\pi_{k}^{\prime})]^{\ell}}{j!\ell!}\right]\right|,$
where the last equation follows from $\sum_{i}\exp(-\epsilon\lambda)\frac{(\epsilon\lambda)^{i}}{i!}=1$. Since $\mathbb{E}[\pi_{k}^{j}]=\mathbb{E}[\pi_{k}^{\prime j}],1\leq j\leq 2L$, we have
$\,\text{TV}(\mathbf{N}_{k},\mathbf{N}_{k}^{\prime})$
$=$ $\,\frac{1}{2}\sum_{j+\ell>2L}\left|\mathbb{E}\left[\exp(-(1-\epsilon)\lambda)\frac{[\lambda(\pi_{k}-\epsilon)]^{j}[\lambda(1-\pi_{k})]^{\ell}}{j!\ell!}\right]\right.$
$\,\left.-\mathbb{E}\left[\exp(-(1-\epsilon)\lambda)\frac{[\lambda(\pi_{k}^{\prime}-\epsilon)]^{j}[\lambda(1-\pi_{k}^{\prime})]^{\ell}}{j!\ell!}\right]\right|$
$\leq$ $\,\frac{1}{2}\left\{\sum_{j+\ell>2L}\mathbb{E}\left[\exp(-(1-\epsilon)\lambda)\frac{[\lambda(\pi_{k}-\epsilon)]^{j}[\lambda(1-\pi_{k})]^{\ell}}{j!\ell!}\right]\right.$
$\,\left.+\mathbb{E}\left[\exp(-(1-\epsilon)\lambda)\frac{[\lambda(\pi_{k}^{\prime}-\epsilon)]^{j}[\lambda(1-\pi_{k}^{\prime})]^{\ell}}{j!\ell!}\right]\right\}.$
Since given $\pi_{k}$, $N_{k11}\sim\text{Poi}(\lambda(\pi_{k}-\epsilon)),N_{k10}\sim\text{Poi}(\lambda(1-\pi_{k})),N_{k11}+N_{k10}\sim\text{Poi}(\lambda(1-\epsilon))$, we have
$\,\sum_{j+\ell>2L}\mathbb{E}\left[\exp(-(1-\epsilon)\lambda)\frac{[\lambda(\pi_{k}-\epsilon)]^{j}[\lambda(1-\pi_{k})]^{\ell}}{j!\ell!}\right]$
$=$ $\,\mathbb{E}[\mathbb{P}(N_{k11}+N_{k10}>2L|\pi_{k})]$
$\leq$ $\,\exp(-\lambda(1-\epsilon))\left(\frac{e\lambda(1-\epsilon)}{2L}\right)^{2L},$
where the last inequality follows from Chernoff bound (17). Hence in the regime $n\lesssim d^{1-\beta}$ with the choice $L=\lfloor 1/\beta\rfloor+1$, the total variation distance is bounded as
$\text{TV}(\mathbf{N}_{k},\mathbf{N}_{k}^{\prime})\leq\left(\frac{en(1-\epsilon)}{2dL}\right)^{2L}\lesssim d^{-2\beta L}=o(1/d^{2}).$
Since $(\mathbf{N}_{1},\dots,\mathbf{N}_{d})$ are independent, we have
$\text{TV}(\mathbf{N},\mathbf{N}^{\prime})\leq\sum_{k=1}^{d}\text{TV}(\mathbf{N}_{k},\mathbf{N}_{k}^{\prime})=o(1/d)\rightarrow 0.$
The functional separation between the null and the alternative is
$\psi_{1}(P)-\psi_{1}(P^{\prime})=\sum_{k=1}^{d}\frac{\epsilon}{d}\left(\frac{1}{\pi_{k}}-\frac{1}{\pi_{k}^{\prime}}\right)$
with expectation
$\mathbb{E}[\psi_{1}(P)-\psi_{1}(P^{\prime})]=\epsilon\mathbb{E}\left[\frac{1}{\pi_{k}}-\frac{1}{\pi_{k}^{\prime}}\right]=\epsilon c_{1}(\beta,\epsilon):=c_{2}(\beta,\epsilon)=c_{2}.$
Define two events
$E=\left\{\left|\frac{1}{d}\sum_{k=1}^{d}\frac{1}{\pi_{k}}-\mathbb{E}\left[\frac{1}{\pi_{k}}\right]\right|\leq\frac{c_{2}}{4}\right\},$
$E^{\prime}=\left\{\left|\frac{1}{d}\sum_{k=1}^{d}\frac{1}{\pi_{k}^{\prime}}-\mathbb{E}\left[\frac{1}{\pi_{k}^{\prime}}\right]\right|\leq\frac{c_{2}}{4}\right\}.$
By Chebyshev’s inequality, we have
$\mathbb{P}(E^{c})\leq\frac{16}{c_{2}^{2}d}\operatorname{Var}\left(\frac{1}{\pi_{k}}\right)\leq\frac{16}{\epsilon^{2}c_{2}^{2}d}.$
$\mathbb{P}(E^{\prime c})\leq\frac{16}{c_{2}^{2}d}\operatorname{Var}\left(\frac{1}{\pi_{k}^{\prime}}\right)\leq\frac{16}{\epsilon^{2}c_{2}^{2}d}.$
We put the following prior distributions induced by $\{\pi_{k},1\leq k\leq d\}$ and $\{\pi_{k}^{\prime},1\leq k\leq d\}$ on $P$ and $P^{\prime}$, respectively:
$\pi\stackrel{{\scriptstyle d}}{{=}}P\mid E,\pi^{\prime}\stackrel{{\scriptstyle d}}{{=}}P^{\prime}\mid E^{\prime}.$
Note that under $\pi,\pi^{\prime}$,
$|\psi_{1}(P)-\psi_{1}(P^{\prime})|\geq c_{2}/2.$
By triangle inequality, the total variation distance of the sufficient counting statistics $\mathbf{N}$ and $\mathbf{N}^{\prime}$ under two priors is bounded by
$\text{TV}\left({\mathbf{N}\mid E},{\mathbf{N}^{\prime}\mid E^{\prime}}\right)\leq$ $\,\text{TV}\left({\mathbf{N}\mid E},{\mathbf{N}}\right)+\text{TV}({\mathbf{N}},{\mathbf{N}^{\prime}})+\text{TV}\left({\mathbf{N}^{\prime}\mid E^{\prime}},{\mathbf{N}^{\prime}}\right)$
$\leq$ $\,\mathbb{P}(E^{c})+\mathbb{P}(E^{\prime c})+\text{TV}({\mathbf{N}},{\mathbf{N}^{\prime}})$
$\leq$ $\,\frac{32}{\epsilon^{2}c_{2}^{2}d}+\text{TV}({\mathbf{N}},{\mathbf{N}^{\prime}})\rightarrow 0.$
By method of fuzzy hypotheses, we conclude
$\,\inf_{\widehat{\psi}_{1}}\sup_{\mathbb{P}\in\mathcal{D}^{U}(\epsilon)}\mathbb{E}_{\mathbb{P}}\left[\left(\widehat{\psi}_{1}-\psi_{1}\right)^{2}\right]$
$\geq$ $\,\frac{c_{2}^{2}}{32}\left(1-\text{TV}\left({\mathbf{N}\mid E},{\mathbf{N}^{\prime}\mid E^{\prime}}\right)\right).$
$\geq$ $\,c(\beta,\epsilon),$
for some constant $c(\beta,\epsilon)$ when $d$ is large enough in the regime $n\lesssim d^{1-\beta}$. ∎
B.9 Proof of Theorem 5
Proof.
Note that $\eta=\mathbb{E}[(A-\pi_{X})(Y-\mu_{X})]$ The conditional bias is (Let $\mathbf{Z}=(X,A,Y),\mathbf{Z}_{1}=(X_{1},A_{1},Y_{1}),\mathbf{Z}_{2}=(X_{2},A_{2},Y_{2})$ be samples independent of $D$)
$\,\mathbb{E}[\widehat{\eta}\mid D]-\eta$
$=$ $\,\mathbb{E}[(A-\widehat{\pi}_{X})(Y-\widehat{\mu}_{X})\mid D]-\mathbb{E}[(A-\pi_{X})(Y-\mu_{X})\mid D]$
$\,-\mathbb{E}\left[\frac{(A_{1}-\widehat{\pi}_{X_{1}})I(X_{1}=X_{2})(Y_{2}-\widehat{\mu}(X_{2}))}{\widehat{p}_{X_{1}}}\mid D\right]$
Note that
$\,\mathbb{E}[(A-\widehat{\pi}_{X})(Y-\widehat{\mu}_{X})\mid D]-\mathbb{E}[(A-\pi_{X})(Y-\mu_{X})\mid D]$
$=$ $\,\mathbb{E}[(A-\widehat{\pi}_{X})(Y-\widehat{\mu}_{X})\mid D]-\mathbb{E}[(A-\widehat{\pi}_{X})(Y-{\mu}_{X})\mid D]$
$\,+\mathbb{E}[(A-\widehat{\pi}_{X})(Y-{\mu}_{X})\mid D]-\mathbb{E}[(A-{\pi}_{X})(Y-{\mu}_{X})\mid D]$
$=$ $\,\mathbb{E}[(A-\widehat{\pi}_{X})(\mu_{X}-\widehat{\mu}_{X})\mid D]+\mathbb{E}[(\pi_{X}-\widehat{\pi}_{X})(Y-{\mu}_{X})\mid D]$
$=$ $\,\mathbb{E}[(\pi_{X}-\widehat{\pi}_{X})(\mu_{X}-\widehat{\mu}_{X})\mid D],$
the last equation follows from conditioning on $X$. By conditioning on $X_{1}$ we have
$\mathbb{E}\left[I(X_{2}=X_{1})(Y_{2}-\widehat{\mu}(X_{2}))\mid D,X_{1}\right]=p_{X_{1}}(\mu_{X_{1}}-\widehat{\mu}_{X_{1}}).$
Hence condition on $\mathbf{Z}_{1}$ we have
$\,\mathbb{E}\left[\frac{(A_{1}-\widehat{\pi}_{X_{1}})I(X_{1}=X_{2})(Y_{2}-\widehat{\mu}(X_{2}))}{\widehat{p}_{X_{1}}}\mid D\right]$
$=$ $\,\mathbb{E}\left[\frac{p_{X_{1}}}{\widehat{p}_{X_{1}}}(A_{1}-\widehat{\pi}_{X_{1}})(\mu_{X_{1}}-\widehat{\mu}_{X_{1}})\mid D\right]$
$=$ $\,\mathbb{E}\left[\frac{p_{X_{1}}}{\widehat{p}_{X_{1}}}(\pi_{X_{1}}-\widehat{\pi}_{X_{1}})(\mu_{X_{1}}-\widehat{\mu}_{X_{1}})\mid D\right].$
Hence the conditional bias is
$\mathbb{E}[\widehat{\eta}\mid D]-\eta=\mathbb{E}\left[(\widehat{\mu}_{X}-\mu_{X})(\widehat{\pi}_{X}-\pi_{X})\left(1-\frac{p_{X}}{\widehat{p}_{X}}\right)\mid D\right].$ (21)
By Cauchy-Schwarz inequality one can bound the conditional bias as
$|\mathbb{E}[\widehat{\eta}\mid D]-\eta|\leq\|\widehat{\mu}-\mu\|_{2}\|\widehat{\pi}-\pi\|_{2}\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|$
We next bound the variance. Let
$h_{2}(\mathbf{Z}_{1},\mathbf{Z}_{2})=(A_{1}-\widehat{\pi}_{X_{1}})(Y_{1}-\widehat{\mu}_{X_{1}})-\frac{(A_{1}-\widehat{\pi}_{X_{1}})I(X_{1}=X_{2})(Y_{2}-\widehat{\mu}_{X_{2}})}{\widehat{p}_{X_{1}}}$
and note that the estimator can be expressed as
$\widehat{\eta}=\mathbb{U}_{n}[h_{2}(\mathbf{Z}_{1},\mathbf{Z}_{2})].$
We will use the variance of U-statistics (Lemma 6 in Robins et al., (2009)) to bound the conditional variance of $\widehat{\eta}$. Let
$h_{1}(\mathbf{Z}_{1})=\mathbb{E}[h_{2}(\mathbf{Z}_{1},\mathbf{Z}_{2})\mid\mathbf{Z}_{1}]=(A_{1}-\widehat{\pi}_{X_{1}})(Y_{1}-\widehat{\mu}_{X_{1}})-\frac{(A_{1}-\widehat{\pi}_{X_{1}})p_{X_{1}}(\mu_{X_{1}}-\widehat{\mu}_{X_{1}})}{\widehat{p}_{X_{1}}}.$
$\sigma_{1}^{2}=$ $\,\mathbb{E}[h_{1}^{2}(\mathbf{Z}_{1})\mid D]$
$\leq$  $\,2\left(\mathbb{E}[(A_{1}-\widehat{\pi}_{X_{1}})^{2}(Y_{1}-\widehat{\mu}_{X_{1}})^{2}\mid D]+\mathbb{E}\left[\frac{(A_{1}-\widehat{\pi}_{X_{1}})^{2}p_{X_{1}}^{2}(\mu_{X_{1}}-\widehat{\mu}_{X_{1}})^{2}}{\widehat{p}_{X_{1}}^{2}}\mid D\right]\right)$
$\lesssim$  $\,1+\mathbb{E}\left[\frac{p_{X}^{2}}{\widehat{p}_{X}^{2}}\mid D\right]$
$\lesssim$  $\,1+\left(1+\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|\right)^{2}$
$\lesssim$ $\,\left(1+\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|\right)^{2}.$
$\sigma_{2}^{2}=$ $\,\mathbb{E}[h_{2}^{2}(\mathbf{Z}_{1},\mathbf{Z}_{2})\mid D]$
$\leq$  $\,2\left(\mathbb{E}[(A_{1}-\widehat{\pi}_{X_{1}})^{2}(Y_{1}-\widehat{\mu}_{X_{1}})^{2}\mid D]+\mathbb{E}\left[\frac{(A_{1}-\widehat{\pi}_{X_{1}})^{2}I(X_{1}=X_{2})(Y_{2}-\widehat{\mu}_{X_{2}})^{2}}{\widehat{p}_{X_{1}}^{2}}\mid D\right]\right)$
$\lesssim$  $\,1+\mathbb{E}\left[\frac{I(X_{1}=X_{2})}{\widehat{p}_{X_{1}}^{2}}\mid D\right]$
$\lesssim$  $\,1+\sum_{k=1}^{d}\frac{p_{k}^{2}}{\widehat{p}_{k}^{2}}$
$\leq$  $\,1+d\left(1+\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|\right)^{2}$
$\lesssim$ $\,d\left(1+\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|\right)^{2}.$
By Lemma 6 in Robins et al., (2009) the conditional variance is upper bounded as
$\operatorname{Var}(\widehat{\eta}\mid D)\lesssim\frac{1}{n}\left(1+\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|\right)^{2}+\frac{d}{n^{2}}\left(1+\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|\right)^{2}.$
B.10 Proof of Theorem 6
Proof.
We will use the following equation for MSE
$\,\mathbb{E}[(\widehat{\eta}-\eta)^{2}]$ (22)
$=$ $\,\mathbb{E}\left\{\mathbb{E}[(\widehat{\eta}-\eta)^{2}\mid D]\right\}$
$=$ $\,\mathbb{E}[\operatorname{Var}(\widehat{\eta}\mid D)]+\mathbb{E}\{(\mathbb{E}[\widehat{\eta}\mid D]-\eta)^{2}\}.$
For the conditional bias derived in (21) we have (using the property $(\mathbb{E}[X])^{2}\leq\mathbb{E}[X^{2}]$)
$\,\mathbb{E}\{(\mathbb{E}[\widehat{\eta}\mid D]-\eta)^{2}\}$
$\leq$ $\,\mathbb{E}\left\{\mathbb{E}\left[(\widehat{\mu}_{X}-\mu_{X})^{2}(\widehat{\pi}_{X}-\pi_{X})^{2}\left(1-\frac{p_{X}}{\widehat{p}_{X}}\right)^{2}\mid D\right]\right\}$
$=$ $\,\mathbb{E}\left[\sum_{k=1}^{d}p_{k}(\widehat{\mu}_{k}-\mu_{k})^{2}(\widehat{\pi}_{k}-\pi_{k})^{2}\left(1-\frac{p_{k}}{\widehat{p}_{k}}\right)^{2}\right]$
$=$ $\,\sum_{k=1}^{d}\mathbb{E}\left[p_{k}(\widehat{\mu}_{k}-\mu_{k})^{2}(\widehat{\pi}_{k}-\pi_{k})^{2}\left(1-\frac{p_{k}}{\widehat{p}_{k}}\right)^{2}\right]$
$\leq$ $\,\xi_{n}^{2}\sum_{k=1}^{d}p_{k}\mathbb{E}[(\widehat{\pi}_{k}-\pi_{k})^{2}],$
where in the last inequality we use the bound
$\max_{k}\left|1-\frac{p_{k}}{\widehat{p}_{k}}\right|\leq\xi_{n}.$
and $(\widehat{\mu}_{k}-\mu_{k})^{2}\leq 1$. A naive bound
$\sum_{k=1}^{d}p_{k}\mathbb{E}[(\widehat{\pi}_{k}-\pi_{k})^{2}]\leq 1$
holds for both empirical average estimator $\widehat{\pi}_{k},\widehat{\mu}_{k}$ or simply letting $\widehat{\pi}_{k}=\widehat{\mu}_{k}=0$. For empirical average estimator $\widehat{\pi}_{k}$ from a sample of size $n$ we can derive an alternative bound. By the property of conditional variance, we have
$\operatorname{Var}(\widehat{\pi}_{k})=\mathbb{E}[\operatorname{Var}(\widehat{\pi}_{k}\mid\mathbf{X}^{n})]+\operatorname{Var}(\mathbb{E}[\widehat{\pi}_{k}\mid\mathbf{X}^{n}]).$
Recall $\mathbb{E}[\widehat{\pi}_{k}\mid\mathbf{X}^{n}]=\pi_{k}I(\widehat{p}_{k}>0),\operatorname{Var}(\widehat{\pi}_{k}\mid\mathbf{X}^{n})=\frac{\pi_{k}(1-\pi_{k})}{n\widehat{p}_{k}}I(\widehat{p}_{k}>0)$ we obtain
$\operatorname{Var}(\widehat{\pi}_{k})\leq\frac{1}{4}\mathbb{E}\left[\frac{I(\widehat{p}_{k}>0)}{n\widehat{p}_{k}}\right]+\pi_{k}^{2}(1-p_{k})^{n}\leq\frac{1}{2(n+1)p_{k}}+(1-p_{k})^{n},$
$(\mathbb{E}[\widehat{\pi}_{k}]-\pi_{k})^{2}=\pi_{k}^{2}(1-p_{k})^{2n}\leq(1-p_{k})^{2n}.$
Thus we have
$\mathbb{E}[(\widehat{\pi}_{k}-\pi_{k})^{2}]\leq\frac{1}{2(n+1)p_{k}}+(1-p_{k})^{n}+(1-p_{k})^{2n}$
$\sum_{k=1}^{d}p_{k}\mathbb{E}[(\widehat{\pi}_{k}-\pi_{k})^{2}]\leq\frac{d}{2(n+1)}+\sum_{k=1}^{d}p_{k}(1-p_{k})^{n}+\sum_{k=1}^{d}p_{k}(1-p_{k})^{2n}.$
Let $f_{6}(x)=x(1-x)^{n},x\in[0,1]$ and $f_{6}^{\prime}(x)=(1-nx)(1-x)^{n-1}$. Hence $f_{6}(x)\leq f_{6}(1/n)<1/n$, which implies
$\sum_{k=1}^{d}p_{k}(1-p_{k})^{n}\leq\frac{d}{n},\sum_{k=1}^{d}p_{k}(1-p_{k})^{2n}\leq\frac{d}{2n}.$
We conclude that
$\sum_{k=1}^{d}p_{k}\mathbb{E}[(\widehat{\pi}_{k}-\pi_{k})^{2}]\leq\frac{2d}{n},$
Combining this bound the naive bound
$\sum_{k=1}^{d}p_{k}\mathbb{E}[(\widehat{\pi}_{k}-\pi_{k})^{2}]\leq 1$
we have
$\mathbb{E}\{(\mathbb{E}[\widehat{\eta}\mid D]-\eta)^{2}\}\lesssim\xi_{n}^{2}\frac{n\wedge d}{n}.$
The bound on conditional variance in proof of Theorem 5 can be reduced to
$\operatorname{Var}(\widehat{\eta}\mid D)\lesssim\frac{1}{n}+\frac{d}{n^{2}}.$
The proof is completed by combining the bounds on conditional bias with variance as in (22).
B.11 Proof of Theorem 7
Proof.
We will show the plug-in estimator of $\psi_{1}=\sum_{k}p_{k}\mu_{1k}$ satisfies
$\widehat{\psi}_{1}-\psi_{1}=(\mathbb{P}_{n}-\mathbb{P})[\varphi_{1}(\mathbf{Z})]+o_{\mathbb{P}}(1/\sqrt{n}),$
where
$\varphi_{1}(\mathbf{Z})=\frac{A(Y-\mu_{1X})}{\pi_{X}}+\mu_{1X}.$
Similar argument can be applied to $\psi_{0}=\sum_{k}p_{k}\mu_{0k}$. By Proposition 1 we will use the doubly robust form of $\widehat{\psi}=\mathbb{P}_{n}[\widehat{\varphi}_{1}(\mathbf{Z})]$. Using the following decomposition
$\widehat{\psi}_{1}-\psi_{1}=(\mathbb{P}_{n}-\mathbb{P})[\varphi_{1}(\mathbf{Z})]+(\mathbb{P}_{n}-\mathbb{P})[\widehat{\varphi}_{1}(\mathbf{Z})-\varphi_{1}(\mathbf{Z})]+\mathbb{P}[\widehat{\varphi}_{1}(\mathbf{Z})-\varphi_{1}(\mathbf{Z})].$
Step1: Bound $\mathbb{P}[\widehat{\varphi}_{1}(\mathbf{Z})-\varphi_{1}(\mathbf{Z})]$. We first show $\mathbb{P}[\widehat{\varphi}_{1}(\mathbf{Z})-\varphi_{1}(\mathbf{Z})]=\int\widehat{\varphi}_{1}(\mathbf{z})-\varphi_{1}(\mathbf{z})d\mathbb{P}(\mathbf{z})=o_{\mathbb{P}}(1/\sqrt{n})$. By direct calculations, one can show
$|\mathbb{P}[\widehat{\varphi}_{1}(\mathbf{Z})-\varphi_{1}(\mathbf{Z})]|=\bigg{|}\mathbb{P}\left[(\widehat{\mu}_{1X}-\mu_{1X})\left(1-\frac{\pi_{X}}{\widehat{\pi}_{X}}\right)\right]\bigg{|}\leq\|\widehat{\mu}_{1X}-\mu_{1X}\|_{2}\left\|1-\frac{\pi_{X}}{\widehat{\pi}_{X}}\right\|_{2}.$
Note that since $n\widehat{p}_{k}\widehat{\pi}_{k}\widehat{\mu}_{1k}|\mathbf{X}^{n},\mathbf{A}^{n}\sim\text{Bin}(n\widehat{p}_{k}\widehat{\pi}_{k},\mu_{1k})$, we have
$\mathbb{E}[\widehat{\mu}_{1k}\mid\mathbf{X}^{n},\mathbf{A}^{n}]=\mu_{1k}I(\widehat{p}_{k}\widehat{\pi}_{k}>0),\,\operatorname{Var}(\widehat{\mu}_{1k}\mid\mathbf{X}^{n},\mathbf{A}^{n})=\frac{\mu_{1k}(1-\mu_{1k})I(\widehat{p}_{k}\widehat{\pi}_{k}>0)}{n\widehat{p}_{k}\widehat{\pi}_{k}}.$
The bias of $\widehat{\mu}_{1k}$ is
$\mathbb{E}[\widehat{\mu}_{1k}-\mu_{1k}]=-\mu_{1k}\mathbb{P}(\widehat{p}_{k}\widehat{\pi}_{k}=0).$
By conditioning on $\mathbf{X}^{n}$, we have
$\mathbb{P}(\widehat{p}_{k}\widehat{\pi}_{k}=0)=\mathbb{E}[\mathbb{P}(\widehat{p}_{k}\widehat{\pi}_{k}=0\mid\mathbf{X}^{n})]=\mathbb{E}[(1-\pi_{k})^{n\widehat{p}_{k}}]=(1-p_{k}\pi_{k})^{n}\leq(1-\epsilon p_{k})^{n},$
where we use the fact $\mathbb{E}[c^{V}]=(1-p+pc)^{n}$ for $V\sim B(n,p)$. Similarly, we have
$\operatorname{Var}(\mathbb{E}[\widehat{\mu}_{1k}\mid\mathbf{X}^{n},\mathbf{A}^{n}])=\mu_{1k}^{2}\mathbb{P}(\widehat{p}_{k}\widehat{\pi}_{k}=0)\mathbb{P}(\widehat{p}_{k}\widehat{\pi}_{k}>0)\leq(1-\epsilon p_{k})^{n}.$
For the expected conditional variance,
$\mathbb{E}\left[\operatorname{Var}(\widehat{\mu}_{1k}\mid\mathbf{X}^{n},\mathbf{A}^{n})\right]\leq\frac{1}{4}\mathbb{E}\left[\frac{I(\widehat{p}_{k}\widehat{\pi}_{k}>0)}{n\widehat{p}_{k}\widehat{\pi}_{k}}\right]$
In the following derivation, we need the next lemma.
Lemma 7.
(Lemma A.2 in Devroye et al., (2013))If $X\sim B(n,p)$, then
$\mathbb{E}\left\{\frac{I(X>0)}{X}\right\}\leq\frac{2}{(n+1)p}$
Note that $n\widehat{p}_{k}\widehat{\pi}_{k}\sim B(n,p_{k}\pi_{k})$, we have
$\mathbb{E}\left[\operatorname{Var}(\widehat{\mu}_{1k}\mid\mathbf{X}^{n},\mathbf{A}^{n})\right]\leq\frac{1}{2(n+1)p_{k}\pi_{k}}\leq\frac{1}{2\epsilon(n+1)p_{k}}.$
Thus the variance of $\widehat{\mu}_{1k}$ is bounded by
$\operatorname{Var}(\widehat{\mu}_{1k})\leq(1-\epsilon p_{k})^{n}+\frac{1}{2\epsilon(n+1)p_{k}}.$
We conclude
$\mathbb{E}[(\widehat{\mu}_{1k}-\mu_{1k})^{2}]\leq(1-\epsilon p_{k})^{n}+\frac{1}{2\epsilon(n+1)p_{k}}+(1-\epsilon p_{k})^{2n}=O(1/n)$
since when $d$ is fixed, $p_{k}$’s are considered as fixed. Hence
$\|\widehat{\mu}_{1X}-\mu_{1X}\|_{2}^{2}=\sum_{k}p_{k}(\widehat{\mu}_{1k}-\mu_{1k})^{2}=O_{\mathbb{P}}(1/n),$
$\|\widehat{\mu}_{1X}-\mu_{1X}\|_{2}=O_{\mathbb{P}}(1/\sqrt{n}).$
Similarly $\mathbb{E}[\widehat{\pi}_{k}\mid\mathbf{X}^{n}]=\pi_{k}I(\widehat{p}_{k}>0),\operatorname{Var}(\widehat{\pi}_{k}\mid\mathbf{X}^{n})=\frac{\pi_{k}(1-\pi_{k})}{n\widehat{p}_{k}}I(\widehat{p}_{k}>0)$, we obtain
$\operatorname{Var}(\widehat{\pi}_{k})\leq\frac{1}{4}\mathbb{E}\left[\frac{I(\widehat{p}_{k}>0)}{n\widehat{p}_{k}}\right]+\pi_{k}^{2}(1-p_{k})^{n}\leq\frac{1}{2(n+1)p_{k}}+(1-p_{k})^{n},$
$(\mathbb{E}[\widehat{\pi}_{k}]-\pi_{k})^{2}=\pi_{k}^{2}(1-p_{k})^{2n}\leq(1-p_{k})^{2n}.$
Thus we have
$\mathbb{E}[(\widehat{\pi}_{k}-\pi_{k})^{2}]\leq\frac{1}{2(n+1)p_{k}}+(1-p_{k})^{n}+(1-p_{k})^{2n}=O(1/n)$
$(\widehat{\pi}_{k}-\pi_{k})^{2}=O_{\mathbb{P}}(1/n).$
Since $\pi_{k}\geq\epsilon$, this shows $1/\widehat{\pi}_{k}=O_{\mathbb{P}}(1)$ and
$\frac{(\widehat{\pi}_{k}-\pi_{k})^{2}}{\widehat{\pi}_{k}^{2}}=O_{\mathbb{P}}(1/n).$
We conclude
$\left\|1-\frac{\pi_{X}}{\widehat{\pi}_{X}}\right\|_{2}^{2}=\sum_{k}p_{k}\frac{(\widehat{\pi}_{k}-\pi_{k})^{2}}{\widehat{\pi}_{k}^{2}}=O_{\mathbb{P}}(1/n),$
$\left\|1-\frac{\pi_{X}}{\widehat{\pi}_{X}}\right\|_{2}=O_{\mathbb{P}}(1/\sqrt{n}).$
This shows
$\mathbb{P}[\widehat{\varphi}_{1}(\mathbf{Z})-\varphi_{1}(\mathbf{Z})]=O_{\mathbb{P}}(1/n)=o_{\mathbb{P}}(1/\sqrt{n}).$
Step2: Bound $(\mathbb{P}_{n}-\mathbb{P})[\widehat{\varphi}_{1}(\mathbf{Z})-\varphi_{1}(\mathbf{Z})]$.We then show
$(\mathbb{P}_{n}-\mathbb{P})[\widehat{\varphi}_{1}(\mathbf{Z})-\varphi_{1}(\mathbf{Z})]=o_{\mathbb{P}}(1/\sqrt{n}).$
Since $X$ is discrete we can write the nuisance functions $(\widehat{\pi},\widehat{\mu}_{1})$ as saturated linear models, i.e.
${\pi}_{x}=\pi(x;\boldsymbol{\alpha})=\boldsymbol{\alpha}^{\top}\boldsymbol{w},$
where $\boldsymbol{w}^{\top}=\{I(x=1),\ldots,I(x=d)\}\in\{0,1\}^{d}$ and
${\mu}_{1x}=\mu_{1}(x;\boldsymbol{\beta})=\boldsymbol{\beta}^{\top}\boldsymbol{w}.$
Here the components of $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$ are simply propensity scores and regression functions within different categories and $\|\boldsymbol{w}\|_{2}=1$. Define the function class =$\mathcal{F}=\{\varphi_{1}(\mathbf{z};\boldsymbol{\gamma}),\boldsymbol{\gamma}=(\boldsymbol{\alpha},\boldsymbol{\beta}),\alpha_{k}\in[\epsilon/2,1],\beta_{k}\in[0,1]\}$. Since propensity scores are lower bounded by $\epsilon/2$ for $\boldsymbol{\alpha}\in\mathcal{F}$, one can show there is a constant $K$ that depends on $\epsilon$ such that
$|\varphi_{1}(\mathbf{z};\boldsymbol{\gamma}_{1})-\varphi_{1}(\mathbf{z};\boldsymbol{\gamma}_{2})|\leq K\|\boldsymbol{\gamma}_{1}-\boldsymbol{\gamma}_{2}\|_{2}.$
Since $d$ is fixed as constant, by example 19.7 in Van der Vaart, (2000), $\mathcal{F}$ is Donsker. Let
$\widetilde{\pi}_{k}=\widehat{\pi}_{k}I(\widehat{\pi}_{k}\geq\epsilon/2)+\frac{\epsilon}{2}I(\widehat{\pi}_{k}<\epsilon/2),\,\widetilde{\mu}_{1k}=\widehat{\mu}_{1k},$
so that we truncate the estimated propensity score $\widehat{\pi}_{k}$ to obtain $\widetilde{\pi}_{k}$. Let
$\widetilde{\varphi}_{1}(\mathbf{Z})=\varphi_{1}(\mathbf{Z};\widetilde{\boldsymbol{\gamma}})=\frac{A(Y-\widetilde{\mu}_{1X})}{\widetilde{\pi}_{X}}+\widetilde{\mu}_{1X}.$
Clearly $\widetilde{\varphi}_{1}(\cdot)=\varphi_{1}(\cdot,\widetilde{\boldsymbol{\gamma}})\in\mathcal{F}$. We have
$\|\widetilde{\boldsymbol{\gamma}}-\boldsymbol{\gamma}\|_{2}^{2}=\|\widetilde{\boldsymbol{\alpha}}-\boldsymbol{\alpha}\|_{2}^{2}+\|\widetilde{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{2}^{2}=\sum_{k=1}^{d}\left[(\widetilde{\pi}_{k}-\pi_{k})^{2}+(\widetilde{\mu}_{1k}-\mu_{1k})^{2}\right].$
Since we assume $\pi_{k}\geq\epsilon$, truncating $\widehat{\pi}_{k}$ yields smaller error and hence $|\widetilde{\pi}_{k}-\pi_{k}|\leq|\widehat{\pi}_{k}-\pi_{k}|$. By the consistency of $\widehat{\pi}_{k}$ and $\widehat{\mu}_{1k}$ shown above, we have
$\|\widetilde{\boldsymbol{\gamma}}-\boldsymbol{\gamma}\|_{2}^{2}\leq\sum_{k=1}^{d}\left[(\widehat{\pi}_{k}-\pi_{k})^{2}+(\widehat{\mu}_{1k}-\mu_{1k})^{2}\right]=o_{\mathbb{P}}(1)$
again because $d$ is fixed. Thus
$\|\widetilde{\varphi}_{1}-\varphi\|_{2}\leq K\|\widetilde{\boldsymbol{\gamma}}-\boldsymbol{\gamma}\|=o_{\mathbb{P}}(1)$
and Lemma 19.24 in Van der Vaart, (2000) shows
$(\mathbb{P}_{n}-\mathbb{P})[\widetilde{\varphi}_{1}(\mathbf{Z})-\varphi_{1}(\mathbf{Z})]=o_{\mathbb{P}}(1/\sqrt{n}).$ (23)
Now consider
$(\mathbb{P}_{n}-\mathbb{P})[\widehat{\varphi}_{1}(\mathbf{Z})-\widetilde{\varphi}_{1}(\mathbf{Z})].$
By strong law of large numbers, we have
$\widehat{\pi}_{k}=\frac{\sum_{i=1}^{n}I(X_{i}=k,A_{i}=1)}{\sum_{i=1}^{n}I(X_{i}=k)}\rightarrow\pi_{k}$
for any $k\in[d]$ almost surely. Thus for almost every $\omega\in\Omega$ (sample space), one can find a $N(\omega,\epsilon,d)\in\mathbf{N}^{+}$ such that for all $n\geq N(\omega,\epsilon,d)$, we have $|\widehat{\pi}_{k}-\pi_{k}|\leq\epsilon/2$ for any $k\in[d]$. This together with $\pi_{k}\geq\epsilon$ shows $\widehat{\pi}_{k}\geq\epsilon/2$ and hence $\widehat{\pi}_{k}=\widetilde{\pi}_{k}$ holds for all $k$ when $n\geq N(\omega,\epsilon,d)$. This implies
$\widetilde{\varphi}_{1}(\mathbf{z})\equiv\widehat{\varphi}_{1}(\mathbf{z}),$
$(\mathbb{P}_{n}-\mathbb{P})[\widehat{\varphi}_{1}(\mathbf{Z})-\widetilde{\varphi}_{1}(\mathbf{Z})]=0$
when $n\geq N(\omega,\epsilon,d)$. This clearly implies
$\sqrt{n}(\mathbb{P}_{n}-\mathbb{P})[\widehat{\varphi}_{1}(\mathbf{Z})-\widetilde{\varphi}_{1}(\mathbf{Z})]\rightarrow 0$
almost surely since the left-hand side is exactly 00 when $n$ is large enough. We conclude
$(\mathbb{P}_{n}-\mathbb{P})[\widehat{\varphi}_{1}(\mathbf{Z})-\widetilde{\varphi}_{1}(\mathbf{Z})]=o_{\mathbb{P}}(1/\sqrt{n})$ (24)
Combining (23) and (24), we have
$(\mathbb{P}_{n}-\mathbb{P})[\widehat{\varphi}_{1}(\mathbf{Z})-\varphi_{1}(\mathbf{Z})]=o_{\mathbb{P}}(1/\sqrt{n}).$
Appendix C Proof of Auxiliary Results
C.1 Proof of Lemma 2
Proof.
First consider the boundary. If $y=0$ then the function $f_{2}$ is reduced to
$g_{1}(x)=(n-1)(1-x)^{n-2}-n(1-x)^{n-1},0\leq x\leq 1,$
$g_{1}^{\prime}(x)=(n-1)(1-x)^{n-3}(2-nx).$
So $|g_{1}(x)|\leq\max(|g_{1}(0)|,|g_{1}(1)|,|g_{1}(2/n)|)$ and $g_{1}(0)=-1,g_{1}(1)=0,g_{1}(2/n)=\left(\frac{n-2}{n}\right)^{n-2}\leq 1$. This shows $|g_{1}(x)|\leq 1$. On the boundary $x=0$ similar arguments hold. Now consider the boundary $x+y=1$, the function $f_{2}$ is reduced to
$|f_{2}(x,y)|=nx^{n-1}(1-x)^{n-1}\leq\frac{n}{4^{n-1}}\leq 1$
Now we consider the interior of the triangle. By taking the partial derivative (or by noting that $x,y$ have symmetric roles in the function $f_{2}$) we see the maximizer of $|f_{2}(x,y)|$ must lie on the line $x=y$. So we define
$g_{2}(x)=(n-1)(1-2x)^{n-2}-n(1-x)^{2n-2},0\leq x\leq 1/2$
and only need to show $|g_{2}(x)|\leq 1$. Let $X_{0}$ be the set of stationary points of $g_{2}$ on $[0,1/2]$. Any stationary point $x_{0}\in X_{0}$ must satisfy $g_{2}^{\prime}(x_{0})=0$, i.e.
$n(1-x_{0})^{2n-3}=(n-2)(1-2x_{0})^{n-3}.$
So for any $x_{0}\in X_{0}$,
$g_{2}(x_{0})=(n-1)(1-2x_{0})^{n-2}-(n-2)(1-2x_{0})^{n-3}(1-x_{0})=(1-2x_{0})^{n-3}(1-nx_{0}).$
We then define a new function to show $|g_{2}(x_{0})|\leq 1$ for any $x_{0}\in X_{0}$. Let
$g_{3}(x)=(1-2x)^{n-3}(1-nx),0\leq x\leq 1/2$
$g_{3}^{\prime}(x)=(n-2)(2nx-3)(1-2x)^{n-4}.$
So we have $|g_{3}(x)|\leq\max(|g_{3}(0)|,|g_{3}(1/2)|,|g_{3}(3/2n)|)$ and $g_{3}(0)=1,g_{3}(1/2)=0,|g_{3}(3/2n)|=|\frac{1}{2}(\frac{n-3}{n})^{n-3}|\leq 1/2$. This shows $|g_{3}(x)|\leq 1$ on $[0,1/2]$, which implies $|g_{2}(x_{0})|\leq 1$ for any $x\in X_{0}$. Note that $g_{2}(0)=-1,g_{2}(1/2)=-n/4^{n-1}$ We conclude
$|g_{2}(x)|\leq\max\left(\max_{x_{0}\in X_{0}}|g_{2}(x)|,1,n/4^{n-1}\right)\leq 1$
C.2 Proof of Lemma 3
Proof.
For any $\gamma>0$, let $\widehat{\psi}_{1}(n)$ be a near-minimax optimal estimator of $\psi_{1}({\boldsymbol{p}})$ for fixed sample $n$, i.e.,
$\sup_{\mathbb{P}\in\mathcal{D}(\epsilon)}\mathbb{E}_{\mathbb{P}}\left[\left(\widehat{\psi}_{1}(n)-\psi_{1}({\boldsymbol{p}})\right)^{2}\right]\leq\gamma+R^{*}(d,n).$
Note we emphasize the dependency of $\psi_{1}$ on ${\boldsymbol{p}}=(p_{1},\dots)$. Under a Poisson-sampling model $\mathbb{P}\in\mathcal{D}(\epsilon,\delta)$, let $n^{\prime}=\sum_{k}n\widehat{p}_{k}\sim\text{Poi}(n\sum_{k}p_{k})$ and construct an estimator $\widetilde{\psi}_{1}=\widehat{\psi}_{1}(n^{\prime})$. Note that conditioned on $n^{\prime}=m,(n\widehat{p}_{1},\dots,)\sim\text{Multinomial}\left(m,\frac{{\boldsymbol{p}}}{\sum_{k}p_{k}}\right)$. By definition, the model with probability vector $\frac{{\boldsymbol{p}}}{\sum_{k}p_{k}}$ and the same propensity score, regression functions as $\mathbb{P}$ is in $\mathcal{D}(\epsilon)$. Under the Poisson-sampling model we have
$\,\mathbb{E}\left[\left(\widetilde{\psi}_{1}-\psi_{1}\left(\frac{{\boldsymbol{p}}}{\sum_{k}p_{k}}\right)\right)^{2}\right]$
$=$ $\,\sum_{m=0}^{\infty}\mathbb{E}\left[\left(\widetilde{\psi}_{1}-\psi_{1}\left(\frac{{\boldsymbol{p}}}{\sum_{k}p_{k}}\right)\right)^{2}\mid n^{\prime}=m\right]\mathbb{P}(n^{\prime}=m)$
$\leq$ $\,\sum_{m=0}^{\infty}R^{*}(d,m)\mathbb{P}(n^{\prime}=m)+\gamma.$
Since $n\mapsto R^{*}(d,n)$ is non-increasing for fixed $d$ and $R^{*}(d,n)\leq 1$ , we have
$\,\mathbb{E}\left[\left(\widetilde{\psi}_{1}-\psi_{1}\left(\frac{{\boldsymbol{p}}}{\sum_{k}p_{k}}\right)\right)^{2}\right]$
$\leq$ $\,\sum_{m\geq n/2}R^{*}(d,m)\mathbb{P}\left[n^{\prime}=m\right]+\mathbb{P}\left[n^{\prime}\leq\frac{n}{2}\right]+\gamma$
$\leq$ $\,R^{*}(d,n/2)+\exp(-n/50)+\gamma.$
The last inequality follows from Chernoff bound and $|\sum_{k}p_{k}-1|\leq\delta<1/3$. The difference between $\psi_{1}\left(\frac{{\boldsymbol{p}}}{\sum_{k}p_{k}}\right)$ and $\psi_{1}\left({{\boldsymbol{p}}}\right)$ is
$\,\bigg{|}\psi_{1}\left(\frac{{\boldsymbol{p}}}{\sum_{k}p_{k}}\right)-\psi_{1}\left({{\boldsymbol{p}}}\right)\bigg{|}$
$=$ $\,\big{|}\sum_{k}p_{k}-1\big{|}\frac{\sum_{k}p_{k}\mu_{1k}}{\sum_{k}p_{k}}$
$\leq$ $\,\delta.$
The last inequality follows since $\frac{\sum_{k}p_{k}\mu_{1k}}{\sum_{k}p_{k}}$ is a weighted average of quantities bounded by 1. Thus we have
$\,\frac{1}{2}\mathbb{E}\left[\left(\widetilde{\psi}_{1}-\psi_{1}\left({{\boldsymbol{p}}}\right)\right)^{2}\right]$
$\leq$ $\,\mathbb{E}\left[\left(\widetilde{\psi}_{1}-\psi_{1}\left(\frac{{\boldsymbol{p}}}{\sum_{k}p_{k}}\right)\right)^{2}\right]+\left(\psi_{1}\left(\frac{{\boldsymbol{p}}}{\sum_{k}p_{k}}\right)-\psi_{1}\left({{\boldsymbol{p}}}\right)\right)^{2}$
$\leq$ $\,R^{*}(d,n/2)+\exp(-n/50)+\gamma+\delta^{2}.$
Take supremum over $\mathbb{P}\in D(\epsilon,\delta)$ and since $\gamma$ is arbitrary, we have
$\frac{1}{2}\widetilde{R}^{*}(d,n,\delta)\leq R^{*}(d,n/2)+\exp(-n/50)+\delta^{2}.$
C.3 Proof of Lemma 4
Proof.
We claim it suffices to find probability measures $\omega_{0},\omega_{1}$ on $[c/K^{2},1]$ such that
  1. 1.
    $\mathbb{E}_{\omega_{0}}[X^{l}]=\mathbb{E}_{\omega_{1}}[X^{l}]\text{ for all }l=-1,0,1,\dots,3K.$ (25)
  2. 2.
    $\left|\mathbb{E}_{\omega_{0}}\left[\frac{X}{X+c/K^{2}}\right]-\mathbb{E}_{\omega_{1}}\left[\frac{X}{X+c/K^{2}}\right]\right|\geq c^{\prime}.$
Note that here we use $X$ to denote a different random variable from the covariate in the maintexts. We first show the claim leads to the results in Lemma 4. With $\omega_{i}$ ($i=0,1$) we define $\widetilde{\omega}_{i}$ supported on $\{0\}\cup[c/K^{2},1]$ such that
$\widetilde{\omega}_{i}(dx)=\frac{c}{K^{2}x}\omega_{i}(dx)+\left(1-\mathbb{E}\left[\frac{c}{K^{2}X}\right]\right)\delta_{0}(dx).$
And for $X\sim\widetilde{\omega}_{i}$, let $\nu_{i}$ be the distribution of $\frac{c_{1}\log n}{n}X$. We let $p\sim\nu_{i}$,
$\pi=\left\{\begin{array}[]{rl}\epsilon\left(1+\frac{c_{1}\log n}{np}\frac{c}{K^{2}}\right)&\text{if }p>0,\\ \epsilon&\text{if }p=0\end{array}\right.$
and $\mu=\epsilon/\pi$. Note that $\pi,\mu$ are both functions of $p$. We then verify the properties in Lemma 4 with joint distribution $\mu_{i}$ defined above.
  1. 1.
    $\text{supp}(\nu_{i})\subseteq\{0\}\cup\frac{c_{1}\log n}{n}[c/K^{2},1]$ implies the range of $p$.
  2. 2.
    For $i,j,k\geq 0$ and $1\leq i+j+k\leq 3K$ we have
    $\,\mathbb{E}_{\mu_{0}}[p^{i}(p\pi)^{j}(p\pi\mu)^{k}]$
    $=$ $\,\mathbb{E}_{\mu_{0}}[p^{i}(p\pi)^{j}(p\pi\mu)^{k}I(p>0)]$
    $=$ $\,\mathbb{E}_{\nu_{0}}\left[p^{i+k}\epsilon^{j+k}\left(p+\frac{cc_{1}\log n}{nK^{2}}\right)^{j}I(p>0)\right]$
    $=$ $\,\mathbb{E}_{\widetilde{\omega}_{0}}\left[\left(\frac{c_{1}\log n}{n}\right)^{i+j+k}\epsilon^{j+k}X^{i+k}\left(X+\frac{c}{K^{2}}\right)^{j}I(X>0)\right]$
    $=$ $\,\int\left(\frac{c_{1}\log n}{n}\right)^{i+j+k}\epsilon^{j+k}\frac{c}{K^{2}}x^{i+k-1}\left(x+\frac{c}{K^{2}}\right)^{j}\omega_{0}(dx)$
    $=$ $\,\int\left(\frac{c_{1}\log n}{n}\right)^{i+j+k}\epsilon^{j+k}\frac{c}{K^{2}}x^{i+k-1}\left(x+\frac{c}{K^{2}}\right)^{j}\omega_{1}(dx)$
    $=$ $\,\mathbb{E}_{\mu_{1}}[p^{i}(p\pi)^{j}(p\pi\mu)^{k}].$
    The first four equations follow from the definition of distributions above and the fifth follows from (25).
  3. 3.
    $\,\mathbb{E}_{\mu_{i}}[p]$
    $=$ $\,\int_{p>0}p\nu_{i}(dp)$
    $=$ $\,\int_{x>0}\frac{c_{1}\log n}{n}x\widetilde{\omega}_{i}(dx)$
    $=$ $\,\int_{x>0}\frac{c_{1}\log n}{n}\frac{c}{K^{2}}\widetilde{\omega}_{i}(dx)$
    $=$ $\,\frac{cc_{1}}{c_{2}^{2}}\frac{1}{n\log n}\leq\frac{cc_{1}c_{3}}{c_{2}^{2}d}\leq\frac{1}{d}$
    as long as $cc_{1}c_{3}/c_{2}^{2}\leq 1$.
  4. 4.
    $\,|\mathbb{E}_{\mu_{0}}[p\mu]-\mathbb{E}_{\mu_{1}}[p\mu]|$
    $=$ $\,\left|\mathbb{E}_{\nu_{0}}\left[\frac{p^{2}}{p+\frac{c_{1}\log n}{n}\frac{c}{K^{2}}}\right]-\mathbb{E}_{\nu_{1}}\left[\frac{p^{2}}{p+\frac{c_{1}\log n}{n}\frac{c}{K^{2}}}\right]\right|$
    $=$ $\,\frac{c_{1}\log n}{n}\left|\mathbb{E}_{\widetilde{\omega}_{0}}\left[\frac{X^{2}}{X+\frac{c}{K^{2}}}\right]-\mathbb{E}_{\widetilde{\omega}_{1}}\left[\frac{X^{2}}{X+\frac{c}{K^{2}}}\right]\right|$
    $=$ $\,\frac{cc_{1}\log n}{nK^{2}}\left|\mathbb{E}_{{\omega}_{0}}\left[\frac{X}{X+\frac{c}{K^{2}}}\right]-\mathbb{E}_{{\omega}_{1}}\left[\frac{X}{X+\frac{c}{K^{2}}}\right]\right|$
    $\geq$ $\,\frac{cc^{\prime}c_{1}}{c_{2}^{2}n\log n}.$
    And we define $c_{4}=cc^{\prime}c_{1}/c_{2}^{2}$.
Thus we only need to prove the claim. By duality of polynomial approximation and moment matching (see e.g., Lemma 19 in Han et al., (2020)) it suffices to prove
$\inf_{P\in\text{span}\{1/x,1,x,\dots,x^{3K}\}}\max_{x\in[c/K^{2},1]}\left|\frac{x}{x+c/K^{2}}-P(x)\right|\geq c^{\prime}/2.$
Let
$E_{n}(f;S)=\min_{\text{deg}(P)\leq n}\max_{x\in S}|f(x)-P(x)|$ (26)
be the best polynomial approximation error (with polynomial’s degree smaller than $n$) of $f$ on a set $S$. We have the following lemma.
Lemma 8.
There exists $c_{0},c^{\prime}>0$ such that
$\liminf_{n\rightarrow\infty}\inf_{\alpha\in\mathbb{R}}E_{n}\left(\frac{x}{x+c_{0}n^{-2}}+\frac{\alpha}{x};[c_{0}n^{-2},1]\right)\geq c^{\prime}>0.$
Let $c=c_{0}/9$, we then have for any $P\in\text{span}\{1/x,1,x,\dots,x^{3K}\}$, write $P(x)=\frac{\alpha_{-1}}{x}+\sum_{k=0}^{3K}\alpha_{k}x^{k}$ and let $P_{\geq}(x)=\sum_{k=0}^{3K}\alpha_{k}x^{k}$
$\,\max_{x\in[c/K^{2},1]}\left|\frac{x}{x+c/K^{2}}-P(x)\right|$
$\geq$ $\,\inf_{\alpha\in\mathbb{R}}\max_{x\in[c/K^{2},1]}\left|\frac{x}{x+c/K^{2}}+\frac{\alpha}{x}-P_{\geq}(x)\right|$
$=$ $\,\inf_{\alpha\in\mathbb{R}}\max_{x\in[c_{0}/(3K)^{2},1]}\left|\frac{x}{x+c_{0}/(3K)^{2}}+\frac{\alpha}{x}-P_{\geq}(x)\right|$
$\geq$ $\,\inf_{\alpha\in\mathbb{R}}E_{3K}\left(\frac{x}{x+c_{0}/(3K)^{2}}+\frac{\alpha}{x};[c_{0}/(3K)^{2},1]\right)$
$\geq$ $\,\frac{c^{\prime}}{2},$
as $K=c_{2}\log n$ is large enough.
We then prove Lemma 8. By Theorem 7.2.4 in Ditzian and Totik, (2012) (and use the notation there), we have
$\omega_{\varphi}^{1}\left(f,\frac{1}{n}\right)_{\infty}\leq\frac{M}{n}\sum_{\ell=0}^{n}E_{\ell}(f;[0,1]),$ (27)
where
$\omega_{\varphi}^{1}\left(f,\frac{1}{n}\right)_{\infty}=\sup_{0<h\leq 1/n}\sup_{x,x+h\varphi(x)\in[0,1]}|f(x+h\varphi(x))-f(x)|,\,\varphi(x)=\sqrt{x(1-x)}$
and $M$ is an absolute constant. Let $f_{\alpha}(x)=\frac{x}{x+c_{0}n^{-2}}+\frac{\alpha}{x}$ and $\widetilde{f}_{\alpha}(x)=f_{\alpha}\left(c_{0}n^{-2}+(1-c_{0}n^{-2})x\right),x\in[0,1]$. Denote $m=\lceil n/\sqrt{c_{0}}\rceil$. Here $c_{0}$ is a constant to be chosen. Consider two cases of $\alpha$ separately:
Case 1: $|\alpha|\leq c_{0}/n^{2}$. By (27) we have
$\,E_{n}(f_{\alpha};[c_{0}n^{-2},1])$
$=$ $\,E_{n}(\widetilde{f}_{\alpha};[0,1])$
$\geq$ $\,\frac{1}{m-n+1}\sum_{\ell=n}^{m}\mathbb{E}_{\ell}(\widetilde{f}_{\alpha};[0,1])$
$\geq$ $\,\frac{1}{m}\sum_{\ell=n}^{m}\mathbb{E}_{\ell}\left(\widetilde{f}_{\alpha};[0,1]\right)$
$\geq$ $\,\frac{1}{M}\omega_{\varphi}^{1}\left(\widetilde{f}_{\alpha},\frac{1}{m}\right)_{\infty}-\frac{1}{m}\sum_{\ell=0}^{n-1}\mathbb{E}_{\ell}\left(\widetilde{f}_{\alpha};[0,1]\right)$
$\geq$ $\,\frac{1}{M}\omega_{\varphi}^{1}\left(\widetilde{f}_{\alpha},\frac{1}{m}\right)_{\infty}-\frac{n}{m}\mathbb{E}_{0}\left(\widetilde{f}_{\alpha};[0,1]\right)$
$\geq$ $\,\frac{1}{M}\omega_{\varphi}^{1}\left(\widetilde{f}_{\alpha},\frac{1}{m}\right)_{\infty}-2\sqrt{c_{0}},$
where the first and the fourth inequality follow from the monotonity of $E_{n}$, the third inequality applies (27) and the last one follows from $|\widetilde{f}_{\alpha}|\leq 2$. For the first term, we have
$\,\omega_{\varphi}^{1}\left(\widetilde{f}_{\alpha},\frac{1}{m}\right)_{\infty}$
$\geq$ $\,\sup_{t_{1},t_{2}\in[3,4]}\left|\widetilde{f}_{\alpha}\left(\frac{t_{1}}{m^{2}}\right)-\widetilde{f}_{\alpha}\left(\frac{t_{2}}{m^{2}}\right)\right|$
$=$ $\,\sup_{t_{1},t_{2}\in[3,4]}\left|{f}_{\alpha}\left(\frac{c_{0}}{n^{2}}+\left(1-\frac{c_{0}}{n^{2}}\right)\frac{t_{1}}{m^{2}}\right)-{f}_{\alpha}\left(\frac{c_{0}}{n^{2}}+\left(1-\frac{c_{0}}{n^{2}}\right)\frac{t_{2}}{m^{2}}\right)\right|,$
where the inequality follows from the fact that when $m$ is sufficiently large,
$\frac{t_{2}-t_{1}}{m^{2}}\leq\frac{1}{m}\sqrt{\frac{t_{2}}{m}\left(1-\frac{t_{2}}{m}\right)}$
holds for all $t_{1}<t_{2}$ and $t_{1},t_{2}\in[3,4]$. Note that as $m\rightarrow\infty$
$1+\left(\frac{n^{2}}{c_{0}}-1\right)\frac{3}{m^{2}}\rightarrow 4,$
$1+\left(\frac{n^{2}}{c_{0}}-1\right)\frac{4}{m^{2}}\rightarrow 5$
Thus we have
$\,\sup_{t_{1},t_{2}\in[3,4]}\left|{f}_{\alpha}\left(\frac{c_{0}}{n^{2}}+\left(1-\frac{c_{0}}{n^{2}}\right)\frac{t_{1}}{m^{2}}\right)-{f}_{\alpha}\left(\frac{c_{0}}{n^{2}}+\left(1-\frac{c_{0}}{n^{2}}\right)\frac{t_{2}}{m^{2}}\right)\right|$
$\geq$ $\,\sup_{t_{1},t_{2}\in[3.5,5]}\left|{f}_{\alpha}\left(\frac{c_{0}}{n^{2}}t_{1}\right)-{f}_{\alpha}\left(\frac{c_{0}}{n^{2}}t_{2}\right)\right|$
$\geq$ $\,\sup_{t\in[3.5,4]}\left|{f}_{\alpha}\left(\frac{c_{0}}{n^{2}}\left(t+\frac{1}{2}\right)\right)-{f}_{\alpha}\left(\frac{c_{0}}{n^{2}}t\right)\right|$
$=$ $\,\sup_{t\in[3.5,4]}\left|\frac{t+1/2}{t+3/2}-\frac{t}{t+1}+\frac{\alpha n^{2}}{c_{0}}\left(\frac{1}{t+1/2}-\frac{1}{t}\right)\right|$
$\geq$ $\,\inf_{\beta}\sup_{t\in[3.5,4]}\left|\frac{t+1/2}{t+3/2}-\frac{t}{t+1}+\beta\left(\frac{1}{t+1/2}-\frac{1}{t}\right)\right|$
$\geq$ $\,a_{1},$
where $a_{1}$ is a positive constant independent of $\alpha,n,c_{0}$ since $\frac{t+1/2}{t+3/2}-\frac{t}{t+1}$ and $\frac{1}{t+1/2}-\frac{1}{t}$ are linearly independent. Hence we have
$E_{n}(f_{\alpha};[c_{0}n^{-2},1])\geq\frac{a_{1}}{M}-2\sqrt{c_{0}},\,|\alpha|\leq\frac{c_{0}}{n^{2}}.$
Case 2: $|\alpha|>{c_{0}}/{n^{2}}$, similar to the proof in Case 1 we have
$\,E_{n}(f_{\alpha};[c_{0}n^{-2},1])$
$\geq$ $\,\frac{1}{M}\omega_{\varphi}^{1}\left(\widetilde{f}_{\alpha},\frac{1}{m}\right)_{\infty}-\sqrt{c_{o}}\mathbb{E}_{0}\left(\widetilde{f}_{\alpha};[0,1]\right)$
$\geq$ $\,\frac{1}{M}\omega_{\varphi}^{1}\left(\widetilde{f}_{\alpha},\frac{1}{m}\right)_{\infty}-\sqrt{c_{o}}\left(1+\frac{|\alpha|n^{2}}{c_{0}}\right).$
The second inequality follows from $|f_{\alpha}|\leq\left(1+\frac{|\alpha|n^{2}}{c_{0}}\right)$. And for the first term by the same argument in Case 1,
$\,\omega_{\varphi}^{1}\left(\widetilde{f}_{\alpha},\frac{1}{m}\right)_{\infty}$
$\geq$ $\,\sup_{t\in[3.5,4]}\left|{f}_{\alpha}\left(\frac{c_{0}}{n^{2}}\left(t+\frac{1}{2}\right)\right)-{f}_{\alpha}\left(\frac{c_{0}}{n^{2}}t\right)\right|$
$=$ $\,\sup_{t\in[3.5,4]}\left|\frac{t+1/2}{t+3/2}-\frac{t}{t+1}+\frac{\alpha n^{2}}{c_{0}}\left(\frac{1}{t+1/2}-\frac{1}{t}\right)\right|$
$\geq$ $\,\frac{\alpha n^{2}}{c_{0}}\inf_{\beta\in\mathbb{R}}\sup_{t\in[3.5,4]}\left|\beta\left(\frac{t+1/2}{t+3/2}-\frac{t}{t+1}\right)+\frac{1}{t+1/2}-\frac{1}{t}\right|$
$\geq$ $\,\frac{\alpha n^{2}}{c_{0}}a_{2},$
where $a_{2}$ is a positive constant independent of $\alpha,n,c_{0}$ again since $\frac{t+1/2}{t+3/2}-\frac{t}{t+1}$ and $\frac{1}{t+1/2}-\frac{1}{t}$ are linearly independent. Hence we have for $|\alpha|>c_{0}/n^{2}$ (choose $c_{0}$ such that $\frac{a_{2}}{M}-\sqrt{c_{0}}>0$)
$\,E_{n}(f_{\alpha};[c_{0}n^{-2},1])$
$\geq$ $\,\frac{\alpha n^{2}}{c_{0}M}a_{2}-\sqrt{c_{o}}\left(1+\frac{|\alpha|n^{2}}{c_{0}}\right)$
$\geq$ $\,\inf_{|\alpha|>c_{0}/n^{2}}\frac{|\alpha|n^{2}}{c_{0}}\left(\frac{a_{2}}{M}-\sqrt{c_{0}}\right)-\sqrt{c_{0}}$
$\geq$ $\,\frac{a_{2}}{M}-2\sqrt{c_{0}}.$
Combining Case 1 and Case 2, we conclude for any $\alpha\in\mathbb{R}$
$E_{n}(f_{\alpha};[c_{0}n^{-2},1])\geq\min\left\{\frac{a_{1}}{M}-2\sqrt{c_{0}},\frac{a_{2}}{M}-2\sqrt{c_{0}}\right\}.$
Choosing $c_{0}>0$ sufficiently small completes the proof. ∎
C.4 Proof of Lemma 5
Proof.
It is known that the conditional distribution of $Y^{n}$ given the summation of its components is multinomial:
$\mathbf{Y}^{n}\mid\sum_{k=1}^{d}Y_{k}^{n}=m\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}^{m}$ (28)
Hence we have
$\,\mathbb{E}[f(Y_{1}^{n},\dots,Y_{d}^{n})]$
$\geq$ $\,\sum_{m=0}^{n}\mathbb{E}\left[f(Y_{1}^{n},\dots,Y_{d}^{n})|\sum_{k=1}^{d}Y_{k}^{n}=m\right]\mathbb{P}\left(\sum_{k=1}^{d}Y_{k}^{n}=m\right)$
$=$ $\,\sum_{m=0}^{n}\mathbb{E}\left[f(X_{1}^{m},\dots,X_{d}^{m})\right]\mathbb{P}\left(\sum_{k=1}^{d}Y_{k}^{n}=m\right)$
$\geq$ $\,\mathbb{E}\left[f(X_{1}^{n},\dots,X_{d}^{n})\right]\mathbb{P}\left(\sum_{k=1}^{d}Y_{k}^{n}\leq n\right)$
$\geq$ $\,\frac{1}{2}\mathbb{E}\left[f(X_{1}^{n},\dots,X_{d}^{n})\right]$
where the first inequality follows from truncating the summation, the first equation follows from (28), the second inequality follows from monotonicity of $\mathbb{E}[f(X_{1}^{n},\dots,X_{d}^{n})]$ and the last inequality follows from $\sum_{k=1}^{d}Y_{k}^{n}\sim\text{Poisson}(n)$ and $\mathbb{P}(\sum_{k=1}^{d}Y_{k}^{n}\leq n)\geq 1/2$. ∎